### v1

Abstract
In 2017, Bertsimas and Dunn proposed the first optimal decision tree (ODT) algorithm using general-purpose solvers, sparking significant interest in subsequent studies. However, existing combinatorial methods—primarily based on branch-and-bound (BnB) method—focus on much simpler variants, such as trees with binary feature data or axis-parallel hyperplanes. Although the first combinatorial algorithm appeared in 1989, little subsequent work has directly built upon it.Researchers have proposed various algorithms with different data structures and frameworks, but few adopt a unified framework comparable to general-purpose solvers. Many so-called optimal algorithms rely on ad-hoc processes and bounding strategies, with complex pseudo-code reducing accessibility and ambiguously stated claims hindering independent validation. As a result, prior BnB algorithms often lack the rigor to prove optimality, and algorithms claiming to solve the same problem may yield inconsistent—or even incorrect—solutions. Ultimately, optimal algorithms demand rigorous scrutiny: they must provably guarantee the optimal solution for a given search space and objective.
We argue that the root cause lies in the absence of a formal framework for designing optimal algorithms using combinatorial methods and standardized criteria for verifying correctness. Any algorithm claiming optimality must be formally proven. To address this, in the first paper (Part I) of this two-part series, we examined the ambiguity and informality in studies of ODT problems and discussed the consequences of relying on informal formalisms in optimal algorithm design. We then presented criteria that an optimal algorithm should satisfy to ensure a verifiable and transparent design process.
Beyond identifying limitations, our primary goal—and the central contribution of this work—is to develop a foundational theory for solving ODT problems. In Part I, we introduce four novel definitions of the ODT problems: three for size-constrained trees and one for depth-constrained trees. These definitions are stated unambiguously through executable recursive programs, satisfying all criteria we propose for a formal specification. In this sense, they resemble the “standard form” used in the study of general-purpose solvers.
Grounded in algebraic programming theory—a relational formalism for deriving correct-by-construction algorithms from specifications—we can not only establish the existence or nonexistence of dynamic programming solutions but also derive them constructively whenever they exist. Consequently, the four generic problem definitions yield four novel optimal algorithms for ODT problems with arbitrary splitting rules satisfies the axioms and objective functions of a given form. These algorithms encompass the known depth-constrained, axis-parallel ODT algorithm as special case, while providing a unified, efficient, and elegant solution for the general ODT problem.
In Part II, we present the first optimal hypersurface decision tree algorithm and provide comprehensive experiments against axis-parallel decision tree algorithms, including heuristic CART and state-of-the-art optimal methods. The results demonstrate the significant potential of decision trees with flexible splitting rules. Moreover, our framework is readily extendable to support algorithms for constructing even more flexible decision trees, including those with mixed splitting rules.
1 Introduction
Decision trees are widely used models in various areas of computer science, including binary space partitioning trees (Tóth, 2005; Fan et al., 2018; De Berg, 2000) and the $K$-$D$-tree (Bentley, 1975) in computational geometry, as well as decision tree models in machine learning (ML) (Breiman et al., 1984).
Due to their simplicity, decision trees are particularly attractive for providing highly interpretable models of the underlying data—an aspect that has gained increasing importance in recent years. Breiman (2001) aptly noted, “On interpretability, trees rate an A+.” The most common approach to optimizing decision trees is to employ top-down greedy algorithms, such as the Classification and Regression Tree (CART) Breiman et al. (1984) and C4.5 (Quinlan, 2014) algorithms, to optimize the tree.
However, with the rapid development of ML research, the greedy top-down optimization algorithms are no longer as effective compared with state-of-the-art ML algorithms. To improve the accuracy of decision tree models, two common approaches are typically adopted. The first is to find the globally optimal decision tree (ODT) for the given data. Alternatively, instead of restricting models to axis-parallel hyperplanes, more complex splitting rules can be applied. For example, a generalization of the classical decision tree problem is the hyperplane (or oblique) decision tree, which employs hyperplane decision boundaries to potentially simplify complex boundary structures. However, the optimization problems in both approaches are computationally challenging. For the optimal axis-parallel decision tree (AODT) problem, Laurent and Rivest (1976) showed that determining whether there exists an axis-parallel decision tree with cost less than or equal to $w$ for a given finite dataset and rule set is NP-complete. Consequently, finding an optimal axis-parallel decision tree for a given dataset is NP-hard. Similarly, for hyperplane decision trees, even the greedy top-down optimization approach is NP-hard, as the number of possible hyperplane splits for a dataset of size $N$ in dimension $D$ is $O\left(N^{D}\right)$, which grows exponentially with $D$.
Nevertheless, the stopping criteria of the CART algorithm recursively split a leaf until one of the following conditions is met: (1) the maximum tree depth is reached, (2) the maximum number of leaves is attained, (3) the minimum number of observations per leaf is violated, or (4) the minimum impurity reduction threshold is satisfied. Upon closer examination, the optimization problem that the CART algorithm attempts to solve differs from the one defined by Laurent and Rivest (1976), as the latter does not impose the constraints specified in CART. This raises an important question: is the optimization problem solved by CART also NP-hard? Interestingly, Ordyniak and Szeider (2021) showed that the decision tree problem with fixed size and depth is fixed-parameter tractable. In other words, if the size or depth of the tree is fixed, there exists a polynomial-time algorithm for solving the problem.
While the result of Ordyniak and Szeider (2021) leaves certain aspects underspecified—for example, some key definitions regarding the problem formulation—it nevertheless offers a valuable perspective on the tractability of decision tree optimization. As a result, the fixed-parameter tractability of ODT problems has led to a significant growth of research interest in recent years. One major line of work focuses on formulating the ODT problem as a mixed-integer programming (MIP) problem, which can then be solved using general-purpose optimization solvers. Following the first MIP formulation proposed by Bertsimas and Dunn (2017), a surge of subsequent studies has investigated this approach (Boutilier et al., 2023; Verwer and Zhang, 2017, 2019; Boutilier et al., 2022; Zhu et al., 2020) leveraging general-purpose MIP solvers. Notably, Bertsimas and Dunn (2017)’s algorithm can not only solve the AODT problem but also the optimal hyperplane decision tree problem.
An alternative line of research, which we refer to as combinatorial methods, seeks to employ approaches such as dynamic programming (DP), greedy strategies, or branch-and-bound (BnB) techniques to exhaustively explore the solution space while eliminating non-optimal solutions through bounding techniques. Unlike the advances made in the study of general-purpose solvers, where many practical solutions have been developed for both the AODT problem and the hyperplane decision tree problem. Although the first combinatorial algorithm for decision tree optimization was introduced as early as 1989 (Cox Jr et al., 1989), most subsequent studies have focused on simpler settings, such as the ODT problem over binary feature data. Only a few studies, such as Brița et al. (2025); Mazumder et al. (2022), have addressed the AODT problem, whereas algorithms for the hyperplane optimal decision tree problem remain unexplored.
Unlike the work of Bertsimas and Dunn (2017) on the ODT problem, which has inspired a substantial body of subsequent research, many later studies on combinatorial algorithms adopt distinct problem definitions. For instance, the approaches of Demirović et al. (2022); Verwer and Zhang (2019); Nijssen and Fromont (2007); Aglin et al. (2020); Nijssen and Fromont (2010) construct optimal decision trees by constraining the tree depth. In contrast, Zhang et al. (2023); Lin et al. (2020); Hu et al. (2019) propose algorithms that constrain the number of leaves, referring to the resulting models as sparse decision trees.
Additionally, the combinatorial methods developed to solve the ODT problem are highly diverse. For example, Nijssen and Fromont (2007); Aglin et al. (2021, 2020); Nijssen and Fromont (2010) characterize decision trees over binary feature data using the concept of itemsets from data, Demirović et al. (2022) represent decision trees via feature vectors, whereas Lin et al. (2020); Hu et al. (2019) define trees in terms of two sets of leaves: those that can be further split and those that are fixed. Unfortunately, the advantages and limitations of these data structures remain insufficiently studied. If these distinct representations of decision trees prove to be indispensable—playing central roles in their respective algorithms and serving different purposes in tree construction—their strengths and weaknesses warrant careful analysis.
At the same time, while using general-purpose MIP solvers for the ODT problem is appealing and can be efficient in many cases, the overall computational complexity remains unpredictable. More importantly, MIP-based approaches provide limited insight into the combinatorial and geometric structure of the problem. Their performance largely depends on advances in general-purpose solvers and, potentially, more sophisticated MIP formulations. To advance theoretical understanding and develop practical algorithms, it is therefore essential to study combinatorial methods, which provide a more direct view of the underlying geometry and algorithmic structure.
In this study, we focus on combinatorial methods for solving the ODT problem, following the axiomatic definition of decision trees established in our previous work He and Little (2025b), which we refer to as proper decision trees, we provide a detailed analysis of the algorithmic, combinatorial, and geometric properties of the ODT problems. Our generic solution not only offers several new perspectives on decision tree optimization but also provides a rigorous framework for future analysis. Building on the algorithmic and geometric results discussed in this paper, another key contribution of this work is demonstrating that the hypersurface decision tree (HODT) model satisfies the axioms we proposed and can therefore be solved using the four ODT algorithms introduced here. Moreover, our algorithms are readily applicable to more flexible decision tree models that incorporate mixed splitting rules. In Part II, we will present the detailed construction of the first optimal hypersurface decision tree (HODT) algorithms, highlighting their practical implementation and applicability.
We believe this research lays a theoretical foundation for the study of ODT problems. A summary of our contributions is as follows:
Achievement in Part I
  • Critique and proposed solutions: We critically examine current approaches to designing algorithms and defining problems in the study of optimal decision trees, highlighting the consequences of relying on informal problem formulations paired with ad-hoc algorithmic solutions—a practice common in combinatorial studies of ODT problems, particularly those employing BnB methods. Beyond this critique, the central distinction of our research lies in philosophy: we argue that deriving truly optimal algorithms demands the highest levels of rigor and clarity. Accordingly, we propose criteria for formally defining problems in the study of optimal algorithms, along with the key characteristics that a good optimal algorithm should possess. These discussions provide a foundational step for advancing research on ODT problems using combinatorial methods.
  • Formal combinatorial definitions: We provide four formal combinatorial definitions for the ODT problem—covering both size- and depth-constrained trees—ensuring unambiguous communication for future research. All definitions are directly executable, allowing verification that the generated trees align with expectations. To our knowledge, no prior studies on using BnB algorithms for ODT have provided a formal problem definition with this level of rigor.
  • Generic optimal decision tree algorithms: We present four novel and generic algorithms for solving ODT problems with arbitrary splitting rules that satisfy our axioms, capable of optimizing arbitrary objective functions of a specified form. All algorithms are constructively derived in a formal manner. Specifically, we provide three algorithms for size-constrained trees and one for depth-constrained trees, formally proving and extending results from previous studies on axis-parallel ODT problems Brița et al. (2025).
  • Filter and thinning fusion: We propose two fusion theorems—filtering and thinning—to enhance the flexibility and practical applicability of ODT algorithms while maintaining formal rigor. The filtering fusion theorem allows additional constraints, such as depth and leaf-size constraints, to be incorporated into ODT algorithms defined over size-constrained trees. Similarly, tree-size constraints can be applied to ODT algorithms defined over depth-constrained trees. The thinning fusion theorem integrates well-known bounding techniques into the framework. Together, these theorems extend the practical utility of the algorithms without compromising optimality.
  • Proving the properness of hypersurface decision trees: By generalizing Xi’s 0-1 loss linear/hypersurface classification theorem from linear models to decision tree models, and combining it with the geometric characterization of polynomial hypersurfaces as combinations of data points, we demonstrate that hypersurface decision trees are proper decision trees and can therefore be solved exactly using our proposed algorithms.
  • Splitting rule generators: Leveraging the sophisticated combination generator introduced by He and Little (2025a), we develop efficient algorithms for generating both hypersurface splitting rules and mixed-splitting rules for a given dataset. The hypersurface splitting rule generator provides the foundation for constructing efficient hypersurface decision tree algorithms developed in Part II, while the mixed-splitting rule generator establish a foundation for future work on constructing optimal ODT algorithms with mixed splitting rules.
  • Complexity analysis for optimal decision tree problems: Most studies on ODT algorithms either do not analyze worst-case complexity or analyze it using incorrect combinatorial characterizations. Our analysis provides an upper bound for size-constrained decision trees, which also implies trivial lower bounds for depth-constrained axis-parallel decision tree problem. Moreover, this lower bound contradicts the function for counting the combinatorial complexity of the problem in Mazumder et al. (2022)’s source code, which may indicate a potential flaw regarding the exhaustiveness of their algorithm.
  • Challenge to existing algorithms: Because our algorithms are derived from a formal axiomatic definition of decision trees, they challenge prior studies that applied a similar algorithmic process to ODT problems with binary feature data (Demirović et al., 2022; Hu et al., 2019; Lin et al., 2020; Zhang et al., 2023; Aglin et al., 2021, 2020). These studies do not satisfy the axioms we propose, which are essential for justifying the correctness of a critical component, yet this component remain unspecified or ambiguously described in their constructions. It is therefore unclear whether their ambiguous algorithmic processes can support the claimed optimality of these algorithms, as none formally specified their problem or proved correctness. Our results call into question the validity of these claims.
Achievement Part II
  • Crossed-hyperplane properties: We identify the crossed-hyperplane properties, which significantly reduce the size of the search space for the HODT problem.
  • Incremental algorithm for ancestry matrix generation: Leveraging the prefix-closed property of crossed hyperplanes, we introduce an incremental algorithm for efficiently generating the ancestry relation matrix, which is crucial for solving size-constrained decision tree problems. This algorithm can be fully vectorized and massively parallelized.
  • Application of size-constrained ODT algorithms to HODT problem: We apply the two rigorously proven algorithms from Part I to solve the size-constrained HODT problem, encompassing classical axis-parallel hyperplanes, general hyperplanes, and hypersurfaces. Both algorithms have polynomial complexity (for fix tree size and data dimension), are easily parallelizable, and amenable to speed-up techniques.
  • Experiments on synthetic data: Using the two heuristic methods we proposed, we conduct comprehensive experiments on hyperplane decision algorithms with synthetic datasets, where the ground truth is generated by hyperplane decision trees (i.e., polygonal regions). The results show that hyperplane decision trees recover the ground truth more accurately than both heuristic methods and optimal axis-parallel tree algorithms under various noise conditions. Notably, our algorithms can even achieve solutions that are more accurate than the ground truth on noisy training data, refuting the notion that optimal algorithms necessarily overfit.
  • Benchmarking on real-world datasets: We perform a comprehensive analysis on 30 real-world datasetss, comparing hypersurface decision tree algorithms with CART and Brița et al. (2025)’s AODT algorithm (ConTree algorithm). The results demonstrate that hyperplane decision trees achieve superior generalization accuracy compared with axis-parallel decision trees when tree size is properly controlled.
A summary of Part I of this paper is as follows. In Section 2, we examine the informality of problem definitions in existing studies of optimal decision tree algorithms. Subsection 2.1 analyzes the consequences of adopting such informal definitions. In Subsection 2.2, we discuss potential remedies and outline the desired characteristics of a sound optimal algorithm. Section 3 introduces the notations and auxiliary functions used throughout the paper.
Section 4 presents the main algorithmic foundations. We begin by defining a proper decision tree, which allows decision trees to be formalized using datatypes and axioms. Based on these axioms, Subsection 4.2 constructs four formal definitions of both size-constrained and depth-constrained ODT problems, each yielding a generic solution algorithm. The construction of these algorithms is detailed in Subsection 4.3. In Subsection 4.4, we extend the framework by introducing filtering and thinning algorithms, enabling constraints and bounding techniques to be applied via simple conditions derived from our theorems. Subsection 4.5 provides a thorough comparison of the four algorithms.
In Section 5, we extend the results of He and Little (2023) to establish the geometric foundation for hypersurface decision tree problems. We show that, when the data are in general position, decision trees with hypersurfaces defined by data points are proper. Building on the combination and nested combination generators of He and Little (2025a), Subsection 5.2 introduces two splitting rule generators: one for hypersurface splits and another for mixed splits, allowing a single tree to incorporate axis-parallel, hyperplane, and hypersurface divisions.
Finally, Section 6 presents the conclusion.
2 How to define an optimal decision tree problem formally and why it matters
A useful analogy for understanding formal problem definitions comes from the study of general-purpose solvers, where strict rules ensure that problems are expressed in a precise form, often referred to as the “standard form.” For instance, the standard form of a linear programming problem requires the objective function to be minimized, all constraints to be equalities, and all variables to be non-negative. These strict requirements ensure unambiguous communication among researchers and reproducibility of results.
In contrast, there appears to be no widely accepted standard for formally defining optimal algorithms that use combinatorial methods. To appreciate the importance of formal definitions, it is first necessary to clarify what constitutes an informal definition and why such informality can lead to significant consequences. We then discuss the key criteria that must be satisfied to resolve ambiguity, as well as the characteristics that a good optimal algorithm should possess.
2.1 Informality and consequences
Data back to the dawn of 20th century, after Gödel (1931) announced his well-known incompleteness theorem, there is a problem determining a formal definition of “effectively calculable.” To address this, three foundational models of computation were proposed: lambda calculus, invented by (Church, 1936), general recursive functions, proposed by Gödel (Gandy, 1988), and Turing machine, by Turing et al. (1936). Church give his definition of $\lambda$-definability and claim all calculation that is $\lambda$-definable are effective calculable. However, his argument was unsatisfactory to Gödel. As (Wadler, 2015) noted:“ Church merely presented the definition of $\lambda$-definability and baldly claimed it corresponded to effective calculability.”
In contrast, Turing conducted a detailed analysis of the capabilities of a “computer”—at the time, the term referred to a human performing a computation assisted by paper and pencil. Gandy (1988) pointed out that Turing’s argument amounts to a theorem asserting any computation a human with paper and pencil can perform can also be performed by a Turing machine. It was finally Turing’s argument convinced Gödel, and eventually prove the three definitions had been proved equivalent.
Returning to the study of the optimal decision tree problem, a similar issue arises in the design of optimal algorithms using combinatorial methods. In many studies, algorithms are directly defined and claimed to be optimal for a given problem, yet a rigorous problem definition is never formally established. While precise problem definitions may seem negligible for most approximate or heuristic algorithms in machine learning, their absence can have serious consequences in the design of true optimal algorithms. Informality in problem definition not only poses risks for high-stakes applications that require strict optimality but can also mislead subsequent research. This is because the design of globally optimal algorithms has a singular goal: to identify a configuration (or multiple configurations with the same optimal objective) over the entire search space, for which there exists a unique optimal value.
Unfortunately, a common trend in the studies of optimal algorithms using combinatorial methods are to focus on proving bounding techniques without establishing the correctness of the main algorithm. However, establishing the optimality of an algorithm requires more than just validating a bounding technique. In essence, much of this work amounts to the design of “optimal bounding techniques” rather than the development of truly “optimal algorithms” for the decision tree problem. These are fundamentally different research directions, as a bounding technique is meaningful only when it supports a provably correct algorithm in the first place. We believe this inclination stems from a programming philosophy that prioritizes efficiency as the primary, and sometimes sole, objective. This tendency is further compounded by the lack of widely accepted criteria for defining an optimal algorithm.
What is an “informal” definition, and what are the consequences of adopting such a definition in optimal algorithm design?
For instance, studies of ODT algorithms using combinatorial methods often leave the problem undefined [^1] We claim that these studies leave the problem undefined because they commonly do not specify the solution space over which they are searching for an optimal decision tree. In some cases, the space is not even ambiguously defined, as in specification (1), making it unclear which ODT problem they are actually attempting to solve. (Lin et al., 2020; Zhang et al., 2023; van der Linden et al., 2023; Demirović et al., 2022; Hu et al., 2019) or specify their problem ambiguously, as in (Mazumder et al., 2022; Brița et al., 2025), where the problem is expressed in a form similar to
$s^{*}=\text{argmin}_{s\in\mathcal{S}}\sum_{n\in\mathcal{N}}E\left(s\left(x_{n}\right)\neq t\right)$ (1)
where $\mathcal{S}$ is often defined as the search space of all possible decision trees with a constrained depth or size. $E$ is the objective function, which is often defined as the 0-1 loss $E_{\text{0-1}}=\mathbf{1}\left(s\left(x_{n}\right)\neq t\right)$ (count the number of misclassification) (counting the number of misclassifications) for classification problems, and the squared loss $E_{\text{0-1}}=\left(s\left(x_{n}\right)\neq t\right)^{2}$ for regression problems.
Although (1) appears intuitively obvious, it suffers from several major flaws. First, the search space $\mathcal{S}$ usually denoted by a symbol and defined through text description. For example, Brița et al. (2025) wrote:“ $\mathcal{T}\left(\mathcal{D},d\right)$ describe the set of all decision trees for the datasets $\mathcal{D}$ with a maximum depth of $d$ .” Adopting such a textual description for $\mathcal{T}\left(\mathcal{D},d\right)$ is inherently ambiguous and dangerous. Even though it appears simple, different researchers may interpret the search space differently due to variations in the constraints imposed on the trees, and some may even adopt entirely different definitions of the decision tree. In contrast, MIP specifications force researchers to define precisely what they mean by a decision tree by specifying the constraints first, enabling an unambiguous definition.
As a result, studies that adopt such informal problem definitions cannot verify the most important step in proving the optimality of an algorithm—showing that the algorithm provably searches all possible solutions in $\mathcal{S}$ ($\mathcal{T}\left(\mathcal{D},d\right)$ in Brița et al. (2025)), because $\mathcal{S}$ is not formally defined. Consequently, research on BnB algorithms often focuses solely on proving the correctness of the bounding techniques and then claims the algorithm is “optimal.”
Additionally, the ambiguous problem definition inevitably leads to inconsistencies in the study of optimal algorithms. For example, although Mazumder et al. (2022) and Brița et al. (2025) claim to solve the same problem, we found that their algorithms return different trees with significantly different objective values on the same datasets—differences too large to be attributed to numerical issues. This suggest that either Mazumder et al. (2022)’s algorithm is non-optimal or that the two studies are solving different ODT problems. Evidence favors the first explanation: a function in Mazumder et al. (2022)’s source code counts the number of possible trees in the search space, which contradicts the lower bound we provide in Subsection 5.3.
Secondly, for most objectives in machine learning, comparisons between two configurations are defined by a preorder: for any pair of configurations $a$ and $b$, $a\leq b$ and $b\leq a$ does not implies $a=b$. In other words, two configurations with the same objective value are not necessarily identical. As a result, the argmin over a preorder does not always yield a unique solution. How we define argmin thus becomes critical, especially in ML applications where we care about the predictive outcomes of the solution: different configurations with the same objective can lead to completely different predictions.
Furthermore, interpretations of what constitutes an “optimal algorithm” vary. Some restrict optimality to a specific objective, while others do not. For instance, in classification tasks, most researchers consider the solution minimizing the 0-1 loss as the standard for optimality, because optimizing convex surrogates such as hinge or squared loss is relatively straightforward. This explains why the support vector machine, which is optimal for hinge loss, is not typically claimed as “optimal algorithm.” In contrast, some studies on empirical risk minimization problem for two-layer neural networks (Arora et al., 2016; Pilanci and Ergen, 2020) optimize convex losses but claim the optimality. This highlights the importance of establishing a common and unambiguous definition of the objective function to avoid misunderstandings about the problem setting.
Finally, if argmin is left undefined, then $E$ is often undefined as well. Although the definitions (or implementations) of $E$ and argmin may seem trivial, they are in fact crucial for proving the correctness of an algorithm. As we will see in our exploration, the existence of a dynamic programming or greedy algorithm is almost entirely determined by the structure of $E$ and $\mathcal{S}$; in other words, it is governed by the structural compatibility between $E$ and $\mathcal{S}$.
2.2 What should we do next?
The difficulty I foresaw were a consequence of the absence of generally accepted criteria, and although I was convinced of the validity of the criteria I had chosen to apply, I feared that my review would I be refused or discarded as “a matter of personal taste.” I still think that such reviews would be extremely useful and i am longing to see them appear, for their accepted appearance would be a sure sign of maturity of the computing community. — Dijkstra (1972)
Criteria of formal specification
In Dijkstra (1972)’s 1972 talk, “ The Humble Programmer,” Dijkstra criticized the programming philosophy that prioritizes efficiency optimization above all else, arguing that it leads to unstructured designs that are hard to program. Similarly, in our study of optimal algorithms for machine learning, an exclusive focus on efficiency when solving optimal algorithms has led to many oversights. These issues, I argue, stem from the facts that algorithm designers also prioritizes the theoretical computational efficiency and absence of widely accepted criteria in the study of optimal algorithms, concerns echoed in Dijkstra’s discussion.
The lack of standardized criteria for validating optimality in studies of combinatorial methods for optimal algorithms has consequences reminiscent of Dijkstra’s observations. I foresaw that my critique of this issue might be dismissed as subjective, much like Dijkstra’s fear that his reviews would be rejected as “a matter of personal taste.” Despite this, I believe such critical reviews are essential and would signal the maturity of the optimization community if widely accepted.
Here, I summarize several key criteria that are essential for the design of optimal algorithms but have often been overlooked. I believe that omitting discussion of these points can lead to severe consequences for the reproducibility and verifiability of optimal algorithm design. To bolster the credibility of these arguments, I draw parallels with insights from Dijkstra (1972) insights. These considerations can be broadly categorized as follows:
  1. 1.
    Exact algorithms must be provable correct: This is the central philosophy we aim to convey: an exact algorithm must be provably correct, which demands the highest level of rigor in establishing its correctness. As Dijkstra emphasized:“ We must not forget that it is not our business to make programs; it is our business to design classes of computations that will display a desired behavior.” When designing optimal algorithms using combinatorial methods, it is essential to rigorously prove that the algorithm exhibits the intended behavior. Specifically, an optimal algorithm must be proved to either (i) exhaustively explore all valid solutions to guarantee identification of the optimal one with respect to the given objective, or (ii) soundly eliminate non-optimal solutions. Without such proof, any claim of optimality remains unfounded.
  2. 2.
    Proof requires an unambiguous problem specification: Logically, a specification (definition) cannot be “proved”; it must be defined or introduced as clearly as possible to allow for verification and rigorous reasoning. In computer science, one of the clearest ways to specify a problem is through an actual program rather than pseudocode, which can obscure crucial details and hinder reproducibility. For combinatorial problems, it is essential to answer: What exactly is the search space? and What objective function is being optimized? Ideally, the search space and objective function should either be defined by a program that are immediately executable or expressed in a standard (MIP) form for studies using general-purpose solvers. Without such clarity, inconsistencies in problem definitions and their corresponding solutions are inevitable.
  3. 3.
    Algorithms should be intellectually manageable: Although a common strategy in algorithm design is to first construct a program and then verify that it satisfies the required properties, this process is often cumbersome and error-prone. As Dijkstra (1972) famously quote:“ Program testing can be a very effective way to show the presence of bugs, but it is hopelessly inadequate for showing their absence. The only effective way to raise the confidence level of a program significantly is to give a convincing proof of its correctness. But one should not first make the program and then prove its correctness, because then the requirement of providing the proof would only increase the poor programmer’s burden.” Similarly, research on combinatorial methods—particularly BnB algorithms—often adopts the practice of constructing ad-hoc algorithms and proving their correctness post-hoc using complex proofs (or simply without proof). However, this approach is problematic. Complex algorithms typically result in complex proofs, which not only make the algorithms less accessible but also harder to verify. For instance, Demirović et al. (2022)’s algorithm spans nearly five pages of pseudocode, making it intellectually challenging to understand, implement, and verify. To address this, Dijkstra—and likewise, this work—advocates a constructive approach to programming: derive programs from well-defined specifications, thereby ensuring correctness by design. In combinatorial optimization, we begin with an unambiguous brute-force specification of the optimization problem and systematically derive a more efficient algorithm. This guarantees that the final implementation inherits the properties of the original specification.
  4. 4.
    The importance of complexity analysis: This issue arises from earlier challenges, such as the lack of clear problem definitions and bounding techniques, which make formal complexity analysis infeasible. The situation is further exacerbated by the fact that many ad-hoc algorithms designed in BnB research are not intellectually manageable. For instance, many operational components—such as the use of depth-first search strategies—are implicitly embedded in the pseudocode of the designed algorithms. While these additional ingredients may provide advantages, such as reducing runtime memory usage, they also introduce side effects: the resulting algorithms become harder to analyze formally. Consequently, most studies on BnB algorithms do not provide time or space complexity analyses. However, complexity analysis remains critically important for algorithm design. Without it, the scalability and practical applicability of the algorithm remain unpredictable.
  5. 5.
    Acceleration techniques must be demonstrably effective: This may seem counterintuitive: if a technique provably reduces the search space, how could it be ineffective? In practice, the overhead of applying the technique can outweigh its benefits. The key is to balance the cost of applying an acceleration method against the potential reduction in search space. For example, Brița et al. (2025) claim that caching is a major factor behind their algorithm’s efficiency. However, He and Little (2025b) provide a simple analysis showing that overlapping subproblems in decision tree construction are rare—particularly for the ODT problem with axis-parallel or more complex splitting rules—making caching or memoization largely ineffective. In our experiments, we even observed cases where Brița et al. (2025)’s algorithm explored over a million decision trees without a single cache hit, suggesting that in such scenarios, disabling caching entirely may improve performance.
  6. 6.
    Designing algorithms and designing bounding techniques are distinct tasks: As discussed, there is widespread confusion in the community regarding the distinction between designing optimal algorithms and developing bounding techniques for them. Bounding techniques are only meaningful when applied to algorithms that are already optimal. Therefore, research on these two aspects should be analyzed and presented separately. Studies on bounding techniques should focus on assessing their effectiveness by providing theoretical guarantees or empirical evidence. In contrast, the design of optimal algorithms must first establish their correctness and optimality, followed by a rigorous analysis of their time and space complexity as well as their empirical efficiency—entirely independent of any bounding techniques.
Characteristics of an good optimal algorithm
Having discussed the critical criteria for designing optimal algorithms, a natural question arises: if several algorithms satisfy these criteria, which one is the best, and what characteristics should a good optimal algorithm possess? Indeed, many algorithms have been proposed for the optimal decision tree problem, but identifying the best approach is often challenging, as comparisons are frequently conducted unfairly or omit essential considerations.
By contrast, in the study of sorting algorithms, several well-established characteristics allow researchers to assess an algorithm’s quality systematically.A good sorting algorithm typically aims to satisfy four key properties, though achieving all simultaneously is nontrivial:
  1. 1.
    Fast: The algorithm should be asymptotically optimal in the number of comparisons, and the constants involved in other operations should also be small.
  2. 2.
    Smooth: Performance should improve on partially sorted inputs; the more sorted the input, the faster the algorithm performs.
  3. 3.
    Stable: Records with equal keys should retain their relative order in the output.
  4. 4.
    Compact: The algorithm should be economical in both space and runtime.
For example, mergesort algorithm is fast, stable, and compact, but not smooth, as it maintains $O\left(N\log\left(N\right)\right)$ complexity even for already sorted inputs.
By analogy, we believe a good optimal algorithm also should satisfy the following properties: provably optimal, conceptually simple, computationally efficient, and inherently parallelizable. In the context of ML research, such as the ODT problem, it is also important to consider an additional property: the generality of the solution. Optimizing a single objective or specific splitting rules does not guarantee robust performance across diverse scenarios. An algorithm capable of handling flexible objective functions or constructing decision trees with complex splitting rules is typically more practical and broadly applicable. While we have already discussed the importance of simplicity and provable optimality, we now clarify what we mean by efficiency, generality, and parallelizability.
  1. 1.
    Efficiency: The efficiency of an optimal algorithm should not be judged solely by its theoretical worst-case complexity. Factors such as hidden constants in the big-O notation, and more importantly, compatibility with hardware and software, significantly influence practical performance. Algorithms with predictable blocked array usage and structured computational processes often achieve higher efficiency in practice. Thus, we argue that designing hardware-friendly algorithms is as important as designing algorithms that achieve superior theoretical efficiency. In Part II, we demonstrate that an algorithm with superior theoretical performance may still run slower than a less theoretically efficient one if the latter is better aligned with hardware structure.
  2. 2.
    Abstraction and generality. Abstraction not only simplifies complex problems but also enhances the generality of solutions. As Dijkstra (1972) noted, “the effective exploitation of his powers of abstraction must be regarded as one of the most vital activities of a competent programmer. the purpose of abstracting is not to be vague, but to create a new semantic level in which one can be absolutely precise.” In algorithm design, abstraction provides substantial advantages. In Part I, we will demonstrate that adopting a generic principle can lead to a deeper understanding of problem structure, ultimately enabling a generic algorithm capable of solving the ODT problem with arbitrary splitting rules and a broad class of objective functions.
  3. 3.
    Parallelism: For problems with intractable combinatorics—such as NP-hard optimization problems—parallelism is perhaps the only practical way to scale the algorithm to handle large scale datasetss. We do not merely advocate designing complex algorithms and parallelizing them post-hoc; rather, we argue that algorithms should be inherently parallelizable from the beginning of the design process. In other words, parallelizability should be treated as a fundamental property during algorithm design. Designing inherently parallel algorithms while maintaining optimality is challenging and requires a deep understanding of both the problem and algorithmic structure. The success of neural networks illustrates this well: GPU hardware significantly enhanced their effectiveness by aligning with their computational structure. Similarly, we argue that designing hardware-friendly algorithms is as important as achieving theoretical efficiency.
Rather than merely offering criticism or vacuous suggestions, we aim to demonstrate the effectiveness of the criteria we proposed. In the following paper, we formally address the ODT problem, thereby ensuring that these criteria are satisfied and that the resulting algorithms are intellectually manageable. We show that approaching the problem in such an intellectually manageable manner not only simplifies the derivation and makes it more accessible, but also enables the construction of more generic and efficient solutions that exhibit the desired properties of a “good optimal algorithm” as we proposed—even though not all properties may be fully achieved simultaneously.
3 Notations and auxiliary functions
3.1 Notations
In this paper, the symbol :: is used in two distinct contexts. When appearing in code expressions, :: denotes the list-construction (cons) operator, which prepends an element to an existing list. For example, $a:\left[b,c\right]=\left[a,b,c\right]$. When appearing in type signatures, the same symbol :: specifies the type of an expression; for example, $x:\mathbb{N}$ means “$x$ is a natural number.”
The types of real and natural numbers are denoted by $\mathbb{R}$ and $\mathbb{N}$, respectively. We use square brackets $\left[\mathcal{A}\right]$ to denote the set of all finite lists of elements $a:\mathcal{A}$, where $\mathcal{A}$ (or letters $B$ and $\mathcal{C}$ at the front of the alphabet) represent type variables. A data point is denoted as a $D$-tuple $x=\left(x_{1},\ldots,x_{D}\right):\mathbb{R}^{D}$. Hence, $\left[\mathcal{R}\right]$ and $\left[\mathbb{R}^{D}\right]$ (We use $\mathcal{D}$ as a short-hand synonym for $\left[\mathbb{R}^{D}\right]$), $\left[\mathcal{H}^{M}\right]$, denote the set of all finite lists of splitting rules, lists of data in $\mathbb{R}^{D}$, and hypersurfaces defined by degree $M$ polynomials, respectively. Two special type synonyms, $C=\left[\mathbb{N}\right]$ and $NC=\left[\left[\mathbb{N}\right]\right]$ are used to denote combinations of natural numbers and nested combinations (i.e., combinations of combinations).
Variables of these types are denoted using their corresponding lowercase letters; for example, $r:\mathcal{R}$, $h:\mathcal{H}$, $h^{M}:\mathcal{H}^{M}$. In particular, a degree-one polynomial hypersurface is defined as $h=\boldsymbol{w}^{T}x+c$ , where $\boldsymbol{w},x:\mathbb{R}^{D}$ and $c\in\mathbb{R}$. Such a hypersurface is known as a hyperplane; we denote $\mathcal{H}^{1}=\mathcal{H}$ by default. Similarly, $\mathcal{H}^{0}$ denotes the set of all axis-parallel hyperplanes, where an axis-parallel hyperplane corresponds to fixing one coordinate xi to a constant value, defined as $h=\boldsymbol{e}_{i}^{T}\boldsymbol{x}+c$, with $e_{i}$ being a unit vector along the $i$-th coordinate axis.
We denote a list of elements of the same type as $\mathit{as}:\left[\mathcal{A}\right]$; for instance, $rs:\left[\mathcal{R}\right]$, $xs:\mathcal{D}$, $hs:\left[\mathcal{H}\right]$, and $\mathit{hs}^{M}:\left[\mathcal{H}^{M}\right]$. The size of a list as is denoted $\left|as\right|$ , and $as!d$ indexes the $d$-th element of the list. Similarly, $\mathit{ass}$ denotes a list of lists of elements of the same type. The operations +++\!\!+ list concatenation; for example, $a:\left[b,c\right]=\left[a\right]+\!\!+\left[b,c\right]=\left[a,b,c\right]$.
Both the affine $D$-space, for $D\geq 1\in\mathbb{N}$, and the vector $D$-space over $\mathbb{R}$ are denoted by $\mathbb{R}^{D}$. Both are the set of all $D$-tuples of elements of $\mathbb{R}$. To distinguish a vector in the vector space and a point in affine space, an element $x=\left(x_{1},x_{2}\ldots,x_{D}\right)\in\mathbb{R}^{D}$ will be called a point in the affine space $\mathbb{R}^{D}$, an element $x=\left(x_{1},x_{2}\ldots,x_{D}\right)^{T}\in\mathbb{R}^{D}$ is called a vector in the vector space $\mathbb{R}^{D}$, where $x_{i}$ is called the coordinate of $x$. Note that, except for the data point vector $x$, we denote all other vectors, such as the normal vector $\boldsymbol{w}$, and the polynomial coefficient vector $\boldsymbol{\alpha}$, using bold symbols. The data point $x$ is written in plain format, rather than bold, to maintain consistency with the data list $\mathit{xs}$. Moreover, when referring to a data point vector, we omit subscripts to avoid confusion with coordinate indices $x_{i}$.
3.2 Function deifinitions and point-free style functional composition
One of the most common ways of defining a list or set is through list/set comprehension. List/set comprehension is a mathematical notation for defining a set by specifying a distinct property that all its members must satisfy. Set comprehension takes the form
$\left\{f\left(x\right)\mid x\in\mathcal{X},p\left(x\right)=\text{True}\right\}.$
This is read as:“ The set of all outputs $f\left(x\right)$ such that element $x$ is a member of the source set $\mathcal{X}$ and the predicate $p\left(x\right)=\text{True}$ holds.” Similarly, list comprehension takes the form
$\left[f\left(x\right)\mid x\leftarrow\mathit{xs},p\left(x\right)=\text{True}\right].$
This is read as:“ The list of all output $f\left(x\right)$ such that element $x$ is drawn from source list $\mathit{xs}$ sequentially and the predicate $p\left(x\right)=\text{True}$ holds.”
Throughout the discussion, we will use some common functions from functional programming languages. One frequently used function is the map function over list
$\mathit{mapL}\left(f,\mathit{xs}\right)=\left[f\left(x\right)\mid x\leftarrow\mathit{xs}\right],$
which applies function $f:\mathcal{A}\to\mathcal{B}$ to each element $x$ in the list $\mathit{xs}$.
Similarly, another common function used frequently in our exploration is the concatenate (or flatten) function
$\mathit{concat}=[x\mid x\leftarrow\mathit{xs},\mathit{xs}\leftarrow\mathit{xss}],$
which flattens a list of lists into a single list. These two functions can also be defined recursively. For example, the map function can be written as:
$\mathit{mapL}$ $:\left(\mathcal{A}\to\mathcal{B}\right)\times\left[\mathcal{A}\right]\to\left[\mathcal{B}\right]$
$\mathit{mapL}$ $\left(f,\mathit{xs}\right)=\left[f\left(x\right)\mid x\leftarrow\mathit{xs}\right]$
and the concatenate function as
$\mathit{concat}$ $:\left[\left[\mathcal{A}\right]\right]\to\left[\mathcal{A}\right]$
[…]
$\mathit{concat}$ $\left(\mathit{as}:\mathit{ass}\right)=\mathit{as}+\!\!+\mathit{mapL}\left(\mathit{as}\right)$
Another useful function is the map function over the tree datatype, which can be defined as follows
$\mathit{mapD}$ $:\left(\mathcal{B}\to\mathcal{C}\right)\times\mathit{DTree}\left(\mathcal{A},\mathcal{B}\right)\to\mathit{DTree}\left(\mathcal{A},\mathcal{C}\right)$
$\mathit{mapD}$ $\left(f,\mathit{DL}\left(a\right)\right)=\mathit{DL}\left(f\left(a\right)\right)$
$\mathit{mapD}$ $\left(f,\mathit{DN}\left(u,r,v\right)\right)=DN\left(\mathit{mapD}\left(f,u\right),r,\mathit{mapD}\left(f,v\right)\right),$
where the “decision tree datatype” $\mathit{DTree}\left(\mathcal{A},\mathcal{B}\right)$ will be formally introduced when we define decision trees.
To maintain the elegance of function definitions, we adopt point-free style programming paradigm, a style in which functions are defined without explicitly naming their arguments. Instead, functions are expressed solely through the composition of other functions, chaining transformations without introducing intermediate variables.
For example, the function $\mathit{concatMap}$ can be defined as
$\mathit{concatMap}_{f}=\mathit{concat}\circ\mathit{mapL}_{f},$
where $\circ$ denotes function composition. This means that $\mathit{mapL}_{f}$ is applied first, followed by $\mathit{\mathit{concat}}$. Here, $\mathit{mapL}_{f}$ is an example of a section—a technique in Haskell for partially applying an operator to create a new function by fixing one of its arguments. Specifically, $\mathit{mapL}_{f}$ is a partial application of the $\mathit{mapL}$ function, parameterized by a function $f:\mathcal{A}\to\mathcal{B}$, resulting in a new function of type $\mathit{mapL}_{f}:\left[\mathcal{A}\right]\to\left[\mathcal{B}\right]$. In the definition of $\mathit{concatMap}_{f}$, the input list $\mathit{xs}$ is not explicitly mentioned; this is why the style is called point-free. Conceptually, $\mathit{concatMap}_{f}$ applies $f$ to each element of the input list and concatenates the resulting lists.
More generally, when a function $g:\mathcal{X}\times\mathcal{Y}\to\mathcal{Z}$ is partially applied to an argument $x:\mathcal{X}$, we obtain a new function $g_{x}:\mathcal{Y}\to\mathcal{Z}$. However, if the subscript of a function name is already used for descriptive purposes—for example, $g_{\text{SomeName}}\left(x,y\right):\mathcal{X}\times\mathcal{Y}\to\mathcal{Z}$—we denote its partially applied form as $g_{\text{SomeName}}\left(x\right):\mathcal{Y}\to\mathcal{Z}$ to avoid ambiguity.
We acknowledge that point-free style may be unfamiliar to some readers. Whenever a function is expressed in point-free form, we will provide its type whenever possible to clarify its behavior.
4 Algorithmic foundation
4.1 Definition of decision tree
Before discussing our formal definition of decision trees and the ODT problem, we first extend the specification in (1) by introducing additional constraints. We then progressively explain how to define decision trees and decision tree problems rigorously.
The ODT problem seeks to identify a binary tree $T$ that minimizes the number of misclassifications, where the internal nodes of $T$ are characterized by a list of splitting rules $\mathit{rs}:\left[\mathcal{R}\right]$. In our exposition, $\mathit{rs}$ is a finite set generated from a finite set of data points $\mathit{xs}$. This size-constrained ODT problem typically involves two hyperparameters. The trade-off between accuracy and tree complexity is managed either by:
  1. 1.
    Fixing the tree size $\left|T\right|$ (the number of branching nodes, with the number of leaves equal to $\left|T\right|+1$), or
  2. 2.
    Adding a penalty term $\lambda\left|T\right|$ to the objective function.
Both approaches are equivalent because a fixed $\lambda$ will corresponds to a tree with a fixed size $\left|T\right|$. In this exposition, we focus on the case where $\left|T\right|=K$. The second hyperparameter controls the minimum number of data items required in each leaf, denoted $N_{\text{min}}$. The size-constrained ODT problem can therefore be formulated as follows
$T^{*}=$  $\underset{T\in\mathcal{S}_{\text{size}}\left(K,\mathit{rs}\right)}{\text{ argmin }}E\left(T\right)$ (2)
s.t., $\left|l\right|\geq N_{\text{min}},\forall l\in\text{leaves}\left(T\right)$
where $E\left(T\right)$ is the objective function. In classification problems, it typically counts the number of misclassifications (i.e., the 0-1 loss), which will be the primary focus of this paper. The set $\mathcal{S}_{\text{size}}\left(K,\mathit{rs}\right)$ denotes the combinatorial search space of all decision trees of size $K$ with respect to a list of splitting rules $rs$, and $\text{leave}\left(T\right)$ represents all leaf nodes of the tree $T$.
Alternatively, one can define the ODT problem with a depth constraint $D$ as
$T^{*}=$  $\underset{T\in\mathcal{S}_{\text{depth}}\left(D,\mathit{xs}\right)}{\text{ argmin }}E\left(T\right)$ (3)
s.t., $\left|l\right|\geq N_{\text{min}},\forall l\in\text{leaves}\left(T\right)$
where $\mathcal{S}_{\text{depth}}\left(D,\mathit{xs}\right)$ denotes the search space of decision trees of depth $D$ with respect to the data list $\mathit{xs}$ .
It is important to recognize that these definitions do not explicitly define $T$ beyond stating that it is a binary tree. Different authors may have slightly different interpretations of a decision tree. Some define a decision tree purely in terms of its splitting rules, where predictions are determined solely by the rules, while others consider the predictions themselves as an essential part of the tree’s structure.
The primary difference between our study and previous work on combinatorial methods is that all previously undefined terms in (2) and (3)—such as leave, $E$, $T$, and $\mathcal{S}_{\text{size}}$ (or alternatively, $\mathcal{S}_{\text{depth}}\left(K,\mathit{xs}\right)$ for depth-constrained tree)—will be defined unambiguously using programs or datatypes with formal axioms.
Some astute readers may wonder why we define the search space of depth-constrained trees using the data list $\mathit{xs}$ and size-constrained trees using $\mathit{rs}$. This is an operational choice aimed at maintaining elegance in the algorithm’s expression, as we will explain in Subsection 4.2. Another reason for treating $\mathit{xs}$ and $\mathit{rs}$ separately is to avoid arbitrary choices of splitting rules, which could lead to constructing an algorithm that does not terminate. If the set of possible splitting rules were infinite, it would be impossible to design an algorithm that constructs an ODT in finite time.
Therefore, the first and most important step is to define unambiguously what we mean by a decision tree $T$. This step is often overlooked in prior literature using combinatorial methods. Without a formal definition of a “decision tree,” it is impossible to rigorously define any other terms related to it.
Proper decision tree
As mentioned, a decision tree is a binary tree that can be defined recursively using the following datatype
$\mathit{DTree}\left(\mathcal{A},\mathcal{B}\right)=DL\left(\mathcal{B}\right)\mid DN\left(\mathit{DTree}\left(\mathcal{A},\mathcal{B}\right),\mathcal{A},\mathit{DTree}\left(\mathcal{A},\mathcal{B}\right)\right).$
[…]
However, not every binary tree qualifies as a valid decision tree. To characterize decision trees precisely, the datatype must be accompanied by a set of axioms. We have previously formalized this notion by defining a proper decision tree, associating $\mathit{DTree}\left(\mathcal{A},\mathcal{B}\right)$with four axioms introduced in He and Little (2025b), as described below.
Definition 1.
[…]
We call a decision tree that satisfies the following axioms, a proper decision tree with respect to $rs$:
  1. 1.
    Each branch node is defined by a single splitting rule $r:\mathcal{R}$, and each splitting rule subdivides the ambient space into two disjoint and connected subspaces, $r^{+}$ and $r^{-}$. The descendant rules of $r$ can be generated only from within $r^{+}$ and $r^{-}$,
  2. 2.
    Each leaf $L$ is defined by the intersection of subspaces $\bigcap_{p\in P_{L}}r_{p}^{\pm}$ for all the splitting rules $\left\{r_{p}\mid p\in P_{L}\right\}$ in the path $P_{L}:\mathcal{P}\left(\mathcal{A}\right)$ from the root to leaf $L$. The connected region (subspace) defined by $\bigcap_{p\in P_{L}}r_{p}^{\pm}$ is referred to as the decision region,
  3. 3.
    The ancestry relation between any pair of splitting rules $r_{i}\left(\swarrow\vee\searrow\right)r_{j}$ is transitive; in other words, if $r_{i}\left(\swarrow\vee\searrow\right)r_{j}$ and $r_{j}\left(\swarrow\vee\searrow\right)r_{k}$ then $r_{i}\left(\swarrow\vee\searrow\right)r_{k}$,
  4. 4.
    For any pair of splitting rules $r_{i}$ and $r_{j}$, only one of the following three cases is true: $r_{i}\swarrow r_{j}$, $r_{i}\searrow r_{j}$, and $r_{i}\overline{\left(\swarrow\vee\searrow\right)}r_{j}$; additionally, $r_{i}\overline{\left(\swarrow\vee\searrow\right)}r_{i}$ is always true;.
Consider a tree with splitting rules $r$ defined as hyperplanes $h$, referred as the decision hyperplanes $h_{1}$, $h_{2}$, and $h_{3}$ defined by normal vector $\boldsymbol{w}_{1}$, $\boldsymbol{w}_{2}$, and $\boldsymbol{w}_{3}$, respectively. A tree of depth two can be illustrated as follows
$h_{1}$$h_{2}$$h_{1}^{+}\cap h_{2}^{+}$$\bar{\boldsymbol{w}}_{2}^{T}\boldsymbol{x}\geq 0$$h_{1}^{+}\cap h_{2}^{-}$$\bar{\boldsymbol{w}}_{2}^{T}\boldsymbol{x}<0$$\bar{\boldsymbol{w}}_{1}^{T}\boldsymbol{x}\geq 0$$h_{3}$$h_{1}^{-}\cap h_{3}^{+}$$\bar{\boldsymbol{w}}_{3}^{T}\boldsymbol{x}\geq 0$$h_{1}^{-}\cap h_{3}^{-}$$\bar{\boldsymbol{w}}_{3}^{T}\boldsymbol{x}<0$$\bar{\boldsymbol{w}}_{1}^{T}\boldsymbol{x}<0$
Each hyperplane $h_{i}$ partitions $\mathbb{R}^{D}$ into two regions: the positive region $h_{i}^{+}=\left\{x\mid x\in\mathbb{R}^{D},\boldsymbol{w}_{i}^{T}\bar{x}\geq 0\right\}$ and the negative region $h_{i}^{-}=\left\{x\in\mathbb{R}^{D},\boldsymbol{w}_{i}^{T}\bar{x}<0\right\}$, where $\bar{x}=\left(x,1\right)\in\mathbb{R}^{D+1}$ denotes the data points in homogeneous coordinates. For brevity, we omit the labels on the branch arrows $\boldsymbol{w}^{T}x<0$ and $\boldsymbol{w}^{T}x\geq 0$, assuming by default that the left branch corresponds to the “less than” case and the right branch to the “greater than or equal to” case.
In ML, the most commonly used decision rules are defined by axis-parallel hyperplanes, denoted by $\mathcal{H}^{0}$. In our discussion we will generalize the classical use of $\mathcal{H}^{0}$ to $M$-degree polynomial hypersurfaces $\mathcal{H}^{M}$ and mixed splitting rules $\mathcal{H}^{m}$, for all $0\leq m\leq M$.
The set $\mathcal{H}^{m}$ consists of arbitrary polynomial hypersurfaces defined as $h^{m}=\left\{x\mid\boldsymbol{w}^{T}\tilde{\boldsymbol{x}}<0\right\}$, where $\tilde{\boldsymbol{x}}=\rho_{m}\left(\bar{\boldsymbol{x}}\right)$ is the $m$-tuple embedding of $\bar{\boldsymbol{x}}$ (The embedding will be defined formally in Subsection 5.1). This definition holds for all $0\leq m\leq M$, implying that a decision tree can simultaneously incorporate axis-parallel hyperplane $\mathcal{H}^{0}$, general hyperplane $\mathcal{H}$, and hypersurfaces $\mathcal{H}^{M}$.
Ancestry relation matrix
Figure 1: The ancestry relation graph (left) captures all ancestry relations between four splitting rules $\left[r_{0},r_{1},r_{2},r_{3}\right]$. In this graph, nodes represent rules, and arrows represent ancestral relations. An incoming arrow from $r_{j}$ to a node $r_{i}$ indicates that $r_{j}$ is the right-child of $r_{i}$. The absence of an arrow indicates no ancestral relation. An outgoing arrows from $r_{i}$ to a node $r_{j}$ indicates that $r_{j}$ is the left-child of $r_{i}$. The ancestral relation matrix (right) $\boldsymbol{K}$, where the elements $\boldsymbol{K}_{ij}=1$, $\boldsymbol{K}_{ij}=-1$, and $\boldsymbol{K}_{ij}=0$ indicate that $r_{j}$ lies on the positive side, negative side of $r_{i}$, or that there is no ancestry relation between them, respectively.
[…]
Definition 2.
Ancestry relation matrix. Given a list of $K$ rules $\mathit{rs}=\left[r_{1},r_{2},\ldots r_{K}\right]$, the ancestry relations between any pair of rules can be characterized as a $K\times K$ square matrix $\boldsymbol{K}$, with elements defined as follows:
[…]
  • $\boldsymbol{K}_{ij}=0$ if $r_{i}\overline{\left(\swarrow\vee\searrow\right)}r_{j}$, where $\overline{R}$ represent the complement relation of $R$. According to De Morgan’s law $r_{i}\overline{\left(\swarrow\vee\searrow\right)}r_{j}=\left(r_{i}\overline{\swarrow}r_{j}\right)\wedge\left(r_{i}\overline{\searrow}r_{j}\right)$. In other words, $\boldsymbol{K}_{ij}=0$ if and only if $h_{j}$ is not a branch node in both the left and right subtree of $h_{i}$ and $\boldsymbol{K}_{ij}\neq 0$, $i\neq j$, if $r_{i}$ is the ancestor of $r_{j}$.
As we will see shortly, abstracting the ancestry relation using a matrix provides a compact and elegant representation for comparing different splitting rules and serves as a key component in constructing efficient algorithms for solving the size-constrained ODT problem.
4.2 Four definitions for optimal decision tree problems
After clearly defining what we mean by a “decision tree,” we now turn to constructing an unambiguous and intellectually manageable definition for optimal decision tree problem. In combinatorial optimization, the brute-force algorithm (also known as the exhaustive enumeration paradigm) is widely recognized as one of the most transparent approaches to solving combinatorial optimization problems. It offers a simple yet provably correct formalism: systematically enumerate all candidate solutions in the search space and select the optimal one.
In this section, we first introduce a general formalism for programmatically specifying brute-force algorithms and explain its significance. We then demonstrate how the ODT problem can be specified unambiguously using brute-force algorithms.
4.2.1 Defining combinatorial optimization problem formally through brute-force enumeration
In the theory of transformational programming (also known as constructive algorithmics) (Bird and De Moor, 1996), combinatorial optimization problems such as 1 is solved using the following, generic, generate-filter-select paradigm,
$s^{*}=\mathit{min}_{E}\left(\mathit{filter}_{p}\left(\mathit{gen}\left(\mathit{xs}\right)\right)\right)$ (4)
The combinatorial generator $\mathit{gen}:\left[\mathcal{A}\right]\to\left[S\right]$ generates a list of all solutions of type $S$ corresponding to elements in the search space $\mathcal{S}\left(\mathit{xs}\right)$ (parameterized by $\mathit{xs}$), with respect to a data list $\mathit{xs}:\left[\mathcal{A}\right]$. Infeasible solutions are then filtered out using the function $\mathit{filter}:\left(S\to\mathit{Bool}\right)\times\left[S\right]\to\left[S\right]$ by retaining only those solutions that satisfy a predicate $p:S\to\mathit{Bool}$. Finally the function $\mathit{min}_{E}$ select the optimal solution with respect to objective $E$. In other words, (4) is essentially a brute-force program.
Astute readers may observe that (4) differs from (1) only in form; they are not fundamentally different. In other words, specification (4) remains ambiguous, since we have not yet defined $\mathit{min}_{E}$, $\mathit{filter}_{p}$, and $\mathit{gen}$. The definitions of $\mathit{min}_{E}$ and $\mathit{filter}_{p}$ are relatively straightforward. We can define $\mathit{min}$ recursively with respect to an objective function $E$ as
$\mathit{min}_{E}$ $\left(\left[a\right]\right)=a$ (5)
$\mathit{min}_{E}$ $\left(a:\mathit{as}\right)=\mathit{smaller}_{E}\left(a,\mathit{min}_{E}\left(\mathit{as}\right)\right),$
where $\mathit{smaller}_{E_{\text{0-1}}}\left(a,b\right)=a$ if $E_{\text{0-1}}\left(a\right)\leq E_{\text{0-1}}\left(b\right)$ and $b$ otherwise. Similarly the $\mathit{filter}_{p}$ is defined recursively as
[…]
The definition of $\mathit{gen}$ will be the main focus of the remainder of this subsection. We provide four distinct ways of defining $\mathit{gen}$ each yielding a corresponding definition of the optimal decision tree problem. Before presenting the definitions of $\mathit{gen}$ in the context of ODT problems, we first consider the following three questions:
Why is (4) important for the study of optimal algorithms? How does (4) differ from MIP specifications or the ambiguous definitions used in BnB methods? Most importantly, how can we construct an efficient program that guarantees solutions aligned with Equation (4)?
Importance of brute-force specifications in deriving efficient solutions
The primary challenge in combinatorial optimization is to construct efficient, optimal solutions that satisfy the constrains of the original problem definition, such as that given in (4). A MIP specification in standard form can directly yield efficient solutions when using a general-purpose solver. However, a brute-force specification, as defined in (4) may initially seem computationally infeasible due to the intractable combinatorics of many problems. The deceptive simplicity of the brute-force algorithm can make it seem irrelevant to the construction of efficient combinatorial algorithms, such as dynamic programming, branch-and-bound (BnB), greedy methods, or divide-and-conquer (D&C). However, the relationship between brute-force specifications and efficient combinatorial optimization algorithms is far more nuanced than this deceptive appearance suggests.
Within the constructive programming community (Bird and De Moor, 1996; De Moor, 1994; Meertens, 1986), it has long been recognized that brute-force algorithms can serve as a foundation for deriving efficient optimization algorithms through constructive proofs. A constructive proof is a form of mathematical reasoning in which one explicitly demonstrates how to construct the object or solution being claimed to exist. In other words, it not only establishes existence but also provides a computational procedure, typically derived through equations between programs. In computer science, this process, known as program calculus, involves establishing equivalences between programs. The Bird-Meertens formalism (Bird, 1987; Meertens, 1986; Bird and De Moor, 1996) is a notable framework that systematically derives programs from their specifications through equational reasoning.
Using this formalism, the derived program inherently satisfies the properties of the original specification, ensuring both computational efficiency and formal correctness. Thus, for combinatorial optimization problems, a rigorous and systematic approach begins with a provably correct brute-force specification, from which efficient solutions are derived constructively. This method ensures that the resulting algorithms are both computationally efficient and formally verified, serving as the central formalism for algorithm derivation in this paper.
A natural question arises: Can efficient solutions (efficient combinatorial optimization algorithms) always be derived from brute-force specifications, and how comprehensive is this derivation paradigm across combinatorial methods? Bird and De Moor (1996) formalized classical combinatorial methods, such as dynamic programming, greedy algorithms, and D&C, using a generic recursive program (hylomorphism) and derived efficient solutions as theorems. Building on this, He (2025) extended the framework to include BnB by incorporating search strategies into Bird’s theorems, further broadening the applicability of this approach. In practice, one only needs to verify the conditions of these theorems, after which program fusion follows automatically.
Furthermore, He (2025) established the following inclusion relationships among different combinatorial optimization methods:
$\text{SDP}\subseteq\text{Greedy algorithm}\subseteq\text{BnB}\subseteq\text{General SDP}\subseteq\text{ DP},\text{Classical D\&C}\subseteq\text{\text{General D\&C}},$ (7)
Where SDP stands for “sequential decision process,” a special form of recursion characterized by sequential and deterministic recursive steps. In simple, the inclusion relations in (7) shows that a general definition of D&C algorithms encompasses all classical combinatorial optimization methods. Both DP and classical D&C correspond to problems with overlapping and non-overlapping subproblems, respectively, and both can be viewed more generally than the BnB method, which is a sequential decision process combined with additional techniques, such as thinning and search strategies. Greedy algorithms represent the simplest instance, corresponding to the fused program of an SDP generator with the selector $\mathit{min}$. Importantly, this inclusion is not based on intuitive classification but on a rigorous formal abstraction of each method using abstract recursion, and these inclusion relations naturally emerge from the abstract level of these recursions. See Section II.2.9 of He (2025) for precise definitions of these terms and explanations.
In order to construct an efficient combinatorial optimization algorithm using constructive approach, two key aspects must be considered in the design of the generator: (i) the efficiency of the generator itself, and (ii) its feasibility for program fusion, i.e., the compatibility between the structures of $\mathit{gen}$ and the objective function $E$.
A carefully designed generator whose structure aligns with both the algorithmic process and the underlying hardware/software (e.g., enabling vectorization) often yields more efficient solutions than algorithms with better theoretical complexity but poor hardware compatibility. Therefore, when designing $\mathit{gen}$, hardware compatibility and parallelizability should be considered from the outset, rather than designing an ad-hoc sequential algorithm and attempting to optimize and parallelize it post-hoc. Moreover, structural compatibility between the objective function $E$ and the generator $\mathit{gen}$ enables program fusion for $\mathit{min}_{E}$ and $\mathit{gen}$. Such fusion produces programs that are not only more succinct but also more efficient, as infeasible or suboptimal configurations can be pruned before being extended into complete solutions.
Comparison of brute-force and mixed integer programming Specifications
Compared to a MIP specification, such as that presented in Bertsimas and Dunn (2017), a brute-force specification defined programmatically (rather than through pseudo-code) can be equally or more rigorous. Programmatic definitions often provide greater clarity in computational contexts because the minimization function in MIP specifications may lack precise definition due to its reliance on set theory. Set theory, while mathematically robust, is less suitable for computer science applications, as algorithms require deterministic procedures that account for element ordering, which set theory typically disregards.
In contrast to BnB algorithms, where the search space is often ambiguously described using symbolic notation, a brute-force specification can define the search space $\mathcal{S}\left(\mathit{xs}\right)$ clearly using a combinatorial generator $\mathit{gen}\left(\mathit{xs}\right)$. This allows researchers to verify whether the search space aligns with the problem’s requirements. If the search space $\mathcal{S}\left(\mathit{xs}\right)$ for an ODT problem is not explicitly defined, or if an algorithm is proposed without a clear problem formulation, verifying its correctness becomes impossible. In the following subsection, we present three distinct definitions for $\mathcal{S}_{\text{size}}\left(K,\mathit{rs}\right)$ (size-constrained ODT) and one definition for $\mathcal{S}_{\text{depth}}\left(d,\mathit{xs}\right)$ (depth-constrained ODT). These definitions serve as the foundation for future research on ODT problems. They not only resolve ambiguities of problem definitions in previous studies that relied on the BnB method but also enable researchers to approach ODT problems from different perspectives.
4.2.2 Defining optimal decision tree problem through brute-force specifications
Given a list of rule $\mathit{rs}:\left[\mathcal{R}\right]$, and the size constraint $K:\mathbb{N}$, if we can define a combinatorial generator $\mathit{genDTSs}\left(K,\mathit{rs}\right)$ (short for “generate decision trees with size constraints”) that exhaustively enumerates all possible decision trees with exactly $K$ leaves with respect to $\mathit{rs}$, we can define the size-constrained decision tree problem as
$\mathit{odk}_{\text{size}}\left(K\right)$ $:\left[\mathcal{R}\right]\to\mathit{DTree}\left(\mathcal{R},\mathcal{D}\right)$ (8)
$\mathit{odk}_{\text{size}}\left(K\right)$ $=\mathit{min}_{E}\circ\mathit{genDTSs}\left(K\right)$ (9)
Since there are three possible definitions for $\mathit{odk}_{\text{size}}$, we replace its subscript with the name of the construction method. For example, $\mathit{odk}_{\text{x}}$ denotes the “x method for optimal decision tree algorithm with size constraints”. Similarly, the depth-constrained tree is defined as
$\mathit{odk}_{\text{depth}}\left(K\right)$ $:\mathcal{D}\to\mathit{DTree}\left(\mathcal{R},\mathcal{D}\right)$ (10)
$\mathit{odk}_{\text{depth}}\left(K\right)$ $=\mathit{min}_{E}\circ\mathit{genDTDs}\left(K\right)$ (11)
where $\mathit{genDTSDs}:\mathbb{N}\times\mathcal{D}\to\left[\mathit{DTree}\left(\mathcal{R},\mathcal{D}\right)\right]$ (short for “generate decision trees with depth constraints”) exhaustively enumerates all possible decision trees of a given depth. We note that $\mathit{odk}_{\text{depth}}$ and $\mathit{odk}_{\text{size}}$ take different types of input for the sake of program elegance. The input $\mathit{rs}:\left[\mathcal{R}\right]$ of $\mathit{odk}_{\text{size}}$ can be generated from data, which will be discussed in Subsection 5.2.
For a decision tree satisfying Axiom 1, He and Little (2025b) demonstrated the following important theorem, which reveals a deep connection between the combinatorics of decision trees and permutations.
Theorem 1.
A decision tree consisting of $K$ splitting rules corresponds to a unique $K$-permutation if and only if it is proper. In other words, there exists an injective mapping from proper decision trees to valid $K$-permutation (i.e., $K$-permutations that satisfies the proper decision tree axiom).
Proof.
See He and Little (2025b) for more details. ∎
As a consequence, the following lemma is immediate.
Lemma 1.
The search space of decision trees with $K$ splitting rule $\mathcal{S}_{\text{size}}\left(K,\mathit{rs}\right)$ with respect to a list of rules $rs$ is included in $\mathcal{S}_{\text{kperms}}\left(K,\mathit{rs}\right)$, the space of $K$-permutations with respect to $rs$. In other words, we have
$\mathcal{S}_{\text{size}}\left(K,\mathit{rs}\right)\subseteq\mathcal{S}_{\text{kperms}}\left(K,\mathit{rs}\right)$ (12)
Moreover, it can be shown that a decision tree attains maximal combinatorial complexity, i.e., $\mathcal{S}_{\text{DTree}}\left(K,\mathit{rs}\right)=\mathcal{S}_{\text{kperms}}\left(K,\mathit{rs}\right)$, when any rule can serve as the root and each branch node has exactly one child He and Little (2025b).
Since permutations are among the most extensively studied combinatorial structures, this characterization provides a fundamental basis for analyzing the combinatorial and algorithmic properties of decision trees. He and Little (2025a) present an elegant and efficient $K$-permutation generator expressed in a divide-and-conquer form. For simplicity, we denote this generator as $\mathit{kperms}\left(K,rs\right)$.
As a result, we can formally define the search space $\mathcal{S}_{\text{size}}\left(K,\mathit{rs}\right)$ using a combinatorial generator $\mathit{genDTSs}_{\text{kperms}}$ (short for “generate size-constrained decision trees based on $K$-permutation.”), defined as
[…]
$\mathit{genDTSs}_{\text{kperms}}\left(K\right)$ $=\mathit{filter}_{p}\circ\mathit{kperms}_{K}.$
The program first generates all possible $K$-permutations using $\mathit{kperms}_{K}$, and then filters out, via $\mathit{filter}_{p}$, those permutations that cannot be used to construct proper decision trees. This two-step process ensures that only valid permutations—i.e., those $K$-permutations of splitting rules satisfying the decision tree axiom 4.2, and meeting the structural and combinatorial requirements of proper decision trees—are retained for further computations.
Therefore, the following specification for the optimal decision tree problem follows immediately:
$\mathit{odk}_{\text{size}}\left(K\right)$ :[][$\left$$\mathcal{R}\right$$\to\left$$\mathcal{R}\right$]
$\mathit{odk}_{\text{size}}\left(K\right)$ $=\mathit{min}_{E_{\text{0-1}}}\circ\mathit{genDTSs}_{\text{kperms}}\left(K\right)$
where $\mathit{odk}_{\text{kperms}}\left(K,\mathit{rs}\right)$ select the optimal $K$-permutation from all feasible $K$-permutations generated by $\mathit{genDTKs}_{\text{kperms}}\left(K,\mathit{rs}\right)$.
Definition of the generator based on tree datatype for size-constrained tree
The definition based on $K$-permutations remains unsatisfactory. The number of proper decision trees for a given set of rules is typically much smaller than the total number of permutations, and determining whether a permutation is feasible is often non-trivial. In the hyperplane decision tree problem, for example, the feasibility test requires $O\left(K^{2}D^{2}\right)$ operations in the worst-case, where $D$ is the dimension of the data. As predicted by He and Little (2025b), if each hyperplane classifies a data point into the positive or negative class with equal probability (i.e., 1/2 for each class), the probability of a decision tree achieving maximal combinatorial complexity is only $O\left(\left(\frac{1}{2}\right)^{D\times K^{2}}\right)$.
Nevertheless, the following observation suggests a more efficient approach. Consider the search space $\mathcal{S}_{\text{size}}\left(K,\mathit{rs}\right)$ for $\left|\mathit{rs}\right|\geq K$. The following equivalence holds
$\mathcal{S}_{\text{size}}\left(K,\mathit{rs}\right)=\bigcup_{\mathit{rs}_{K}\in\mathcal{S}_{\text{kcombs}}\left(K,\mathit{rs}\right)}\mathcal{S}_{\text{size}}\left(K,\mathit{rs}_{K}\right),$ (14)
where $\mathit{rs}_{K}=\left[r_{1},r_{2},\ldots r_{K}\right]$ denote a list of $K$ rules. By assuming that $\mathcal{S}_{\text{size}}\left(K,\mathit{rs}_{K}\right)$ can be generated from $\mathit{genDTs}\left(\mathit{rs}_{K}\right)$ (short for “generate decision trees with exactly $K$ splitting rules”, where the parameter $K$ is implicit, since $\left|\mathit{rs}_{K}\right|=K$) where $\mathit{genDTs}\left(\mathit{rs}_{K}\right)=\mathit{genDTSs}_{\text{kperms}}\left(K,\mathit{rs}_{K}\right)$ is one possible definition, we can define 14 programmatically
$\mathit{genDTSs}\left(K\right)=\mathit{concatMapL}_{\mathit{genDTs}}\circ\mathit{kcombs}\left(K\right),$ (15)
Interestingly, He and Little (2025b) showed that $\mathcal{S}_{\text{size}}\left(K,\mathit{rs}_{K}\right)$ can be generated more efficiently by using a function $\mathit{genDTs}_{\text{rec}}$ (short for “recursive method for generating decision tree with exactly $K$ splitting rules”) than with $\mathit{genDTSs}_{\text{kperms}}\left(K,\mathit{rs}_{K}\right)$. The recursive generator $\mathit{genDTs}_{\text{rec}}:\left[\mathcal{R}\right]\times\mathcal{D}\to\left[\mathit{DTree}\left(\mathcal{R},\mathcal{D}\right)\right]$ is based on the tree datatype $\mathit{DTree}\left(\mathcal{R},\mathcal{D}\right)$, which is defined as
$\mathit{genDTs}_{\text{rec}}$ $:\mathcal{D}\times\left[\mathcal{R}\right]\to\left[\mathit{DTree}\left(\mathcal{R},\mathcal{D}\right)\right]$ (16)
$\mathit{genDTs}_{\text{rec}}$ $\left(\mathit{xs},\left[\;\right]\right)=\left[\mathit{DL}\left(\mathit{xs}\right)\right]$
$\mathit{genDTs}_{\text{rec}}$ $\left(\mathit{xs},\left[r\right]\right)=\left[DN\left(\mathit{DL}\left(\mathit{xs}^{+}\right),r,\mathit{DL}\left(\mathit{xs}^{-}\right)\right)\right]$
[…]
$\begin{aligned} &\quad\quad\quad\left(\mathit{rs}^{+},r_{i},\mathit{rs}^{-}\right)\leftarrow\mathit{splits}\left(\mathit{rs}\right),u\leftarrow\mathit{genDTs}\left(\mathit{xs},\mathit{rs}^{+}\right),v\leftarrow\mathit{genDTs}\left(\mathit{xs},\mathit{rs}^{-}\right)\bigg{]}.\end{aligned}$
where $splits\left(\mathit{rs}\right)=\left[\left(\mathit{rs}^{+},r_{i},\mathit{rs}^{-}\right)\mid r_{i}\leftarrow rs,\mathit{all}\left(r_{i},\mathit{rs}\right)=\mathit{True}\right]$ ($\mathit{all}\left(r_{i},rs\right)$ returns true if all rules $r_{j}$ in $rs$ satisfy $K_{ij}\neq 0$ for $i\neq j$, and false otherwise), $\mathit{rs}^{\pm}=\left[r_{j}\mid r_{j}\leftarrow rs,\boldsymbol{K}_{ij}=\pm 1\right]$, $\mathit{xs}\cap\mathit{ys}=\left[x\mid x\leftarrow\mathit{xs},x\in\mathit{ys}\right]$ (all elements that both in list $\mathit{xs}$ and $\mathit{ys}$), and $\mathit{xs}^{\pm}=\mathit{xs}\cap r^{\pm}=\left[x\mid x\leftarrow\mathit{xs},x\in r^{\pm}\right]$ ($\mathit{xs}^{+}$ collect all data points lies in the space defined by $r^{+}$) . Assume that the ancestry relation matrix $\boldsymbol{K}$ can be pre-stored in memory (as shown for hypersurface splitting rules in Part II). Under this assumption, the $\mathit{splits}$ function requires only $O\left(K\right)$ operations.
The $\mathit{genDTs}$ generator function recursively constructs larger proper decision trees $N\left(u,r_{i},v\right)$ from smaller proper decision trees $\mathit{genDTs}\left(\mathit{rs}^{+}\right)$ and $\mathit{genDTs}\left(\mathit{rs}^{-}\right)$, with the $\mathit{splits}$ function ensuring that only feasible splitting rules can become subtree roots during recursion. At each step, $\mathit{genDTs}$ accumulates information by creating a root $r_{i}$ for every proper decision subtree generated by $\mathit{genDTs}\left(rs^{+},xs\right)$, using the $\mathit{mapD}{}_{\cap_{r_{i}^{+}}}$ function. The $\mathit{genDTs}$ function will generate $K!$ number of tree in the worst-case, the worst-case is achieved is precisely when the tree has a “chain-like” structure (every nodes of the tree has at most one child), in this case the proper decision tree established a one-to-one correspondence to the $K$-permutation.
This above observation leads to the following program for generating $\mathcal{S}_{\text{size}}\left(K,\mathit{rs}\right)$
$\mathit{genDTSs}_{\text{rec}}\left(K,\mathit{rs}\right)=\mathit{concatMapL}_{\mathit{genDTs}_{\text{rec}}\left(\mathit{xs}\right)}\left(\mathit{kcombs}\left(K,\mathit{rs}\right)\right),$ (17)
such that $\left|\mathit{rs}\right|\geq K$, where $\mathit{genDTSs}_{\text{rec}}\left(K\right)$ are short for “recursive method for generating decision trees of $K$ splitting rules.”
After formalizing the search space $\mathcal{S}_{\text{size}}\left(K,\mathit{rs}\right)$ by defining it as a concrete program (17), now we can formally define what we mean by optimal decision problem by the following brute-force algorithm
$\mathit{odt}_{\text{rec}}\left(K\right)$ $:\left[\mathcal{R}\right]\to\mathit{DTree}\left(\mathcal{R},\mathcal{D}\right)$ (18)
$\mathit{odt}_{\text{rec}}\left(K\right)$ $=\mathit{min}_{E_{\text{0-1}}}\circ\mathit{genDTKs}_{K}.$
$=\mathit{min}_{E_{\text{0-1}}}\circ\mathit{concatMapL}{}_{\mathit{genDTs}_{\text{rec}}\left(\mathit{xs}\right)}\circ\mathit{kcombs}\left(K\right).$
Alternatively, if the definition $\mathit{genDTs}$ is defined by $\mathit{genDTSs}_{\text{kperms}}\left(K,\mathit{rs}_{K}\right)$, we have
$\mathit{odt}_{\text{kperms}}\left(K\right)$ $=\mathit{min}_{E_{\text{0-1}}}\circ\mathit{concatMapL}{}_{\mathit{genDTSs}_{\text{kperms}}\left(K\right)}\circ\mathit{kcombs}_{K}$
$=\mathit{min}_{E_{\text{0-1}}}\circ\mathit{concatMapL}{}_{\mathit{filter}_{p}\circ\mathit{kperms}_{K}}\circ\mathit{kcombs}_{K}.$
A sequential definition based on tree datatype for size-constrained tree
For a size-constrained decision tree, the number of leaves is fixed, so we know that there are at most $K$ recursive steps to construct a full decision tree with $K$-fixed splitting rules. This provides us an iterative method of defining the decision tree generator
$\mathit{genDTs}_{\text{vec}}$ $:\mathcal{D}\times\left[\mathcal{R}\right]\to\left[\mathit{DTree}\left(\mathcal{R},\mathcal{D}\right)\right]$ (19)
$\mathit{genDTs}_{\text{vec}}$ $\left(\mathit{xs},\left[\;\right]\right)=\left[\mathit{DL}\left(\mathit{xs}\right)\right]$
$\mathit{genDTs}_{\text{vec}}$ $\left(\mathit{xs},\mathit{rs}\right)=\mathit{concat}\circ\left[\mathit{updates}\left(r,\mathit{rs}^{\prime},xs\right)|\left(r,\mathit{rs}^{\prime}\right)\longleftarrow\mathit{candidates}\left(\mathit{rs}\right)\right]$
[…]
$\mathit{updates}$ $:\mathcal{R}\times\left[\mathcal{R}\right]\to\left[\mathit{DTree}\left(\mathcal{R},\mathcal{D}\right)\right]$ (20)
$\mathit{updates}$ $\left(r,\mathit{rs}\backslash r,\mathit{xs}\right)=\mathit{catMaybes}\left[\mathit{update}\left(r,t\right)\mid t\leftarrow\mathit{genDTs}_{\text{vec}}\left(\mathit{rs}\backslash r,\mathit{xs}\right)\right]$
which call $\mathit{genDTs}_{\text{vec}}$ recursively, and append a new rule $r$ to every tree $t$ generated by $\mathit{genDTs}_{\text{vec}}\left(\mathit{rs}\backslash r,xs\right)$ by using $\mathit{update}$ function
$\mathit{update}$ $:\mathcal{R}\to\mathit{DTree}\left(\mathcal{R},\mathcal{D}\right)\to\mathit{Maybe}\left(\mathit{DTree}\left(\mathcal{R},\mathcal{D}\right)\right)$
$\mathit{update}$ $\left(r,\mathit{DL}\left(\mathit{xs}\right)\right)=\mathit{Just}\left(DN\left(\mathit{DL}\left(xs^{+}\right),r,\mathit{DL}\left(xs^{-}\right)\right)\right)$
[…]
Although the definition of e $\mathit{update}$ update appears complex, it is applied recursively along the path of the old tree $\mathit{DN}\left(u,s,v\right)$ to determine whether adding a new splitting rule $r$ would result in a non-proper tree. If the tree remains proper, rris added to the leaf; otherwise, $\mathit{Nothing}$ is returned, indicating a non-proper decision tree.
Because the computation in $\mathit{genDTs}_{\text{vec}}$ is sequential, it can be fully vectorized. The generator $\mathit{genDTs}_{\text{vec}}$ (short for “vectorized method generating decision trees with $K$ splitting rules.”) benefits particularly from this property, as it allows full utilization of modern hardware, such as GPUs. We discuss this advantage in detail in Subsection 4.5.
One can check $\mathit{genDTs}_{\text{vec}}$ will generate the same set of proper decision trees as $\mathit{genDTs}_{\text{rec}}$, but will produce different list of trees. This is because different valid $K$-permutation can correspondent to the same tree (vice is not true). For instance, the tree
$r_{1}$$r_{2}$$r_{3}$
can be constructed by sequentially traversing list $\left[r_{1},r_{2},r_{3}\right]$ or $\left[r_{1},r_{3},r_{2}\right]$. To remove duplicate trees, we need an additional $\mathit{unique}:\left[\mathit{DTree}\left(\mathcal{R},\mathcal{D}\right)\right]\to\left[\mathit{DTree}\left(\mathcal{R},\mathcal{D}\right)\right]$ operation which removes all duplicate trees in a list. This gives us
$\mathit{genDTs}_{\text{vec}}$ $:\mathcal{D}\times\left[\mathcal{R}\right]\to\left[\mathit{DTree}\left(\mathcal{R},\mathcal{D}\right)\right]$ (21)
$\mathit{genDTs}_{\text{vec}}$ $\left(\mathit{xs},\left[\;\right]\right)=\left[\mathit{DL}\left(\mathit{xs}\right)\right]$
$\mathit{genDTs}_{\text{vec}}$ $\left(\mathit{xs},\mathit{rs}\right)=\mathit{unique}\circ\mathit{concat}\circ\left[\mathit{updates}\left(r,\mathit{rs}^{\prime},xs\right)|\left(r,\mathit{rs}^{\prime}\right)\longleftarrow\mathit{candidates}\left(\mathit{rs}\right)\right]$
Now this new $\mathit{genDTs}_{\text{vec}}$ will produce the same trees as $\mathit{genDTs}_{\text{rec}}$, albeit in different ordering.
Finally, a sequential definition (a brute-force program) for the size-constrained optimal decision tree problem is defined as
$\mathit{odt}_{\text{vec}}\left(K\right)=\mathit{min}_{E_{\text{0-1}}}\circ\mathit{concatMapL}{}_{\mathit{genDTs}_{\text{vec}}\left(\mathit{xs}\right)}\circ\mathit{kcombs}_{K}.$ (22)
Definition based on tree datatype for depth-constrained tree
Depth-constrained decision trees are often used in the study of ODT problems with smaller combinatorial complexity, such as axis-parallel decision trees (Mazumder et al., 2022; Brița et al., 2025) or decision trees over binary feature data Demirović et al. (2022); Verwer and Zhang (2019); Nijssen and Fromont (2007); Aglin et al. (2020); Nijssen and Fromont (2010). However, none of these studies formally define their problem or explain why their algorithm correctly solves the desired problem. In this subsection, we provide a formal characterization of decision trees defined in this manner.
Moreover, while depth-constrained trees are widely used in previous studies on axis-parallel and binary-data decision trees, they are unsuitable for ODT problems with more complex splitting rules, such as hyperplanes or hypersurfaces. We explain this limitation from an algorithmic perspective in Subsection 4.5.1.
A depth-constrained tree generator $\mathit{genDTDs}_{\text{depth}}$ receives a pre-specified tree depth $d:\mathbb{N}$ and a list of data $\mathit{xs}$ as input, and is defined as
$\mathit{genDTDs}_{\text{depth}}$ $:\mathbb{N}\times\mathcal{D}\to\left[\mathit{DTree}\left(\mathcal{R},\mathcal{D}\right)\right]$ (23)
$\mathit{genDTDs}_{\text{depth}}$ $\left(0,\mathit{xs}\right)=\left[\mathit{DL}\left(\mathit{xs}\right)\right]$
$\mathit{genDTDs}_{\text{depth}}$ $\left(1,\mathit{xs}\right)=\left[DN\left(\mathit{DL}\left(\mathit{xs}^{+}\right),r,\mathit{DL}\left(\mathit{xs}^{-}\right)\right)\mid r\leftarrow\mathit{gen_{splits}}\left(\mathit{xs}\right)\right]$
$\mathit{genDTDs}_{\text{depth}}$ $\left(d,\mathit{xs}\right)=\left[\mathit{DN}\left(u,r,v\right)\mid u\leftarrow\mathit{genDTDs}_{\text{depth}}\left(\mathit{xs}^{+},d-1\right),v\leftarrow\mathit{genDTDs}_{\text{depth}}\left(\mathit{xs}^{-},d-1\right),r\leftarrow\mathit{gen_{splits}}\left(\mathit{xs}\right)\right]$
This definition states that all possible decision trees of depth $d$ with respect to data list $\mathit{xs}$ can be constructed from decision trees of depth d1d-1 on the subsets $\mathit{xs}^{+}$ and $\mathit{xs}^{-}$. Currently, a concrete definition for $\mathit{gen_{splits}}$ is not provided, as the splitting rule $r$ defining the decision tree problem has not yet been formalized. For the classical axis-parallel ODT problem, $\mathit{gen_{splits}}$ is straightforward: splitting rules can be generated from axis-parallel hyperplanes based on the data points. Equivalently, Brița et al. (2025); Mazumder et al. (2022) use the medians of adjacent pairs of data points for each dimension $D$.
For a data list of size $N$, there are $N\times D$ possible splitting rules. Each rule $r$ corresponds to an integer value defined by the $d$-th feature of the data points $x\in\mathbb{R}^{D}$, with $r^{+}=\left\{x\mid x_{d}>r\right\}$ and $r^{-}=\left\{x\mid x_{d}\leq r\right\}$. We will define $\mathit{gen_{splits}}$ for hypersurface decision trees in Subsection 5.2, after discussing the geometry of hypersurface splitting rules.
Therefore, the optimal decision tree problem with depth constraint $d:\mathbb{N}$ is defined as
$\mathit{odt}_{\text{depth}}$ $:\mathbb{N}\times\mathcal{D}\to\mathit{DTree}\left(\mathcal{R},\mathcal{D}\right)$ (24)
$\mathit{odt}_{\text{depth}}$ $\left(d\right)=\mathit{min}{}_{E}\circ\mathit{genDTDs}_{\text{depth}}\left(d\right)$
Some readers may find the recursion 23 closely related to the algorithm given by Brița et al. (2025), which has similar recursive structure. We will see shortly that Brița et al. (2025)’s algorithm can indeed be derived from the specification (24), which proves why Brița et al. (2025)’s algorithm is optimal.
4.2.3 Remarks of potential flaws for optimal decision tree problems over binary feature data
The recursions 23 and 18 are valid because, according to Axiom 1 of decision trees, subtrees can only be constructed from the subregions defined by $r$, which partition the data list $\mathit{xs}$ to $\mathit{xs}^{+}$ and $\mathit{xs}^{-}$. This allows us to fix a root $r$, and safely divide the remaining splitting rules into two disjoint lists of splitting rules, they correspond to $\mathit{gen_{splits}}\left(\mathit{xs}^{\pm}\right)$ in $\mathit{odt}_{\text{depth}}$ and $r^{\pm}$ in $\mathit{odt}_{\text{rec}}$. In particular, we do not need to worry about elements of $\mathit{gen_{splits}}\left(\mathit{xs}^{\pm}\right)$ or $r^{\pm}$ appearing in $\mathit{gen_{splits}}\left(\mathit{xs}^{\mp}\right)$ or $r^{\mp}$ because of Axiom 4, which ensures that only one of the three possible cases occurs.
However, in the decision tree problem over binary feature data, a selected feature may appear in either the left or right subtree, depending on the presence or absence of the feature, regardless of the root $r$. Consequently, for any chosen root $r\in\mathit{gen_{splits}}\left(\mathit{xs}\right)$ (or $r\in\mathit{rs}$ for $\mathit{odt}_{\text{rec}}$) any subset of $\r\mathit{gen_{splits}}\left(\mathit{xs}\right)\backslash$ r or $\r\mathit{rs}\backslash$ r can form a feasible subtree. This substantially increases the complexity of splitting in binary-feature decision trees. It is therefore unclear why existing studies on ODT over binary feature data—such as Zhang et al. (2023); Demirović et al. (2022); Aglin et al. (2021); Lin et al. (2020); Hu et al. (2019); Verwer and Zhang (2019); Nijssen and Fromont (2010, 2007)—define splitting rules in ways that are either ambiguous or entirely unspecified.
For instance, Demirović et al. (2022) provide a relatively clear definition using a recursive formula (Equation (1) in their paper). Their method proceeds by sequentially scanning the features (denoted as $i\in\left[0,n-1\right]$ in Equation (1), where $n$ is the number of features) and partitioning the remaining elements by selecting a data point with index $i$. Here is our hypothesis: for each fix index $i$, Demirović et al. (2022) partition the index list $\left[0,n-1\right]$ into subsets of indices smaller than $i$ (size $i$) and greater than $i$ (size ni1n-i-1), reminiscent of Catalan-style splitting. However, the trees constructed over binary feature data do not satisfy the standard decision tree axioms: a descendant feature $j$ can simultaneously appear as both the left and right child of its ancestor $i$, and fixing a root $i$ does not uniquely determine how the remaining n1n-1 features are split. The only plausible explanation for their algorithmic process is that they implicitly fix an ordering of features by arranging them as a list $\left[0,n-1\right]$, which explains their use of list notation rather than set notation. In this sense, a fixed-order labeled set (i.e., a list) is equivalent to an unlabeled set, which justifies our description of their recursion as “Catalan-style splitting,” originally designed for unlabeled trees. Moreover, line 9 of Algorithm 2 in their paper explicitly uses a for-loop to scan all features, which further supports our interpretation. Therefore, unless Demirović et al. (2022) can demonstrate that such splitting rules exhaustively enumerate all possible labeled decision trees over binary feature data, their claim of optimality remains, at best, premature.
Also, Demirović et al. (2022) state that their algorithms can be considered starting points for their work, which may suggest that the same splitting technique was employed. Verwer and Zhang (2019) also adopt a splitting principle similar to that of Demirović et al. (2022).
Hu et al. (2019) define leaves as either unsplittable ($d_{\text{un}}$) or splittable ($d_{\text{split}}$), but the criteria for assigning leaves to $d_{\text{un}}$ remain unclear. Their recursive process first scans all possible combinations of $d_{\text{split}}$ and then scans subsets of these combinations. It is unclear why this two-step approach is necessary, as considering all subsets directly would be equivalent. Moreover, Hu et al. (2019) present a highly complex formula (Equation (1) in their paper) for counting the combinatorics of the decision tree problem over binary feature data, yet they provide neither explanation nor justification. Consequently, it remains unclear why the recursive process in their algorithm can exhaustively enumerate the recursion it describes. In their subsequent study, Lin et al. (2020) modify the splitting function to resemble the approach of Aglin et al. (2021); Nijssen and Fromont (2007, 2010) scanning all possible splitting rules sequentially and partitioning the remaining data using an undefined $\mathit{split}$ function (line 13-14 in Algorithm 1 GOSDT).
Without formal problem definitions and rigorous proofs of algorithmic optimality, these studies on ODT over binary feature data may be incorrect. Even if their final algorithms or implementations produce correct results, ambiguities in definitions can mislead newcomers and impede progress in the field. Therefore, it is highly desirable to provide formal proofs demonstrating the correctness—or incorrectness—of these algorithms. Unfortunately, this is currently impossible due to ambiguous problem definitions in these works
4.3 Four solutions for optimal decision tree problems
As mentioned earlier, the efficiency of an algorithm for solving a problem is largely determined by the design of its generator, which implicitly defines the problem. In this section, we show that the four definitions introduced in 4.2 each lead to an efficient algorithm for solving the ODT problem.
4.3.1 Optimal decision tree algorithms for size-constrained tree
Simplified decision tree problem
The ODT problem defined in (18) is inherently difficult to solve directly. Fortunately, He and Little (2025b) demonstrated that the original ODT problem can be reduced to solving a simplified optimal decision tree problem, without loss of generality. This reduction not only makes the resulting algorithm significantly more efficient but also provides greater flexibility for designing parallel algorithms.
Theorem 2.
Simplified decision tree problem. Given a list of rules $\mathit{rs}:\left[\mathcal{R}\right]$ and a size constraint $K:\mathbb{N}$, we have following equivalence
$\mathit{min}_{E}\circ\mathit{concatMapL}{}_{\mathit{sodt}\left(\mathit{xs}\right)}\circ\mathit{kcombs}_{K}=\mathit{min}_{E}\circ\mathit{concatMapL}{}_{genDTs\left(\mathit{xs}\right)}\circ\mathit{kcombs}_{K}$ (25)
where $\mathit{sodt}\left(\mathit{xs}\right)=\mathit{min}_{E_{\text{0-1}}}\circ\mathit{genDTs}\left(\mathit{xs}\right)$.
Proof.
The result is a simple consequence of distributivity
$\mathit{min}{}_{E}\circ\mathit{mapL}{}_{genDTs\left(\mathit{xs}\right)}\circ\mathit{kcombs}{}_{K}$
$\equiv$ distributivity $\mathit{min}_{E}\circ\mathit{concat}=\mathit{min}_{E}\circ\mathit{mapL}_{\mathit{min}_{E}}$
$\mathit{min}{}_{E}\circ\mathit{mapL}{}_{\mathit{min}_{E}\circ genDTs\left(\mathit{xs}\right)}\circ\mathit{kcombs}{}_{K}$
$\equiv$ define $\mathit{sodt}\left(\mathit{xs}\right)=\mathit{min}_{E}\circ genDTs\left(\mathit{xs}\right)$
$\mathit{min}_{E}\circ\mathit{mapL}_{\mathit{sodt}\left(\mathit{xs}\right)}\circ\mathit{kcombs}{}_{K}$.
[…]
The left-hand formulation in (25) is computationally more efficient than the right-hand formulation because, for each $K$-combination of $rs$, only a single optimal decision tree is produced by $\mathit{sodt}$. Consequently, the $\mathit{min}{}_{E}$ function on the left-hand side of (25) needs to select the optimal decision tree only from the set generated by $\mathit{mapL}_{\mathit{sodt}\left(\mathit{xs}\right)}\circ\mathit{kcombs}{}_{K}$, which is significantly smaller than the set of all possible decision trees of size $K$.
Obviously, the three size-constrained generators—based on $K$-permutations, based on recursive tree structure, and sequential process, as discussed previously—yield three distinct definitions for $\mathit{sodt}$. For instance, the $K$-permutation generator can be factorized into the following two programs:
$\mathit{kperms}_{K}=\mathit{concatMap}_{\mathit{perms}}\circ\mathit{kcombs}_{K}$ (26)
where $\mathit{perms}:\left[\mathcal{A}\right]\to\left[\left[\mathcal{A}\right]\right]$ receives a list and returns all its permutations. In other words, the set of all possible $K$-permutations is equivalent to the set obtained by first generating all $K$-combinations and then computing the permutations of each combination.
In particular, $\mathit{perms}$ can be defined recursively as
[…]
$\mathit{perms}$ $\left(r:\mathit{rs}\right)=\mathit{ins}\left(r,\mathit{perms}\left(\mathit{rs}\right)\right)$
where function $\mathit{ins}$ insert $r$ into all possible positions for each partial permutations in $\mathit{perms}\left(\mathit{rs}\right)$ ( see subsection II.1.2.4 in He (2025) for details). Comparing (26) with (25), we can define $\mathit{sodt}$ using $\mathit{perms}$ through the following equivalence
$\mathit{sodt}_{\text{kperms}}$ :[][$\left$$\mathcal{R}\right$$\to\left$$\mathcal{R}\right$] (27)
$\mathit{sodt}_{\text{kperms}}$ $=\mathit{min}_{E}\circ\mathit{genDTs}_{\text{kperms}}$
$=\mathit{min}_{E}\circ\mathit{filter}_{p}\circ\mathit{kperms}_{K}$
$=\mathit{min}_{E}\circ\mathit{filter}_{p}\circ\mathit{perms}$
Similarly, by substituting $\mathit{genDTs}_{\text{rec}}$ and $\mathit{genDTs}_{\text{vec}}$ into the definition of $\mathit{sodt}$, we obtain the corresponding recursive and vectorized formulations of $\mathit{sodt}$:
$\mathit{sodt}_{\text{rec}}\left(\mathit{xs}\right)$ $=\mathit{min}_{E}\circ\mathit{genDTs}_{\text{rec}}\left(\mathit{xs}\right)$ (28)
$\mathit{sodt}_{\text{vec}}\left(\mathit{xs}\right)$ $=\mathit{min}_{E}\circ\mathit{genDTs}_{\text{vec}}\left(\mathit{xs}\right)$ (29)
Structure of objective, monotonicity, and dynamic programming
Once we have a formal definition, it is natural to ask whether an efficient solution—such as dynamic programming (DP) or a greedy algorithm—can be derived for solving $\mathit{sodt}$. It has long been recognized that the existence of a DP or greedy solution relies on identifying monotonicity (Bird and De Moor, 1996). Specifically, when the objective function $E$ has an algorithmic structure compatible with that of $\mathit{gen}$, monotonicity can be exploited to enable program fusion. In such cases, the $\mathit{min}_{E}$ function applied within $\mathit{genDTs}$ yields an efficient recursive definition for the original specification. The resulting solution is often referred to as a dynamic programming or greedy algorithm.
To illustrate this, we show that when the objective function is defined according to a given general scheme, two DP algorithms can be derived from specifications (28) and (23). Moreover, we can not only prove the existence of DP algorithms in this case, but also demonstrate the non-existence of a DP formulation for (27) and (29).
We consider that the decision tree problem conforms to the following general scheme:
$E$ $:\left(\mathcal{D}\to\mathbb{R}\right)\times\left(\mathbb{R}\times\mathbb{R}\to\mathbb{R}\right)\times\mathit{DTree}\left(\mathcal{R},\mathcal{D}\right)\to\mathbb{R}$ (30)
$E$ $\left(f,g,\mathit{DL}\left(\mathit{xs}\right)\right)=f\left(\mathit{xs}\right)$
[…]
such that $g\left(a,b\right)\geq\max\left(a,b\right)$. For example, the classical objective function used in classification tasks—central to the discussion in Part II of this paper—is the 0-1 loss objective, which can be defined as
$E_{\text{0-1}}$ $:\mathit{DTree}\left(\mathcal{R},\mathcal{D}\right)\to\mathbb{N}$ (31)
$E_{\text{0-1}}$ $\left(\mathit{DL}\left(\mathit{xs}\right)\right)=\sum_{\left(x_{i},y_{i}\right)\in\mathit{xs}}\mathbf{1}\left[\hat{y}\neq y_{i}\right]$
[…]
where $\hat{y}=\underset{k\in\mathcal{K}}{\text{argmax}}\sum_{\left(x_{i},y_{i}\right)\in\mathit{xs}}\mathbf{1}\left[y_{i}=k\right]$, which represents the majority class in a leaf. The following lemma trivially holds.
Lemma 2.
Monotonicity in the decision tree problem. Given left subtrees $u$ and $u^{\prime}$ and right subtrees $v$ and $v^{\prime}$ rooted at $r$, the implication
$E\left(\mathit{mapD}{}_{\cap_{r^{+}}}\left(u\right)\right)\leq E\left(\mathit{mapD}{}_{\cap_{r^{\prime+}}}\left(u^{\prime}\right)\right)\wedge E\left(\mathit{mapD}{}_{\cap_{r^{-}}}\left(v\right)\right)\leq E\left(\mathit{mapD}{}_{\cap_{r^{\prime-}}}\left(v^{\prime}\right)\right)\Longrightarrow E\left(DN\left(u,r,v\right)\right)\leq E\left(DN\left(u^{\prime},r^{\prime},v^{\prime}\right)\right)$ (32)
only holds, in general, if $r=r^{\prime}$.
Dynamic programming algorithm for size-constrained decision tree
Astute readers may have noticed that the recursive structure of $E$ aligns with that of $\mathit{genDTs}_{\text{rec}}$. We now show that this alignment implies a dynamic programming (DP) solution. By leveraging the monotonicity property 2, we can fuse the minimizer $\mathit{min}_{E}$ with $\mathit{genDTs}_{\text{rec}}$ in the simplified decision tree problem defined in 28. This yields a DP solution $\mathit{sodt}_{\text{rec}}$ for the problem $\mathit{min}_{E}\circ\mathit{genDTs}_{\text{rec}}$ 28. Thus providing an efficient algorithm for solving the optimal decision tree problem by substituting the DP solution $\mathit{sodt}_{\text{rec}}$ into $\mathit{odt}_{\text{rec}}$.
Theorem 3.
Dynamic programming algorithm for size-constrained decision tree. Given a list of $K$ rules $\mathit{rs}_{K}:\left[\mathcal{R}\right]$. If the objective function $E$ can be defined according to (30)—that is, if Lemma Lemma 2 holds—then the solution obtained by
$\mathit{sodt}_{\text{rec}}$ $:\mathcal{D}\times\left[\mathcal{R}\right]\to\mathit{DTree}\left(\mathcal{R},\mathcal{D}\right)$ (33)
$\mathit{sodt}_{\text{rec}}$ $\left(\mathit{xs},\left[\;\right]\right)=\left[\mathit{DL}\left(\mathit{xs}\right)\right]$
$\mathit{sodt}_{\text{rec}}$ $\left(\mathit{xs},\left[r\right]\right)=\left[\mathit{DN}\left(\mathit{DL}\left(xs^{+}\right),r,\mathit{DL}\left(xs^{-}\right)\right)\right]$
[…]
provides a solution to the brute-force specification (28). In other words, we have following relation
$\mathit{sodt}_{\text{rec}}\left(\mathit{xs}\right)\subseteq\mathit{min}_{E}\circ\mathit{genDTs}_{\text{rec}}\left(\mathit{xs}\right),$
where the symbol $\subseteq$” means the solution on the left is one of the solution on the right.
Proof.
Due to the monotonicity of the problem, we can now derive the program by following equational reasoning
$\mathit{min}_{E}\circ genDTs_{\text{rec}}\left(\mathit{xs}\right)$
$\equiv$ definition of $genDTs_{\text{rec}}$
$\begin{aligned} &\mathit{minR}{}_{E}\bigg{[}\mathit{DN}\left(\mathit{mapD}{}_{\cap_{\mathit{xs}^{+}}}\left(u\right),r,\mathit{mapD}{}_{\cap_{\mathit{xs}^{-}}}\left(v\right)\right)\mid\\ &\quad\quad\left(\mathit{rs}^{+},r,\mathit{rs}^{-}\right)\leftarrow\mathit{splits}\left(\mathit{rs}\right),u\leftarrow\mathit{genDTs}_{\text{rec}}\left(\mathit{xs},\mathit{rs}^{+}\right),v\leftarrow\mathit{genDTs}_{\text{rec}}\left(\mathit{xs},\mathit{rs}^{-}\right)\bigg{]}\end{aligned}$
$\subseteq$ monotonicity
$\mathit{minR}_{E}\bigg{[}\mathit{DN}\bigg{(}\mathit{minR}_{E}\left[\mathit{mapD}{}_{\cap_{\mathit{xs}^{+}}}\left(u\right)\mid u\leftarrow\mathit{genDTs}_{\text{rec}}\left(\mathit{xs},\mathit{rs}^{+}\right)\right],r,$
$\quad\quad\mathit{minR}_{E}\left[\mathit{mapD}{}_{\cap_{\mathit{xs}^{-}}}\left(u\right)\mid u\leftarrow\mathit{genDTs}_{\text{rec}}\left(\mathit{xs},\mathit{rs}^{-}\right)\right]\bigg{)}\mid\left(\mathit{rs}^{+},r,\mathit{rs}^{-}\right)\leftarrow\mathit{splits}\left(rs\right)\bigg{]}$
$\equiv$ definition of $\mathit{mapL}$
$\begin{aligned} &\mathit{minR}_{E}\bigg{[}\mathit{DN}\bigg{(}\mathit{minR}_{E}\left(mapL_{\mathit{mapD}{}_{\cap_{\mathit{xs}^{+}}}}\left(\mathit{genDTs}_{\text{rec}}\left(\mathit{xs},\mathit{rs}^{+}\right)\right)\right),r,\\ &\quad\quad\mathit{minR}_{E}\left(mapL_{\mathit{mapD}{}_{\cap_{\mathit{xs}^{-}}}}\left(\mathit{genDTs}_{\text{rec}}\left(\mathit{xs},\mathit{rs}^{-}\right)\right)\right)\bigg{)}\mid\left(\mathit{rs}^{+},r,\mathit{rs}^{-}\right)\leftarrow\mathit{splits}\left(\mathit{rs}\right)\bigg{]}\end{aligned}$
$\equiv$ definition of $\mathit{mapD}$
$\begin{aligned} &\mathit{minR}_{E}\bigg{[}\mathit{DN}\bigg{(}\mathit{minR}_{E}\left(\mathit{genDTs}_{\text{rec}}\left(\mathit{r}^{-}\cap\mathit{xs},\mathit{rs}^{+}\right)\right),r,\\ &\quad\quad\mathit{minR}_{E}\left(\mathit{genDTs}_{\text{rec}}\left(\mathit{r}^{-}\cap\mathit{xs},\mathit{rs}^{-}\right)\right)\bigg{)}\mid\left(\mathit{rs}^{+},r_{i},\mathit{rs}^{-}\right)\leftarrow\mathit{splits}\left(\mathit{rs}\right)\bigg{]}\end{aligned}$
$\equiv$ definition of $\mathit{sodt}_{\text{rec}}$
$\mathit{minR}_{E}\bigg{[}\mathit{DN}\bigg{(}\mathit{sodt}_{\text{rec}}\left(\mathit{xs}^{+},\mathit{rs}^{+}\right),r,\mathit{sodt}_{\text{rec}}\left(\mathit{xs}^{-},\mathit{rs}^{-}\right)\bigg{)}\mid\left(\mathit{rs}^{+},r,\mathit{rs}^{-}\right)\leftarrow\mathit{splits}\left(\mathit{rs}\right)\bigg{]}.$
Clearly, the solution produced by $\mathit{sodt}_{\text{rec}}$ is more efficient than the brute-force algorithm $\mathit{min}_{E}\circ genDTs_{\text{rec}}$. The key difference is that the brute-force approach first exhaustively enumerates all possible solutions using $genDTs_{\text{rec}}$ and then selects the best one, whereas $\mathit{sodt}_{\text{rec}}$ recursively constructs the optimal solution $\mathit{sodt}_{\text{rec}}\left(\mathit{xs},\mathit{rs}\right)$ from the optimal sub-solutions $\mathit{sodt}_{\text{rec}}\left(\mathit{xs}^{+},\mathit{rs}^{+}\right)$ and $\mathit{sodt}_{\text{rec}}\left(\mathit{xs}^{-},\mathit{rs}^{-}\right)$. He and Little (2025b) showed that $\mathit{sodt}_{\text{rec}}$ has a worst-case time complexity of $O\left(K!\times N\right)$, assuming that $\mathit{splits}\left(\mathit{rs}\right)$ can be computed in linear time with respect to the size of $\mathit{rs}$, and that $\mathit{xs}^{\pm}$ can be computed in $O\left(N\right)$ time for a data list $\mathit{xs}$ of size $N$.
Consequently, the optimal decision tree problem can be solved using (25), where the $\mathit{sodt}$ is defined by using (33).
4.3.2 Optimal decision tree algorithm for depth-constrained tree
A similar reasoning applies to the ODT problem with depth constraints, commonly used in axis-parallel and binary-feature decision trees. Brița et al. (2025) present an algorithm for this problem; however, instead of formally deriving it, they present it directly. While their approach can be seen as a special case of the algorithm derived here, a formal derivation is important for understanding the underlying principles and for establishing a rigorous foundation.
The following theorem explains why the algorithm of Brița et al. (2025) is correct and how it can be systematically derived from (24).
Theorem 4.
Dynamic programming algorithm for depth-constrained decision tree. Given a depth constraint $d:\mathbb{N}$ and a data list $\mathit{xs}$. If the objective function $E$ can be defined according to (30), i.e., if Lemma 2 holds, then the solution obtained by
$\mathit{odt}_{\text{depth}}$ $:\mathbb{N}\times\mathcal{D}\to\mathit{DTree}\left(\mathcal{R},\mathcal{D}\right)$ (34)
$\mathit{odt}_{\text{depth}}$ $\left(0,\mathit{xs}\right)=\mathit{DL}\left(\mathit{xs}\right)$
$\mathit{odt}_{\text{depth}}$ $\left(1,\mathit{xs}\right)=\mathit{min}_{E}\left[DN\left(\mathit{DL}\left(\mathit{xs}^{+}\right),r,\mathit{DL}\left(\mathit{xs}^{-}\right)\right)\mid r\leftarrow\mathit{gen_{splits}}\left(\mathit{xs}\right)\right]$
$\mathit{odt}_{\text{depth}}$ $\left(d,\mathit{xs}\right)=\mathit{min}_{E}\left[\mathit{DN}\left(\mathit{odt}_{\text{depth}}\left(d-1,\mathit{xs}^{+}\right),r,\mathit{odt}_{\text{depth}}\left(d-1,\mathit{xs}^{-}\right)\right)\mid,r\leftarrow\mathit{gen_{splits}}\left(\mathit{xs}\right)\right]$
is a solution of $\mathit{min}_{E}\circ\mathit{genDTDs}_{\text{depth}}\left(d\right)$. In other words, we have
$\mathit{odt}_{\text{depth}}\subseteq\mathit{min}_{E}\circ\mathit{genDTDs}_{\text{depth}}\left(d\right).$
Proof.
The base cases ($d=1$ and $d=0$) follow directly from the definition of an optimal decision tree:
For $d=0$, no classification is performed; all data points are assigned to a single leaf.
For $d=1$, the optimal decision tree consists of a single splitting rule. Hence, it is equivalent to selecting the optimal splitting rule from $\mathit{gen_{splits}}\left(\mathit{xs}\right)$
For the recursive case, we have:
$\mathit{odt}_{\text{depth}}\left(d,\mathit{xs}\right)=$ $\mathit{min}_{E}\circ\mathit{genDTDs}_{\text{depth}}\left(d\right)$
$=$ $\{\text{definition of $\mathit{genDTDs}_{\text{depth}}$}\}$
$\mathit{min}_{E}\left[\mathit{DN}\left(u,r,v\right)\mid u\leftarrow\mathit{genDTDs}_{\text{depth}}\left(d-1,\mathit{xs}^{+}\right),v\leftarrow\mathit{genDTDs}_{\text{depth}}\left(d-1,\mathit{xs}^{-}\right),r\leftarrow\mathit{gen_{splits}}\left(\mathit{xs}\right)\right]$
$\subseteq$ { monotonicity (2) }
$\left[\mathit{DN}\left(\mathit{min}_{E}\left(\mathit{genDTDs}_{\text{depth}}\left(d-1,\mathit{xs}^{+}\right)\right),r,\mathit{min}_{E}\left(\mathit{genDTDs}_{\text{depth}}\left(d-1,\mathit{xs}^{-}\right)\right)\right)\mid r\leftarrow\mathit{gen_{splits}}\left(\mathit{xs}\right)\right]$
$=$ { definition of $\mathit{odt}_{\text{depth}}$ }
$\left[\mathit{DN}\left(\mathit{odt}_{\text{depth}}\left(d-1,\mathit{xs}^{+}\right),r,\mathit{odt}_{\text{depth}}\left(d-1,\mathit{xs}^{-}\right)\right)\mid r\leftarrow\mathit{gen_{splits}}\left(\mathit{xs}\right)\right]$
This proved that $\mathit{odt}_{\text{depth}}\left(d,\mathit{xs}\right)$ can be computed recursively using optimal subtrees for $\mathit{xs}^{+}$ and $\mathit{xs}^{-}$ of depth d1d-1. ∎
Interestingly, for AODT and ODT over binary-feature data, Brița et al. (2025) and Demirović et al. (2022) observed that $\mathit{odt}_{\text{depth}}\left(2,\mathit{xs}\right)$ can be solved more efficiently than the classical recursive method. Expressing $\mathit{odt}_{\text{depth}}\left(2,\mathit{xs}\right)$ in a more compact form and substituting it into the recursion 34, may help convey their algorithmic insights more clearly.
4.3.3 Non-existence of monotonicity
Another advantage of a rigorous formulation is that it allows us not only to prove the existence of dynamic programming algorithms but also to demonstrate their non-existence when the objective function is restricted to the form. The following theorem formalizes the non-existence of monotonicity for $\mathit{sodt}_{\text{kperms}}$ and $\mathit{sodt}_{\text{rec}}$.
Theorem 5.
Non-existence of monotonicity. For objective function defined according to (30), the one-step recursive process in $\mathit{sodt}_{\text{kperms}}$ and $\mathit{sodt}_{\text{rec}}$ does not have the monotonicity.
Proof.
It suffices to show that the following implication does not hold:
$E\left(t\right)\leq E\left(t^{\prime}\right)\nRightarrow E\left(\mathit{update}\left(r,t\right)\right)\leq E\left(\mathit{update}\left(r,t^{\prime}\right)\right),$
where $\mathit{update}\left(r,t\right)$ denotes adding the rule $r$ to the corresponding leaf of tree $t$. Intuitively, this means that even if $E\left(t\right)\leq E\left(t^{\prime}\right)$ the new tree obtained by adding $r$ to $t$ may create a new tree with higher error than $E\left(t^{\prime}\right)$.
Similarly, for $\mathit{sodt}_{\text{kperms}}$, the following implication fails:
$E\left(\mathit{rs}\right)\leq E\left(\mathit{rs}^{\prime}\right)\nRightarrow E\left(\mathit{insR}\left(r,\mathit{rs}\right)\right)\leq E\left(\mathit{insR}\left(r,\mathit{rs}^{\prime}\right)\right)$
where $\mathit{insR}:\left[\mathcal{R}\right]\to\left[\mathcal{R}\right]$ is the relational version of the $\mathit{ins}$ operation He (2025), which inserts $r$ to a random position in a existing permutation of rules $\mathit{rs}$. The implication fails because the newly inserted rule $r$ can become the root of the resulting tree, which can change the error non-monotonically.
Therefore, dynamic programming or greedy algorithms do not exist for $\mathit{sodt}_{\text{kperms}}$ and $\mathit{sodt}_{\text{vec}}$, when the objective function is defined according to (30). However, this does not imply that DP or greedy solutions are impossible for objectives defined in other form. For instance, if the objective function is incrementally definable—such as the tree size—then a greedy solution exist. ∎
4.3.4 The use of memoization or caching
Given a DP solution for the optimal decision tree problem and the shared subtrees across different trees, it is natural to consider applying memoization, a well-established technique for improving DP efficiency.
Caching provides a practical alternative when a bottom-up recursion or iterative DP is either unavailable or too complex to implement efficiently. This approach has been widely adopted in ODT algorithms over binary feature data. For instance, Aglin et al. (2021) employ caches over itemsets (collections of Boolean features), while Demirović et al. (2022) note a trade-off between branch and dataset caching, as introduced by Nijssen and Fromont (2007).
The correctness of these caching techniques in ODT studies remains questionable. The derivations of $\mathit{odt}_{\text{depth}}$ and $\mathit{odt}_{\text{size}}$ depend on Lemma 2, which holds only when the roots of two trees, $r$ and $r^{\prime}$ are equivalent. Therefore, correct use of memoization requires storing both the root and its corresponding subtrees, because optimality of subtrees $u$ and $v$ over subregions $r^{\pm}$ implies optimality of the full tree $DN\left(u,r,v\right)$. Thus, no trade-off exists between branch and dataset caching: both subtrees (depth- or size-constrained) and datasets must be stored. In this context, the caching strategy in Brița et al. (2025) appears correct, as their ConTree algorithm reuses cached solutions defined both by sub-dataset and subtree depth.
However, He and Little (2025b) predicted that memoization becomes ineffective when the number of possible splitting rules is large. This is because each rule $r$ generates a distinct sub-dataset, resulting in few identical optimal subtrees for a given sub-dataset. Denote the set of possible splitting rules with respect to data list $\mathit{xs}$ as $\mathcal{S}_{\mathit{\text{split}}}\left(\mathinner{xs}\right)$, to enable full memoization for constructing an optimal decision tree of size $K$, one would need to store at least:
$O\left(\left|\mathcal{S}_{\mathit{\text{split}}}\left(\mathinner{xs}\right)\right|\times\sum_{k=0}^{K-1}\left|\mathcal{S}_{\text{size}}\left(k,\mathinner{xs}\right)\right|\right)$
Similarly, for depth-constrained trees, the combinatorial complexity becomes:
$O\left(\left|\mathcal{S}_{\mathit{\text{split}}}\left(\mathinner{xs}\right)\right|\times\sum_{i=0}^{d-1}\left|\mathcal{S}_{\text{depth}}\left(i,\mathinner{xs}\right)\right|\right)$
This results in formidable combinatorial complexity for most ODT problems. For example, in the AODT problem, the number of possible rules is $\mathcal{S}_{\mathcal{H}^{0}}=O\left(N\times D\right)$. For hyperplane ODT problems, the number of possible splits grows to $\mathcal{S}_{\mathcal{H}}=O\left(N^{{}^{D}}\right)$ (see Section 5). Consequently, the caching approach in Brița et al. (2025), which stores only a few thousand optimal sub-trees, has a very low probability of cache hits. Our tests on multiple data lists confirm this: in extreme cases, their algorithm examined millions of decision trees without achieving a single cache hit. This suggests that the efficiency claimed for their ConTree algorithm due to caching is overstated; most gains likely arise from factors other than caching.
4.4 Further acceleration—Thinning and filtering
4.4.1 Filtering
In machine learning, a common way to prevent overfitting is to require each leaf node to contain at least $N_{\min}$ data points, avoiding situations where a leaf has only a very small number of samples. A straightforward way to enforce this constraint is to introduce a filtering process:
$\mathit{sodt}_{\text{filt}}$ $=\mathit{min}_{E}\circ\mathit{filter}_{p}\circ\mathit{genDTs}$ (35)
$\mathit{odt}_{\text{filtDep}}$ $=\mathit{min}_{E}\circ\mathit{filter}_{p}\circ\mathit{genDTDs}_{\text{depth}}$ (36)
where $p:\mathit{DTree}\left(\mathcal{R},\mathcal{D}\right)\to\mathit{Bool}$ is a Boolean-valued predicate.
However, this direct approach is inefficient, as $\mathit{genDTs}$ and $\mathit{genDTDs}_{\text{depth}}$ generate a huge number of trees, making post-generation filtering costly. When the predicate $p$ satisfies certain conditions, filtering can be fused with tree generation, avoiding unnecessary computation. Specifically, the following condition that help identify how filtering can be fused into the generation process.
Definition 3.
Prefix-closed predicate. A predicate $p:\mathit{DTree}\left(\mathcal{R},\mathcal{D}\right)\to\mathit{Bool}$ is called prefix-closed if for any tree $\mathit{DN}\left(u,r,v\right)$, the following implication holds:
$p\left(\mathit{DN}\left(u,r,v\right)\right)\implies p\left(u\right)\wedge p\left(v\right).$ (37)
For example, the predicate that checks whether a tree contains a leaf with at least $N_{\text{min}}$ data points is prefix-closed, because if a tree $\mathit{DN}\left(u,r,v\right)$ violates this condition, then its subtrees then $u$ and $v$ also violate it. Similarly, a tree-size predicate, which measures the number of splitting rules, is prefix-closed: $\mathit{DN}\left(u,r,v\right)$ has fewer than $K$ rules, then both $u$ and $v$ also have fewer than $K$ rules. The same holds for depth constraints: if a tree has depth less than $d$, its subtrees also satisfy this bound.
Prefix-closed predicates provide an effective mechanism to fuse the filtering operation $\mathit{filter}_{p}$ with the generator $\mathit{gen}$. In particular, the following filter fusion theorem enables the integration of depth- or size-constrained conditions into leaf- or size-constrained ODT problems.
Theorem 6.
Filtering fusion through prefix-predicate. If a predicate $p$ is prefix-closed, Then, the solution obtained using $\mathit{sodt}_{\text{filtRec}}$, defined as
$\mathit{sodt}_{\text{filtRec}}$ $:\left[\mathcal{R}\right]\times\mathcal{D}\to\mathit{DTree}\left(\mathcal{R},\mathcal{D}\right)$ (38)
$\mathit{sodt}_{\text{filtRec}}$ $\left(\left[\;\right],\mathit{xs}\right)=\left[\mathit{DL}\left(\mathit{xs}\right)\right]$
$\mathit{sodt}_{\text{filtRec}}$ $\left(\left[r\right],\mathit{xs}\right)=\left[\mathit{DN}\left(\mathit{DL}\left(xs^{+}\right),r,\mathit{DL}\left(xs^{-}\right)\right)\right]$
[…]
provides a solution to $\mathit{min}_{E}\circ\mathit{filter}_{p}\circ\mathit{genDTs}_{\text{rec}}$. Similarly, the solution obtained using $\mathit{odt}_{\text{filtDep}}$, defined as
$\mathit{odt}_{\text{filtDep}}$ $:\mathbb{N}\times\mathcal{D}\to\mathit{DTree}\left(\mathcal{R},\mathcal{D}\right)$ (39)
$\mathit{odt}_{\text{filtDep}}$ $\left(0,\mathit{xs}\right)=\mathit{DL}\left(\mathit{xs}\right)$
$\mathit{odt}_{\text{filtDep}}$ $\left(1,\mathit{xs}\right)=\mathit{min}_{E}\left[DN\left(\mathit{DL}\left(\mathit{xs}^{+}\right),r,\mathit{DL}\left(\mathit{xs}^{-}\right)\right)\mid r\leftarrow\mathit{gen_{splits}}\left(\mathit{xs}\right)\right]$
$\mathit{odt}_{\text{filtDep}}$ $\left(d,\mathit{xs}\right)=\mathit{min}_{E}\left(\mathit{filter}_{p}\left[\mathit{DN}\left(\mathit{odt}_{\text{filtDep}}\left(d-1,\mathit{xs}^{+}\right),r,\mathit{odt}_{\text{filtDep}}\left(d-1,\mathit{xs}^{-}\right)\right)\mid r\leftarrow\mathit{gen_{splits}}\left(\mathit{xs}\right)\right]\right)$
provides a solution to (36).
Proof.
The proof follows by substituting the fused generator and following the procedure of Theorems (3) and (4). For more general results and formal proofs, see He (2025). ∎
Theorem 6 provides an efficient and convenient method to incorporate any prefix-closed constraint into the ODT algorithm. Any partial solution that does not satisfy the prefix-closed constraint can be filtered out before extending it into a complete tree, thus reducing unnecessary computation while ensuring constraints are incorporated.
The derivation for $\mathit{sodt}_{\text{vec}}$ and $\mathit{sodt}_{\text{kperms}}$ similar to Theorem 6, but requires a different prefix-closed predicate definition, as the recursive structure of this two algorithms are sequential. S $e$ He (2025) for more general results.
4.4.2 Thinning algorithm (bounding technique)
Thin introduction
The thinning algorithm is almost identical to the exploitation of dominance relations in the algorithm design literature (Ibaraki, 1977; Bird and De Moor, 1996). Both thinning and classical dominance relations aim to improve the time complexity of naive dynamic programming algorithms (Bird and De Moor, 1996). The key difference is that thinning considers not only the theoretical effectiveness of bounds but also their practical implementation. This aspect is often overlooked in studies of branch-and-bound algorithms, where researchers focus on the existence alone of a bounding relation but neglect its efficiency. As noted in subSection 2.1, a bounding technique must be proven effective: the computational time it saves must outweigh the time required to execute it. If a bounding check is computationally expensive yet reduces the solution space only marginally, its application becomes ineffective.
The thinning technique leverages the fundamental observation that certain partial configurations are superior to others. Extending non-optimal partial configurations is thus a waste of computational resources. Formally, we define a function $\mathit{thin}:\left(\mathcal{A}\times\mathcal{A}\to\mathit{Bool}\right)\times\left[\mathcal{A}\right]\to\left[\mathcal{A}\right]$ which receives a relation $R:\mathcal{A}\times\mathcal{A}\to\mathit{Bool}$, referred to as a dominance relation, and a sequence $\mathit{xs}$. The output is a subsequence $\mathit{ys}\subseteq\mathit{xs}$, satisfying:
$\forall x\in\mathit{xs}:\exists y\in\mathit{ys}:R\left(y,x\right)$ (40)
In other words, for every element in the original sequence, there exists a dominating element in the thinned subsequence, ensuring that only potentially optimal configurations are retained for further computation. Properties (1) and (2) are known as the universal property of the thinning algorithm (Bird and De Moor, 1996).
Thinning is similar to, but distinct from, $\mathit{min}_{E}$. Indeed, the $\mathit{min}_{E}$ function can be understood as a special case of thinning with respect to a total order defined by the objective function $E$. In contrast, thinning is based on a preorder $R$, because transitivity is required, and total orders are too restrictive for the purposes of thinning (He, 2025; Bird and Gibbons, 2020). In a preorder relation, some configurations may be incomparable. Consequently, $\mathit{thin}_{R}:\left[\mathcal{A}\right]\to\left[\mathcal{A}\right]$ receives a list and returns a list, whereas $\mathit{min}_{E}:\left[\mathcal{A}\right]\to\mathcal{A}$ always returns a single element.
The thinning algorithm must satisfy the following laws:
$\mathit{min}_{E}$ $=\mathit{min}_{E}\circ\mathit{thin}_{R_{2}}$ (41)
$\mathit{mapL}_{f}\circ\mathit{thin}_{R_{1}}$ $\subseteq\mathit{thin}_{R_{2}}\circ\mathit{mapL}_{f}$ (42)
such that $R_{1}\left(a,b\right)\implies R_{2}\left(a,b\right)$. The first law (41) is called thin introduction. It states that applying thinning does not change the optimal solution: the optimal solution after thinning is the same as the original optimal solution. The second law (42), called thin-map, governs the interaction between a map operation and thinning: mapping over a thinned sequence produces a result contained within the thinning of the mapped sequence.
Monotonicity and thinning
Similar to monotonicity used in designing efficient combinatorial optimization algorithms, it also plays a crucial role in identifying valid dominance relations that enable thinning fusion. Intuitively, monotonicity ensures that if a partial configuration $a$ dominates another configuration $b$, then extending both configurations in the same way preserves the dominance relation. Without this property, thinning could discard potentially optimal solutions.
Definition 4.
Monotonicity in thinning. Let $f:\mathcal{A}\to\mathcal{A}$ be a one-step update relation in a recursion (sequential decision process), We say that $f$ is monotonic with respect to a relation $R:\mathcal{A}\times\mathcal{A}\to\mathit{Bool}$ if
$R\left(a,b\right)\Longrightarrow R\left(f\left(a\right),f\left(b\right)\right)$ (43)
In words, if configuration $a$ dominates configuration $b$, then after applying the same update $f$ to both, the dominance relation still holds.
For instance, in $\mathit{sodt}_{\text{vec}}$, the semantic interpretation of monotonicity in thinning is expressed as: $R\left(t,t^{\prime}\right)\Longrightarrow R\left(\mathit{update}\left(r,t\right),\mathit{update}\left(r,t^{\prime}\right)\right)$. Similarly, in $\mathit{sodt}_{\text{kperms}}$, we have: $R\left(\mathit{rs},\mathit{rs}^{\prime}\right)\Longrightarrow R\left(\mathit{insR}\left(r,\mathit{rs}\right),\mathit{insR}\left(r,\mathit{rs}^{\prime}\right)\right)$. Likewise, in $\mathit{sodt}_{\text{rec}}$, the semantic interpretation of monotonicity in thinning can be expressed as $R\left(u,u^{\prime}\right)\wedge R\left(v,v^{\prime}\right)\subseteq R\left(\mathit{DN}\left(u,r,v\right),\mathit{DN}\left(u^{\prime},r,v^{\prime}\right)\right)$.
Finite dominance relation and global upper bound
There are many ways to define a thinning algorithm He (2025), and its effectiveness depends on the context. For our purposes—the study of optimal decision tree problems—the dominance relation allows us to implement $\mathit{thin}$ simply using the $\mathit{filter}$ function. Therefore, we will not discuss the general construction of thinning, but instead focus on formalizing the specific dominance relations studied in ODT algorithms for machine learning research.
He (2025) showed that dominance relations commonly used in ML can be abstracted into two types: finite dominance relation (FDR) and global upper bound (GUB). Both relations rely on two auxiliary functions defined as follows.
Definition 5.
Pessimistic upper bound and optimistic lower bound. Given a configuration $a:\mathcal{A}$, and let $f:\mathcal{A}\to\mathcal{A}$ be a one-step update relation [^2] In the ODT problem, for $\mathit{sodt}_{\text{rec}}$ and $\mathit{odt}_{\text{depth}}$, $f$ corresponds to the tree constructor $\mathit{DN}$ with the same root $r$ and left/right subtrees $t$ but different right/left subtrees $x$ and $y$. In $\mathit{sodt}_{\text{vec}}$ and $\mathit{sodt}_{\text{kperms}}$, $f$ is a parameterized function, either $\mathit{update}\left(r\right)$ or $\mathit{insR}\left(r\right)$ for adding a new splitting rule $r$ to a existing tree. in a recursion. Assume smaller values are better: a function $\mathit{ub}$ is called pessimistic upper bound if $\mathit{ub}\left(f^{n}\left(a\right)\right)\leq\mathit{ub}\left(a\right)$. A function $\mathit{lb}:\mathcal{A}\times\mathcal{D}\to\mathbb{R}$ is called optimistic lower bound if $\mathit{lb}\left(f^{n}\left(a\right)\right)\geq\mathit{lb}\left(a\right)$.
Intuitively, these bounds capture the idea that as more information is accumulated, the bounds become tighter, assume small is better:
  • The pessimistic upper bound provides a “worst-case” estimate of $a$, so any update $f^{n}\left(a\right)$ should yield a tighter (smaller) upper bound.
  • The optimistic lower bound provides a “best-case” estimate of $a$, so any update $f^{n}\left(a\right)$ should yield a looser (larger) lower bound.
Many existing studies on ODT problems focus on developing effective optimistic lower bounds or pessimistic upper bounds. He (2025) demonstrated that the following two generic dominance relations, derived using $\mathit{ub}$ and $\mathit{lb}$, satisfy the thinning laws. These abstract the dominance relations used in most optimal decision tree algorithms for machine learning.
Theorem 7.
Finite dominance relations and global upper bound. Given two partial configuration $a,b:\mathcal{A}$, and a global upper-bound constant $\mathit{ub}^{*}:\mathcal{A}$, defines
  1. 1.
    Finite dominance relation:
    $\mathit{fdom}\left(a,b\right)=\mathit{ub}\left(a\right)\leq\mathit{lb}\left(b\right).$
  2. 2.
    Global upper bounds dominance relation:
    $\mathit{gubdom}\left(a\right)=\mathit{ub}^{*}\leq\mathit{lb}\left(a\right).$
Then both $\mathit{fdom}$ and $\mathit{gubdom}$ satisfies the monotonicity.
Proof.
To prove the dominance valid, we nee to verify the thinning algorithm defined using $\mathit{fdom}$ satisfies the universal property, which is equivalent to verifying the following monotonicity (Bird and De Moor, 1996):
$\mathit{ub}\left(a\right)\leq\mathit{lb}\left(b\right)\Longrightarrow\mathit{ub}\left(f^{n}\left(a\right)\right)\leq\mathit{lb}\left(f^{n}\left(b\right)\right)$
The implication holds because $\mathit{ub}\left(f^{n}\left(a\right)\right)\leq\mathit{ub}\left(a\right)\leq\mathit{lb}\left(b\right)\leq\mathit{lb}\left(f^{n}\left(b\right)\right)$. Similarly, to prove $\mathit{gubdom}$, we need to verify
$\mathit{ub}^{*}\leq\mathit{lb}\left(a\right)\Longrightarrow\mathit{ub}^{*}\leq\mathit{lb}\left(f^{n}\left(a\right)\right)$
which holds because $\mathit{ub}^{*}\leq\mathit{lb}\left(a\right)\leq\mathit{lb}\left(f^{n}\left(a\right)\right)$. ∎
For the global upper bound dominance relation, the thinning function $\mathit{thin}:\left[\mathcal{A}\right]\to\left[\mathcal{A}\right]$ can be efficiently implemented using a simple $\mathit{filter}$ He (2025).
Examples of lower bound in literature
We now introduce two useful and generic lower bounds proposed by Demirović et al. (2022) and Zhang et al. (2023) for classification and regression problems, respectively.
One particularly interesting and widely applicable bounding technique was introduced by Demirović et al. (2022); Lin et al. (2020); Brița et al. (2025) for classification problems. This technique is based on the following observation, although it relies on an assumption that has not yet been formally proven.
Definition 6.
Similarity lower bound. Let $\mathit{xs}_{1}$ and $\mathit{xs}_{2}$ denotes two different sets, and $\mathit{xs}_{\text{1}}-\mathit{xs}_{2}$ denote the list of elements in $\mathit{xs}_{1}$ but not in $\mathit{xs}_{2}$, let $E_{\text{0-1}}\left(\mathit{odt}_{\text{depth}}\left(d,\mathit{xs}_{1}\right),\mathit{xs}_{2}\right)$ denote the number of misclassifications when the optimal decision tree of fixed depth $d$ trained on data list $\mathit{xs}_{1}$, is evaluated on data list $\mathit{xs}_{2}$, (For readability, we write $\mathit{odt}_{\text{depth}}\left(\mathit{xs}\right)$ with $d$ fixed by default.)
The similarity lower bound is based on the following observation:
$E_{\text{0-1}}\left(\mathit{odt}_{\text{depth}}\left(\mathit{xs}_{\text{2}}\right),\mathit{xs}_{\text{2}}\right)\geq E_{\text{0-1}}\left(\mathit{odt}_{\text{depth}}\left(\mathit{xs}_{1}\cap\mathit{xs}_{\text{2}}\right),\mathit{xs}_{1}\cap\mathit{xs}_{\text{2}}\right)\geq E_{\text{0-1}}\left(\mathit{odt}_{\text{depth}}\left(\mathit{xs}_{\text{1}}\right)\right)-\left|\mathit{xs}_{1}-\mathit{xs}_{2}\right|$ (44)
The last inequality states that removing samples $\mathit{xs}_{1}-\mathit{xs}_{2}$ from data list $\mathit{xs}_{1}$ can reduce the number of misclassifications by at most $\left|\mathit{xs}_{1}-\mathit{xs}_{2}\right|$. Although looks intuitively obvious. In the worst case, assume all removed samples were misclassified. Then:
$E_{\text{0-1}}\left(\mathit{odt}_{\text{depth}}\left(\mathit{xs}_{1}\cap\mathit{xs}_{\text{2}}\right),\mathit{xs}_{1}\cap\mathit{xs}_{\text{2}}\right)=E_{\text{0-1}}\left(\mathit{odt}_{\text{depth}}\left(\mathit{xs}_{\text{1}}\right),\mathit{xs}_{\text{1}}\right)-\left|\mathit{xs}_{1}-\mathit{xs}_{2}\right|$
This equality implicitly suggests that the optimal decision tree trained on the larger data list $\mathit{xs}_{\text{1}}$ is also optimal on the smaller data list $\mathit{xs}_{1}\cap\mathit{xs}_{\text{2}}$. If this were not true, then there would exist another tree $t_{\text{opt}}^{\prime}$ such that
$E_{\text{0-1}}\left(t_{\text{opt}}^{\prime},\mathit{xs}_{1}\cap\mathit{xs}_{\text{2}}\right)<E_{\text{0-1}}\left(\mathit{odt}_{\text{depth}}\left(\mathit{xs}_{\text{1}}\right),\mathit{xs}_{\text{1}}\right)-\left|\mathit{xs}_{1}-\mathit{xs}_{2}\right|,$
contradicting the bound. This assumption—that the optimal tree for a larger data list remains optimal for its subset—is not formally proven and is somewhat counterintuitive. Clarifying or proving this property is necessary to ensure the validity of the similarity lower bound.
Example 2: $K$-means lower bound
Another interesting lower bound was introduced in the context of the optimal regression tree problem by Zhang et al. (2023). It captures the key insight that the regression error within a leaf cannot be lower than the optimal solution to the one-dimensional $K$-means clustering problem over the corresponding labels.
Theorem 8.
$K$-means lower bound. Given a set of rules $\mathit{rs}_{K}$ rules, a list of data points $\mathit{xs}$, and a list of corresponding labels $\mathit{ys}:\left[\mathbb{R}\right]$ associated with $\mathit{xs}$, let the objective function $E_{2}$ denote the ordinary squared loss. Define $\text{kmeans-1D}\left(k,\mathit{ys}\right)$ as the optimal cost of performing $k$-means clustering over a one-dimensional data list $\mathit{ys}:\left[\mathbb{R}\right]$. Then, we have the following lower bound:
$E_{2}\left(\mathit{sodt}\left(\mathit{xs},\mathit{rs}_{K}\right),\mathit{ys}\right)\geq\text{kmeans-1D}\left(K+1,\mathit{ys}\right)$
Proof.
The result follows directly, as a decision tree can be viewed as a method of clustering a one-dimensional data list—a label vector. Its clustering performance cannot surpass that of exact 1D $K$-means clustering, which is optimal with respect to squared loss. Consequently, any other clustering over $\mathit{ys}$, including that induced by the decision tree, will exhibit a greater squared loss. ∎
Fusion for thinning algorithm
We now derive the program of fusing thinning algorithm in the recursion of $\mathit{odt}_{\text{depth}}$ when the dominance relation is defined as $\mathit{fdom}$ or $\mathit{gubdom}$. The thinning algorithm for $\mathit{sodt}_{\text{rec}}$ can be derived in a similar manner.
Lemma 3.
Thinning fusion by finite dominance relation. Given a pessimistic upper bound $\mathit{ub}$ and an optimistic lower bound $\mathit{lb}$, which defines a finite dominance relation $\mathit{fdom}_{1}\left(t,\mathit{t}^{\prime}\right)=\mathit{ub}\left(d,\mathit{xs}\right)\leq\mathit{lb}\left(d,\mathit{xs}\right)$ and $\mathit{fdom}_{2}\left(t,\mathit{t}^{\prime}\right)=\mathit{ub}\left(\mathit{odt}_{\text{depth}}\left(d,\mathit{xs}\right)\right)\leq\mathit{lb}\left(\mathit{odt}_{\text{depth}}\left(d,\mathit{xs}\right)\right)$, then the solution of
$\mathit{odt}_{\text{thindep}}$ $:\mathbb{N}\times\mathcal{D}\to\mathit{DTree}\left(\mathcal{R},\mathcal{D}\right)$ (45)
$\mathit{odt}_{\text{thindep}}$ $\left(0,\mathit{xs}\right)=\mathit{DL}\left(\mathit{xs}\right)$
$\mathit{odt}_{\text{thindep}}$ $\left(1,\mathit{xs}\right)=\mathit{min}_{E}\left[DN\left(\mathit{DL}\left(\mathit{xs}^{+}\right),r,\mathit{DL}\left(\mathit{xs}^{-}\right)\right)\mid r\leftarrow\mathit{gen}_{\mathit{splits}}\left(\mathit{xs}\right)\right]$
$\mathit{odt}_{\text{thindep}}$ $\left(d,\mathit{xs}\right)=\mathit{min}_{E}\left(\mathit{mapL}_{f}\left(\mathit{thin}_{\mathit{fdom}_{1}}\left(\left[\left(r,d,\mathit{xs}\right)\mid r\leftarrow\mathit{gen_{splits}}\left(\mathit{xs}\right)\right]\right)\right)\right)$
where $f\left(r,d,\mathit{xs}\right)=\mathit{DN}\left(\mathit{odt}_{\text{depth}}\left(d-1,\mathit{xs}^{+}\right),r,\mathit{odt}_{\text{depth}}\left(d-1,\mathit{xs}^{-}\right)\right)$, is also a solution of $\mathit{odt}_{\text{depth}}$. In particular when $r$ is the defined as the global upper bound with a fiction configuration $\mathit{ub}^{*}$, we can implement $\mathit{thin}$ using $\mathit{filter}$, then we have have
$\mathit{odt}_{\text{thindep}}$ $:\mathbb{N}\times\mathcal{D}\to\mathit{DTree}\left(\mathcal{R},\mathcal{D}\right)$
$\mathit{odt}_{\text{thindep}}$ $\left(0,\mathit{xs}\right)=\mathit{DL}\left(\mathit{xs}\right)$
$\mathit{odt}_{\text{thindep}}$ $\left(1,\mathit{xs}\right)=\mathit{min}_{E}\left[DN\left(\mathit{DL}\left(\mathit{xs}^{+}\right),r,\mathit{DL}\left(\mathit{xs}^{-}\right)\right)\mid r\leftarrow\mathit{gen}_{\mathit{splits}}\left(\mathit{xs}\right)\right]$
[…]
where $E\left(t^{*}\right)=\inf$.
Proof.
We prove 45 by following equational reasoning
$\mathit{min}_{E}\left(\left[\mathit{DN}\left(\mathit{odt}_{\text{depth}}\left(d-1,\mathit{xs}^{+}\right),r,\mathit{odt}_{\text{depth}}\left(d-1,\mathit{xs}^{-}\right)\right)\mid r\leftarrow\mathit{gen_{splits}}\left(\mathit{xs}\right)\right]\right)$
$=$ { thin-introduction law}
$\mathit{min}_{E}\left(\mathit{thin}_{R_{2}}\left(\left[\mathit{DN}\left(\mathit{odt}_{\text{depth}}\left(d-1,\mathit{xs}^{+}\right),r,\mathit{odt}_{\text{depth}}\left(d-1,\mathit{xs}^{-}\right)\right)\mid r\leftarrow\mathit{gen_{splits}}\left(\mathit{xs}\right)\right]\right)\right)$
$=$ { $\mathit{odt}_{\text{depth}}\left(d,\mathit{xs}\right)$ is abstracted as $\mathit{mapL}_{f}$}
$\mathit{min}_{E}\left(\mathit{thin}_{R_{1}}\left(\mathit{mapL}_{f}\left(\left[\left(r,d,\mathit{xs}\right)\mid r\leftarrow\mathit{gen_{splits}}\left(\mathit{xs}\right)\right]\right)\right)\right)$
$\subseteq$ { thin-map law, when $R_{1}=\mathit{fdom}_{1}$}
$\mathit{min}_{E}\left(\mathit{mapL}_{f}\left(\mathit{thin}_{fdom_{1}}\left(\left[\left(r,d,\mathit{xs}\right)\mid r\leftarrow\mathit{gen_{splits}}\left(\mathit{xs}\right)\right]\right)\right)\right)$
In particular, when $R_{1}=\mathit{gubdom}$, then thinning algorithm $\mathit{thin}_{R_{2}}$ is just the filter function $\mathit{filter}$, then we have
$\mathit{min}_{E}\left(\mathit{thin}_{R_{1}}\left(\mathit{mapL}_{f}\left(\left[\left(r,d,\mathit{xs}\right)\mid r\leftarrow\mathit{gen_{splits}}\left(\mathit{xs}\right)\right]\right)\right)\right)$
$\subseteq$ { thin-map law, when $R_{1}=\mathit{gubdom}$}
$\mathit{min}_{E}\left(\mathit{mapL}_{f}\left(\mathit{filter}_{p_{\left(lb,\mathit{ub}^{*}\right)}}\left(\left[\left(r,d,\mathit{xs}\right)\mid r\leftarrow\mathit{gen_{splits}}\left(\mathit{xs}\right)\right]\right)\right)\right)$
$=$ { definition of $\mathit{filter}$}
[…]
Note that the thinning algorithm does not necessarily guarantee a speedup compared to the original definition of $\mathit{odt}_{\text{depth}}$. Its performance depends on both the implementation of $\mathit{thin}$ and the effectiveness of the relation $R$. He (2025) provides various recursive definitions of $\mathit{thin}$ and the effectiveness of each definition is problem-dependent.
4.5 Comparison of four methods
4.5.1 Comparison between size- and depth-constrained trees algorithms
As noted, depth-constrained decision trees are commonly used in ODT studies with relatively small combinatorics, such as axis-parallel trees (Mazumder et al., 2022; Brița et al., 2025) trees over binary-feature data (Demirović et al., 2022; Verwer and Zhang, 2019; Nijssen and Fromont, 2007; Aglin et al., 2020; Nijssen and Fromont, 2010). These existing methods typically process recursion in a depth-first manner, unlike the breadth-first formulation we defined in 34. In contrast, the size-constrained ODT algorithm has been far less studied; Hu et al. (2019); Lin et al. (2020); Zhang et al. (2023) have applied it to solve the ODT problem over binary-feature data, but the correctness of their algorithms remains unresolved.
This preference is not coincidental. The discussion below clarifies why researchers favor depth-constrained approaches and why the depth-constrained ODT algorithm 34 is unsuitable for problems with intractable combinatorics, such as hypersurface decision trees. The primary reason is that such problems require parallelization to achieve reasonable solution times, yet depth-constrained trees are difficult to implement in an embarrassingly parallel manner—parallelization requires minimal or no communication between processors.
We first analyze the size-constrained decision tree and explain why it is suitable for parallel implementation. In 25, the original size-constrained ODT problem can be factorized as:
$\mathit{min}_{E}\circ\mathit{concatMapL}{}_{\mathit{sodt}\left(\mathit{xs}\right)}\circ\mathit{kcombs}_{K}=\mathit{odt}_{K}$
This factorization indicates that solving the challenging $\mathit{odt}$ problem can be achieved by solving many smaller instances of $\mathit{sodt}$. As He and Little (2025a) demonstrated, there exists an elegant and efficient divide-and-conquer definition for $\mathit{kcombs}_{K}$, which decomposes the problem into smaller, independent subproblems. These subproblems can be processed in parallel and mapped efficiently onto modern memory hierarchies, particularly when combined with recursively blocked arrays (Gupta et al., 2000). Unfortunately, no similar factorization has yet been found for the depth-constrained ODT problem that would enable embarrassingly parallel execution. Consequently, the optimal decision tree algorithm derived from 23 can only be executed sequentially or parallelized post hoc, making it difficult to achieve the same level of speed-up as embarrassingly parallel algorithms.
For hyperplane decision trees, each recursive step may involve up to $O\left(N^{D}\right)$ possible rules. Adopting a breadth-first search strategy would result in extremely slow branching and high memory consumption. When exact solutions are infeasible—as is the case for most data lists with large $d$ and NN—depth-first search becomes necessary, as it can provide approximate solutions within a fixed time limit.
Although depth-first search effectively reduces memory usage, it significantly complicates parallelization compared to breadth-first strategies. Previous research on ODT problems with smaller combinatorics remains tractable even without parallelization because the number of possible splits at each recursive step is limited to $O\left(N\times D\right)$ for axis-parallel trees. However, this approach becomes ineffective when applied to the larger combinatorics of 23 where splitting rules are defined as hyperplanes. Hence, parallelization is likely the only viable solution for problems with such large combinatorial complexity.
Nevertheless, developing a combinatorial characterization for depth-constrained decision trees that allows parallel factorization, similar to the size-constrained ODT problem, represents an interesting direction for future research.
4.5.2 Comparison between size-constrained optimal decision tree algorithms
Aspect $\mathit{sodt}_{\text{rec}}$ $\mathit{sodt}_{\text{vec}}$ $\mathit{sodt}_{\text{kperms}}$
Advantages Lower computational cost Elegant and easily implemented recursive definition Facilitates recursive tree evaluation Low probability of worst-case complexity Memory efficient Trackable, iterative computation enabling full vectorization Support incremental objective evaluation Trackable, iterative computation enabling full vectorization Moderate probability of worst-case complexity
Disadvantages Risk of stack overflow in deep recursion due to stack frame allocation Limited support for tail call optimization (TCO) in imperative languages Challenges in parallelization due to sequential call stack Expensive unique operation for large $K$ Costly feasibility testing ($O\left(K^{2}\right)$ in worst case) High probability of worst-case complexity Memory intensive due to vectorized implementation Generates duplicate $K$-permutations corresponds to the same tree Costly feasibility testing ($O\left(K^{2}\right)$ in worst case) Memory intensive due to vectorized implementation Does not support incremental objective evaluation
Table 1: Comparison of advantages and disadvantages of $\mathit{sodt}_{\text{rec}}$, $\mathit{sodt}_{\text{vec}}$ and $\mathit{sodt}_{\text{kperms}}$ algorithms
Methods Theoretical Computational Efficiency Memory Usage Parallelizability Recursive Evaluability
$\mathit{odt}_{\text{kperms}}$ 2 3 1 5
$\mathit{odt}_{\text{rec}}$ 1 2 2 1
$\mathit{odt}_{\text{vec}}$ 2 4 1 3
$\mathit{odt}_{\text{depth}}$ 1 1 4 1
Table 2: Comparison $\mathit{odt}_{\text{rec}}$, $\mathit{odt}_{\text{vec}}$, $\mathit{odt}_{\text{kperms}}$, and $\mathit{odt}_{\text{depth}}$ algorithms, we provide a ranking on a scale from 1 to 5: 1 indicates easy, 2 indicates relatively easy but less efficient than 1, 3 indicates difficult but feasible, 4 indicates very difficult or involving a highly complex process, and 5 indicates infeasible or requiring prohibitive computational resources.
Having discussed the advantages of depth- and size-constrained decision trees, we believe size-constrained ODT algorithms are more suitable candidates for solving complex ODT problems. We present three distinct definitions of $\mathit{odt}_{\text{size}}$: $\mathit{odt}_{\text{rec}}$, $\mathit{odt}_{\text{vec}}$ and $\mathit{odt}_{\text{kperms}}$—naturally raising the question: which is most suitable?
In machine learning research, the interplay between software, hardware, and algorithms is often overlooked. An algorithm’s success may derive not from its intrinsic superiority but from the tools—hardware and software—used in its implementation. This phenomenon, termed the “hardware lottery” by Hooker (2021) highlights how the siloed development of hardware, software, and algorithms has historically shaped which approaches prevail. A prominent example is the success of deep neural networks, which benefited substantially from optimized hardware and software.
Similarly, in the ODT problem, a comparable “lottery” exists between the vectorized method $\mathit{odt}_{\text{vec}}$ and recursive method $\mathit{odt}_{\text{rec}}$ for solving optimal decision tree problems. Because $\mathit{odt}_{\text{size}}$ can be factorized into many smaller instances solved using $\mathit{sodt}$, the algorithmic behavior is largely determined by the performance of $\mathit{sodt}$. Therefore, we focus our discussion on the behavior of these algorithms.
Compare with $\mathit{sodt}_{\text{rec}}$, $\mathit{sodt}_{\text{vec}}$ and $\mathit{sodt}_{\text{kperms}}$, $\mathit{sodt}_{\text{rec}}$ not only requires provably fewer computations due to its DP solution but also has a lower probability of incurring worst-case complexity. Due to the use of tree datatype, the recursive pattern of $\mathit{sodt}_{\text{rec}}$ behaves like a divide-and-conquer algorithm, so the combinatorial complexity of subtrees decreases logarithmically with respect to the length of $\mathit{rs}$ in typical cases, whereas $\mathit{sodt}_{\text{vec}}$ decreases linearly due to its intrinsic sequential nature. Furthermore, $\mathit{sodt}_{\text{rec}}$ creates no duplicate trees, eliminating the need for a $\mathit{unique}$ operation. However, its drawbacks are significant: it is difficult to optimize and carries a risk of stack overflow due to recursive calls. Moreover, $\mathit{sodt}_{\text{rec}}$ is a data-dependent recursion, in which the function’s behavior and termination depend on the structure or values of the input, produces irregular and unpredictable branching and memory-access patterns that are challenging to parallelize efficiently.
In contrast, $\mathit{sodt}_{\text{vec}}$ is purely sequential and can be fully vectorized, offering substantial advantages. It enables incremental evaluation of the tree objective, as feasible splits are added recursively to leaves, requiring only simple leaf splits. Nevertheless, it is less efficient than $\mathit{sodt}_{\text{rec}}$ because no DP solution exists for the objective that we concerns, and additional computations arise from feasibility checks using the ancestry relation matrix $\boldsymbol{K}$ and the $\mathit{unique}$ operation. This method is also more memory-intensive, particularly when processing data in batches. For instance, storing $M$ trees with $K$ splitting rules in a batch requires $N\times M\times\left(K+1\right)$ memory for predictions—a significant burden for large $M$ and $K$. In practice, this can quickly exceed GPU memory limits.
The $\mathit{odt}_{\text{kperms}}$ algorithm faces similar challenges to $\mathit{sodt}_{\text{vec}}$. Representing decision trees via $K$-permutations produces duplicate trees. Unlike $\mathit{sodt}_{\text{vec}}$, duplicates can sometimes be removed using a $\mathit{unique}$ operation, but permutations such as $\left[1,2,3\right]$ and $\left[1,3,2\right]$ correspond to the same tree require reconstructing the tree structure first, which incurs additional computational cost.
Overall, we believe $\mathit{sodt}_{\text{rec}}$ and $\mathit{sodt}_{\text{vec}}$ are the most suitable methods for solving the HODT problem. However, it remains unclear whether the hardware advantages of $\mathit{sodt}_{\text{vec}}$ will outweigh the computational efficiency of $\mathit{sodt}_{\text{rec}}$. We summarized advantages and limitations of $\mathit{odt}_{\text{size}}$ algorithms in Table 1.
[…]
In Part II, we implement optimized versions of $\mathit{sodt}_{\text{rec}}$ and $\mathit{sodt}_{\text{vec}}$ using more efficient array-based tree representations instead of tree structures. Experimental results align with expectations: although $\mathit{sodt}_{\text{rec}}$ is theoretically more efficient $\mathit{sodt}_{\text{vec}}$ outperforms it in both sequential and parallel settings due to better hardware compatibility. Future work may explore hardware tailored to $\mathit{sodt}_{\text{rec}}$, enabling a more balanced comparison between these approaches.
5 Geometric foundation
Having developed a comprehensive algorithmic theory for ODT problems, along with four distinct solutions for both size- and depth-constrained cases, we note that these theories rely on the assumption that the splitting rules ($\mathit{rs}$) or the method for generating these rules ($\mathit{gen}_{\mathit{splits}}$) are known. However, in machine learning, such rules are typically not readily available; they must be inferred from data. Moreover, our theory is grounded in the axioms of the proper decision tree, but not all splitting rules satisfy these axioms—for example, those used in ODT problems over binary feature data.
In Section 5.1, we present a detailed analysis of the combinatorial and geometric aspects of hypersurfaces by drawing on results from the theory of hyperplane arrangements and algebraic geometry. We show that, when data points are in general position, hypersurfaces defined by fixed-degree polynomials can be represented as combinations of data items and can be exhaustively generated in polynomial time. Furthermore, we demonstrate that such hypersurfaces satisfy the axioms of the proper decision tree.
Section 5.1 explains the construction of two splitting rule generators, that is, a concrete implementation of $\mathit{gen}_{\mathit{splits}}$ for hypersurface and mixed splitting rules. Finally, Subsection 5.3 presents a detailed combinatorial analysis of the complexity of various ODT problems.
5.1 Classification meet combinatorial geometry
Polynomial hypersurfaces, and their special case—hyperplanes—are natural candidates for the splitting rules of a proper decision tree, as they naturally satisfy Axioms 1, 2, and 3. However, for Axiom 4, a subtle question arises: why can’t the relations $h_{i}^{M}\searrow h_{j}^{M}$, $h_{i}^{M}\swarrow h_{j}^{M}$ coexist? At first glance, by Bézout’s theorem, any two polynomial hypersurfaces must intersect at some points. Intuitively, this suggests that a hypersurface could lie on both sides of another by being “cut” into different regions along their intersection points, thus both $h_{i}^{M}\searrow h_{j}^{M}$ and $h_{i}^{M}\swarrow h_{j}^{M}$ should coexist because $h_{i}$ can cut $h_{j}$ and $h_{j}$ can cut $h_{i}$ into two parts. While this is true in general, the situation is different in the context of machine learning, where we are primarily concerned with the partition of data induced by hypersurfaces rather than the hypersurfaces themselves.
By enforcing a natural assumption on the data distribution, we show that the hypersurfaces generated from data can only lie on one side of another hypersurface, ensuring that the coexistence of $h_{i}^{M}\searrow h_{j}^{M}$ and $h_{i}^{M}\swarrow h_{j}^{M}$ is impossible.
5.1.1 Polynomial hypersurface and the $W$-tuple Veronese embedding
A monomial $M$ with respect to a $D$-tuple $x=\left(x_{1},x_{2}\ldots,x_{D}\right)$ is a product of the form $M\left(x\right)=x_{1}^{\alpha_{1}}\cdot x_{2}^{\alpha_{2}}\ldots\cdot x_{D}^{\alpha_{D}},$ where $\alpha_{1},\alpha_{2}\ldots,\alpha_{D}$ are non-negative integers, denote $\boldsymbol{\alpha}=\left(\alpha_{1},\alpha_{2}\ldots,\alpha_{D}\right)$. The total degree of this monomial is the sum $\left|\boldsymbol{\alpha}\right|=\alpha_{1}+\cdot\cdot\cdot+\alpha_{n}$.
Definition 7.
Polynomial. A polynomial $P\left(x\right)$, or $P$ in short, is a finite linear combination (with coefficients in the field $\mathbb{R}$) of monomials in the form
$P\left(x\right)=\sum_{i}w_{i}x^{\boldsymbol{\alpha}_{i}},w_{i}\in\mathbb{R}.$
We denote the set of all polynomials with variables $x_{1},x_{2}\ldots,x_{D}$ and coefficients in $\mathbb{R}$ is denoted by $\mathbb{R}\left[x_{1},x_{2}\ldots,x_{D}\right]$ or $\mathbb{R}\left[x\right]$.
The maximal degree of $P$, denoted $deg\left(P\right)$, is the maximum $\left|\boldsymbol{\alpha}_{i}\right|$ such that the coefficient $\alpha_{j}$ is nonzero. For instance, the polynomial $P\left(x\right)=5x_{1}^{2}+3x_{1}x_{2}+x_{2}^{2}+x_{1}+x_{2}+3$ for $x\in\mathbb{R}^{2}$ has six terms and maximal degree two.
The number of possible monomial terms of a degree $K$ polynomial is equivalent to the number of ways of selecting $K$ variables from the multisets of $D+1$ variables [^3] There are $D+1$ variables considering polynomials in homogeneous coordinates, i.e. projective space $\mathbb{P}^{D}$.. This is equivalent to the size $K$ combinations of $D+1$ elements taken with replacement. In other words, selecting $K$ variables from the variable set $\left(x_{0},x_{1},\ldots,x_{D}\right)$ in homogeneous coordinates with repetition. This lead to the following fact.
Fact 1.
If polynomial $P$ in $\mathbb{R}\left[x_{1},x_{2}\ldots,x_{D}\right]$ has maximal degree $M$, then polynomial $P$ has $\left(\begin{array}[]{c}D+M\\ D\end{array}\right)$ monomial terms (including constant term) at most.
Given polynomial $P\in\mathbb{R}\left[x_{1},x_{2}\ldots,x_{D}\right]$ of degree $M$, a hypersurface $h^{M}$ defined by $P$ is the set of zeros of $P$ is $h^{M}=\left\{\left(x_{1},x_{2}\ldots,x_{D}\right)\in\mathbb{R}^{D}:P\left(x_{1},x_{2}\ldots,x_{D}\right)=0\right\}$. Geometrically, the zero set of a polynomial $P$ determines a subset of $\mathbb{R}^{D}$. This subset is called a surface if its dimension is less than $D$ and a hypersurface if its dimension is D1D-1. Specifically, if this subset is an affine subspace defined by a degree-one polynomial, it is referred to as a flat. A D1D-1-dimensional affine flat is known as a hyperplane.
5.1.2 0-1 loss linear classification theorem
A hyperplane arrangement is a dissection of $\mathbb{R}^{D}$ induced by a finite set of hyperplanes. At first glance, a hyperplane arrangement may appear to carry more structural information than a set of data points (a point configuration). However, an effective approach to studying geometric structures involving points and hyperplanes is to examine the transformations between these two representations. By analyzing the duality between point configurations and hyperplane arrangements, it becomes evident that both possess equally rich combinatorial structures.
Although the space of all possible hyperplanes seems to exhibit infinite combinatorial complexity—since each hyperplane is defined by a continuous normal vector $\boldsymbol{w}_{k}$—the finiteness of data imposes strong constraints. Specifically, the number of distinct data partitions induced by hyperplanes is finite, which introduces an equivalence relation among hyperplanes.
In 1965, Cover provided a precise formula for counting the number of possible partitions induced by affine hyperplanes. Given $N$ data points in general position, the number of linearly separable dichotomies is:
$\mathcal{S}_{\mathrm{Cover}}\left(N,D+1\right)=2\sum_{d=0}^{D}\left(\begin{array}[]{c}N-1\\ d\end{array}\right).$
This result implies that the optimal 0-1 loss solution obtained by exhaustively searching over Cover’s space $\mathcal{S}_{\mathrm{Cover}}$ is equivalent to the optimal solution obtained by searching the entire continuous space $\mathbb{R}^{D}$. Formally, given $N$ data points in general position, we have
$\mathit{minR}{}_{E_{\text{0-1}}}\left\{\boldsymbol{w}\mid\boldsymbol{w}\in\mathcal{S}_{\mathrm{Cover}}\left(N,D+1\right)\right\}\subseteq\mathit{minR}{}_{E_{\text{0-1}}}\left\{\boldsymbol{w}\mid\boldsymbol{w}\in\mathbb{R}^{D}\right\},$ (46)
where $\mathit{minR}{}_{E_{\text{0-1}}}$ selects one of the optimal solutions with respect to $E_{\text{0-1}}$, and $E_{\text{0-1}}\left(\boldsymbol{w},\mathit{xs}\right)=\sum_{\left(x,t\right)\in\mathcal{D}}\mathbf{1}\left[\boldsymbol{w}^{T}x\neq t\right]$, we omit dataset $\mathit{xs}$ when evaluating $\boldsymbol{w}$ in $\mathit{minR}{}_{E_{\text{0-1}}}$. Equation (46) shows that exhaustive enumeration of $\mathcal{S}_{\mathrm{Cover}}\left(N,D+1\right)$ yields an optimal solution without searching an infinitely large hypothesis space.
However, Cover’s formula only counts the number of possible partitions; it does not provide an explicit method for enumerating the equivalence classes of hyperplanes. To address this, He and Little (2023) introduced a dual transformation that maps a set of data points into a hyperplane arrangement while preserving the incidence relations between hyperplanes and data points, revealing that hyperplanes and points are isomorphic in a combinatorial sense. In particular, we have following results.
Theorem 9.
Under the general position assumption and when optimizing the 0-1 loss, for any partition induced by a hyperplane $h$ in $\mathbb{R}^{D}$ with respect to a dataset $\mathit{xs}$, there exists $D$ points in $\mathit{xs}$ such that the partition induced by hyperplanes passing through these $D$ points coincides with the partition induced by $h$
Consequently, the 0-1 loss linear classification problem can be solved by exhaustively enumerating $\left(\begin{array}[]{c}N\\ D\end{array}\right)=O\left(N^{D}\right)$ possible combinations of data points. More formally, we have the following theorem.
Theorem 10.
0-1 loss linear classification theorem. Consider a data list $\mathit{xs}:\left[\mathbb{R}^{D}\right]$ of $N$ points in general position, along with their associated labels. Then
$\mathit{minR}{}_{E_{\text{0-1}}}\left\{min_{E_{\text{0-1}}}\left(\boldsymbol{w}_{s},-\boldsymbol{w}_{s}\right)\mid s\in\mathcal{S}_{\text{kcombs}}\left(D,\mathit{xs}\right)\right\}\subseteq\mathit{minR}{}_{E_{\text{0-1}}}\left\{\boldsymbol{w}\mid\boldsymbol{w}\in\mathcal{S}_{\mathrm{Cover}}\left(N,D+1\right)\right\}\subseteq\mathit{minR}{}_{E_{\text{0-1}}}\left\{\boldsymbol{w}\mid\boldsymbol{w}\in\mathbb{R}^{D}\right\}$
and $\boldsymbol{w}_{s}\in\mathbb{R}^{D}$ denotes the normal vector determined uniquely by $D$ data points (in general position, a unique hyperplane passes through any such $D$ points, which can be computed via matrix inversion).
Proof.
See He and Little (2023). ∎
5.1.3 0-1 loss polynomial hypersurface classification theorem
Polynomial hypersurfaces Embedded data Polynomial expression Number of data required
Axis-parallel hyperplane $\left(0,\ldots,x_{i},\ldots 0\right)$ $x_{i}=a$, $\forall i\in\left[D\right]$ 1
Affine hyperplane $\left(x_{1},\ldots,x_{D},1\right)$ $w_{1}x_{1}+w_{2}x_{2}+\ldots+w_{D}x_{D}+1=0$, $D$
Hypersphere $\left(x_{1},\ldots,x_{D},x_{1}^{2},\ldots,x_{D}^{2},1\right)$ $w_{1}x_{1}+\ldots w_{D}x_{D}\ldots+w_{2D}x_{D}^{2}+1=0$ $2D$
Conic section $\left(x_{1},\ldots,x_{1}x_{D},\ldots,x_{D}^{2},1\right)$ $w_{1}x_{1}+\ldots w_{g}x_{1}x_{2}\ldots+w_{G}x_{D}^{2}+1=0$, $G=\left(\begin{array}[]{c}D+2\\ 2\end{array}\right)-1$
Arbitrary degree-$M$ polynomials $\left(x_{1},\ldots,x_{1}x_{2}^{M-1},\ldots,x_{D}^{M},1\right)$ $w_{1}x_{1}+\ldots w_{g}x_{1}x_{2}^{M-1}\ldots+w_{G}x_{D}^{M}+1=0$ $G=\left(\begin{array}[]{c}D+M\\ D\end{array}\right)-1$
Table 3: Examples of polynomial hypersurfaces with the corresponding polynomial expression, and the number data item requires to obtain the embedding.
Based on the point-hyperplane duality, equivalence relations for linear classifiers on finite sets of data were established above. However, a linear classifier is often too restrictive in practice, as many problems require more complex decision boundaries. It is natural to ask whether it is possible to extend the theory to non-linear classification. This section explore a well-known concept in algebraic geometry, the $M$-tuple Veronese embedding, which allows the generalization of the previous strategy for solving classification problem with hyperplane classifier to problems involving hypersurface classifiers, with a worst-case time complexity $O\left(N^{G}\right)$, where $G=\left(\begin{array}[]{c}D+M\\ D\end{array}\right)-1$, and $M$ is the degree of the polynomial defining the hypersurface. If both $M$ and $D$ are fixed constants, this again gives a polynomial-time algorithm for solving the 0-1 loss hypersurface classification problem.
Theorem 11.
Given variables $x_{1},\ldots x_{D}$ in space $\mathbb{R}^{D}$ (which is isomorphic to the affine space $\mathbb{R}^{D}$ when ignoring the points at infinity (Cox et al., 1997)), let $M_{0},M_{1},\ldots M_{G}$ be all monomials of degree $M$ with variables $x_{0},x_{1},\ldots x_{D}$, where $G\leq\left(\begin{array}[]{c}D+M\\ D\end{array}\right)-1$. Define an embedding $\rho_{M}:\mathbb{P}^{D+1}\to\mathbb{P}^{G}$ which sends the point $\bar{x}=\left(1,x_{1},\ldots x_{D}\right)\in\mathbb{R}^{D+1}$ to the point $\rho_{M}\left(\bar{x}\right)=\left(M_{0}\left(\bar{x}\right),M_{1}\left(\bar{x}\right),\ldots M_{G}\left(\bar{x}\right)\right)$, we refer to the data list $\rho_{M}\left(\mathit{xs}\right)=\left[\left(M_{0}\left(\bar{x}\right),M_{1}\left(\bar{x}\right),\ldots M_{G}\left(\bar{x}\right)\right)\mid x\leftarrow\mathit{xs}\right]$ as the embedded data list. The hyperplane classification over the embedded data lists $\rho_{M}\left(\mathit{xs}\right)$ is isomorphic to the polynomial hypersurface classification (defined by a degree $M$ polynomial) over the original data list $\mathit{xs}$.
Proof.
First, we need to prove the existence of the hypersurface. For any $G$ distinct points $\mathit{s}$ in $\mathbb{R}^{D+1}$, there exist a polynomial hypersurface $S$ defined by a polynomial $P$ with degree at most $M$ such that $\mathit{s}$ lies on $S$. This is because the vector space $V$ of a polynomial $P$ in $\mathbb{R}\left[x_{1},\ldots x_{D}\right]$of degree $M$ has a dimension $\left(\begin{array}[]{c}D+W\\ D\end{array}\right)$. For any data list $ds$ of size $G<\left(\begin{array}[]{c}D+M\\ D\end{array}\right)$, the linear map from $V$ to $\mathbb{R}^{G}$ must has non-trivial kernel (the linear system has more variables than equations), thus the polynomial hypersurface $S$ must exist. Next, consider a hyperplane $\boldsymbol{w}^{T}\bar{y}=0$ defined over variables $\bar{y}=\left(1,y_{1},\ldots y_{G}\right)\in\mathbb{R}^{G+1}$, we have $\boldsymbol{w}^{T}\bar{y}\geq$0 or $\boldsymbol{w}^{T}\bar{y}$¡0 implies $\boldsymbol{w}^{T}\rho_{M}\left(\bar{x}\right)\geq 0$ or $\boldsymbol{w}^{T}\rho_{M}\left(\bar{x}\right)<0$, because the closure of the real numbers under polynomial operation, thus $M\left(\bar{x}\right)\in\mathbb{R}$ if $\bar{x}\in\mathbb{R}^{D+1}$, for any monomial $M$. ∎
Theorem 11 show that the $M$-degree polynomial hypersurface classification problem over data lists in $\mathbb{R}^{D}$ is equivalent to a linear classification problem in the embedded space $\mathbb{R}^{G}$. By analogy with the linear classification result in Theorem 10, we can formalize the 0-1 Loss Polynomial Hypersurface Classification Theorem as follows.
Theorem 12.
0-1 loss polynomial hypersurface classification theorem. Consider a data list $\mathit{xs}$ of $N$ data points in $\mathbb{R}^{D}$ in general position, along with their associated labels. Denote $E_{\text{0-1}}\left(\boldsymbol{w}\right)$ as the 0-1 loss for evaluating normal vector $\boldsymbol{w}\in\mathbb{R}^{G}$ with respect to data lists $\rho_{M}\left(\mathit{xs}\right)$. Then the hyper degree $M$ polynomial
$\mathit{minR}{}_{E_{\text{0-1}}}\left\{min_{E_{\text{0-1}}}\left(\boldsymbol{w}_{s},-\boldsymbol{w}_{s}\right)\mid s\in\mathcal{S}_{\text{kcombs}}\left(G,\rho_{M}\left(\mathit{xs}\right)\right)\right\}\subseteq\mathit{minR}{}_{E_{\text{0-1}}}\left\{\boldsymbol{w}\mid\boldsymbol{w}\in\mathbb{R}^{G}\right\}$
where $\boldsymbol{w}_{s}\in\mathbb{R}^{G}$ denote as the normal vector determined by $G$ data points in $\mathcal{S}_{\text{kcombs}}\left(G,\mathit{xs}\right)$.
5.1.4 Hypersurface decision trees are proper
We now show that the hyperplane classification tree satisfies the axioms of a proper decision tree, as verified by the following lemma.
Lemma 4.
Given a set of data $\mathit{xs}$ and list of $K$ hyperplanes $\mathit{hs}_{K}=\left[h_{1},h_{2},\ldots h_{K}\right]$ defined by normal vectors $ws=\left[w_{1},w_{2},\ldots w_{K}\right]$ in $\mathbb{R}^{D+1}$, where each hyperplane passes through exactly $D$ data points. Define the positive and negative regions of each hyperplane as $h_{i}^{+}=\left\{x\mid x\in\mathbb{R}^{D},\boldsymbol{w}_{i}^{T}\bar{x}\geq 0\right\}$ and $h_{i}^{-}=\left\{x\mid x\in\mathbb{R}^{D},\boldsymbol{w}_{i}^{T}\bar{x}<0\right\}$. Then, the decision tree constructed using these hyperplanes as splitting rules is a proper decision tree.
Proof.
Axioms 1 and 2 follow directly from the definition of $h^{\pm}$. Axioms 3 and 4 hold because, for any hyperplane $h$ defined by a $D$-tuple of data points $\mathit{xs}_{D}$ we have
$\mathit{xs}_{D}\subseteq\bigcap_{p\in P}h_{p}^{\pm}\implies\mathit{xs}_{D}\subseteq h_{p}^{\pm},\forall p\in P$
and
$\mathit{xs}_{D}\subseteq h^{\pm}\implies\mathit{xs}_{D}\nsubseteq h^{\mp}.$
Thus axiom 3, 4 holds. ∎
By the equivalence between hyperplane and hypersurface, as established in Theorem 11, it follows immediately that decision trees whose splitting rules are characterized by hypersurfaces passing through $G$-combinations of data points are also proper. We next show that the optimal decision tree constructed from these hypersurfaces, with $G$ data points lying on them, is also optimal among all decision trees constructed from general hypersurfaces in $\mathbb{R}^{D}$.
Lemma 5.
0-1 classification theorem for hypersurfaces decision tree model. Consider a data list $xs$ of $N$ data points of dimension $D$ in general position, along with their associated labels. The optimal decision tree constructed using $M$-degree polynomial hypersurfaces, each passing through $G$ points, achieves the same minimum 0-1 loss as the optimal decision tree constructed using arbitrary hypersurfaces in $\mathbb{R}^{D}$.
Proof.
We prove this by induction. The base case, involving a single linear hypersurface, has already been established in Theorem 12. For the inductive step, assume two subtrees $t_{1}$ and $t_{2}$ of size $i$ and Ki1K-i-1, respectively, are optimal for the corresponding subregions $\mathit{xs}^{+}$ and $\mathit{xs}^{-}$, determined by root $h$. We show that the tree $t=DN\left(t_{1},h,t_{2}\right)$ is optimal with respect to root $h$. Since $t_{1}$ and $t_{2}$ are optimal, any subtree of the same size constructed in $\mathit{xs}^{+}$ and $\mathit{xs}^{-}$ will have a equal or greater 0-1 loss than $t_{1}$ and $t_{2}$. By definition of the objective function, we have
$E\left(t^{\prime},xs\right)=E\left(t_{1}^{\prime},xs^{+}\right)+E\left(t_{2},xs^{-}\right)\geq E\left(t,xs\right)=E\left(t_{1},xs^{+}\right)+E\left(t_{2},xs^{-}\right)$
Thus, $t=DN\left(t_{1},h,t_{2}\right)$ is optimal for the fixed root $h$. Furthermore, for any arbitrary hypersurface defined by normal vector $\boldsymbol{w}\in\mathbb{R}^{G}$, there exists a hypersurface passing through $G$ points that induces the same partition $\mathit{xs}^{+}$ and $\mathit{xs}^{-}$. Therefore, the optimal decision tree constructed using hypersurfaces through $G$ points attains the same 0-1 loss as the optimal tree constructed from arbitrary hypersurfaces. ∎
Although we assume that the data are in general position—which excludes axis-parallel hyperplanes, as they do not satisfy the general position assumption—the results also hold for axis-parallel hyperplanes, since they can be infinitesimally shifted to become hyperplanes in general position without affecting the predictions. Therefore, for axis-parallel hyperplane splitting rule, a single data point is required to characterize the hyperplane. By definition, an axis-parallel hyperplane, $x_{i}=c$ is uniquely determined by the $i$-th coordinate of any point lying on it. That is, for any point $\left(x_{1},x_{2},\ldots,x_{D}\right)\in\mathbb{R}^{D}$, the hyperplane parallel to the $i$-th axis passing through it is uniquely determined by $x_{i}$. Since each data point has $D$ coordinates, there are $N\times D$ possible axis-parallel splitting rules in total.
5.2 Splitting rule generators
5.2.1 Generating fixed polynomial degree hypersurface splitting rules
  1. 1.
    Input: $\mathit{xs}$: input data list of length $N$; $K$: Outer combination size (nested combination); $G$: inner-combination size (ordinary combination), determined by polynomial degree $M$ of hypersurface
  2. 2.
    Output: Array of $\left(K,G\right)$-nested-combinations
  3. 3.
    $\mathit{css}=\left[\left[\left[\>\right]\right],\left[\right]^{k}\right]$ // initialize combinations
  4. 4.
    $\mathit{ncss}=\left[\left[\left[\>\right]\right],\left[\right]^{k}\right]$ // initialize nested-combinations
  5. 5.
    $\mathit{asgn}^{+},\mathit{asgn}^{-}=\mathit{empty}\left(\left(\begin{array}[]{c}N\\ D\end{array}\right),N\right)$ // initialize $\mathit{asgn}^{+},\mathit{asgn}^{-}$ as two empty $\left(\begin{array}[]{c}N\\ D\end{array}\right)\times N$ matrix
  6. 6.
    for $n\leftarrow\mathit{range}\left(0,N\right)$ do: $\mathit{range}\left(0,N\right)=\left[0,1,\ldots,N-1\right]$
  7. 7.
    for $j\leftarrow\mathit{reverse}\left(\mathit{range}\left(G,n+1\right)\right)$ do:
  8. 8.
      $\mathit{updates}=\mathit{reverse}\left(\mathit{map}\left(\cup\rho_{M}\left(\mathit{xs}\right)\left[n\right],\mathit{css}\left[j-1\right]\right)\right)$
  9. 9.
      $\mathit{css}\left[j\right]=\mathit{css}\left[j\right]\cup\mathit{updaets}$ // update $\mathit{css}$ to generate combinations in revolving door ordering,
  10. 10.
    $\mathit{asgn}^{+},\mathit{asgn}^{-}=\mathit{genModels}\left(\mathit{css}\left[G\right],\mathit{asgn}^{+},\mathit{asgn}^{-}\right)$ // use $G$-combination to generate the positive prediction and negative prediction of each hyperplanes and stored in $\mathit{asgn}^{+},\mathit{asgn}^{-}$
  11. 11.
    $C_{1}=\left(\begin{array}[]{c}n\\ G\end{array}\right)$, $C_{2}=\left(\begin{array}[]{c}n+1\\ G\end{array}\right)$
  12. 12.
    for $i\leftarrow\mathit{range}\left(C_{1},C_{2}\right)$ do:
  13. 13.
       for $k\leftarrow\mathit{reverse}\left(\mathit{range}\left(K,i+1\right)\right)$ do:
  14. 14.
        $\mathit{ncss}\left[k\right]=\mathit{map}\left(\cup\left[i\right],\mathit{ncss}\left[k-1\right]\right)\cup\mathit{ncss}\left[k\right]$ // update nested combinations
  15. 15.
    $\mathit{ncss}\left[K\right]=\left[\>\right]$
  16. 16.
    return $\mathit{ncss}\left[K\right]$
Algorithm 1 $\mathit{nestedCombs}$
To solve the original decision tree problem $\mathit{odt}_{\text{size}}$ using $\mathit{sodt}$, a separate generator is required to produce possible splitting rules for a given data list $\mathit{xs}$. Specifically, we need an efficient definition for the following specification:
$\mathit{odk}_{\text{size}}\left(K\right)\circ\mathit{gen}_{\mathit{splits}}^{M}=\mathit{min}_{E_{\text{0-1}}}\circ\mathit{concatMap}_{\mathit{sodt}\left(\mathit{xs}\right)}\circ\mathit{kcombs}\left(K\right)\circ\mathit{gen}_{\mathit{splits}}^{M}$
where $\mathit{gen}_{\mathit{splits}}^{M}:\mathcal{D}\to\left[\mathcal{R}\right]$ denotes the splitting rule generator, precisely the splitting rule generator used in the definition of $\mathit{odk}_{\text{depth}}$.
For the axis-parallel decision tree problem, $\mathit{gen}_{\mathit{splits}}^{0}\left(\mathit{xs}\right)$ can be simply defined as the all possible axis-parallel hyperplanes that go through every data points in $\mathit{xs}$, so there are $N\times D$ such hyperplanes in total. For more complex splitting rules, as established in Theorems 10 and 12, both hyperplane and hypersurface decision boundaries can be exhaustively enumerated by generating $G$-combinations of data points.
Consequently, the search space of hypersurface decision boundaries can be represented as $\mathcal{S}_{\text{kcombs}}\left(G,\mathit{ds}\right)$, and the generator for this search space can be implemented directly using a combination generator:
$\mathit{gen}_{\mathit{splits}}^{M}\left(\mathit{ds}\right)=!\left(G\right)\circ\mathit{kcombs}\left(G,\mathit{ds}\right),$
where $\mathit{kcombs}\left(G,\mathit{ds}\right):\mathbb{N}\times\mathcal{D}\to\left[\left[\mathcal{D}\right]\right]$ generate all combinations with size smaller than $G$, and $\left(G\right)$ extract combinations of size exactly $G$. Hence, the composition
$\mathit{kcombs}\left(K\right)\circ\mathit{gen}_{\mathit{splits}}^{M}=\mathit{kcombs}\left(K\right)\circ!\left(G\right)\circ\mathit{kcombs}\left(G,\mathit{ds}\right),$
corresponds to a combination-of-combinations generator, i.e., nested combinations, denote as $\left(K,G\right)$-nested combinations.
Once $G$-combinations are used to generate $K$-combinations, these intermediate combinations are no longer needed and can be discarded. By introducing an additional set-empty operation, $\mathit{setEmpty}\left(G,\mathit{xs}\right)$ which sets the $G$th element of the list $\mathit{xs}$ to an empty list, we can modify the nested combination generator $\mathit{kcombs}\left(K\right)\circ\mathit{gen}_{\mathit{splits}}^{M}$ by introducing $\mathit{setEmpty}$. This modified generator, denoted as $\mathit{nestedCombs}\left(K,G\right)$, can be compositionally specified as:
$\mathit{nestedCombs}\left(K,G\right)$ $=\left\langle\mathit{setEmpty}\left(G\right)\circ\mathit{kcombs}\left(G\right),\mathit{kcombs}\left(K\right)\circ!\left(G\right)\circ\mathit{kcombs}\left(G\right)\right\rangle$ (47)
$=\left\langle\mathit{setEmpty}\left(G\right),\mathit{kcombs}\left(K\right)\circ!\left(G\right)\right\rangle\circ\mathit{kcombs}\left(G\right)$
where $\left\langle f,g\right\rangle\left(x\right)=\left(f\left(x\right),g\left(x\right)\right)$, and $G=1$ if $M=0$ and $G=\left(\begin{array}[]{c}M+D\\ D\end{array}\right)-1$ if $M\geq 1$.
However, the naive implementation 47 is inefficient due to the overhead of storing intermediate results: $\mathit{kcombs}\left(G\right)$ produces $O\left(N^{G}\right)$ combinations for an input of size $N$, which is both memory-intensive and computationally costly. Additionally, its non-recursive nature prevents the application of thinning techniques.
Nevertheless, He (2025) presents both D&C and sequential versions of $\mathit{nestedCombs}$, while the D&C version is efficient, it complicates the ordering of elements. Element ordering is critical for our applications because combinations represent hyperplanes, which carry essential predictive information. Therefore, we adopt the sequential version of $\mathit{nestedCombs}$ developed by He and Little (2025a). The pseudo-code for this sequential version is presented in Algorithm 1.
One source of efficiency in $\mathit{nestedCombs}$ is the representation of inner combinations as integers ranging from $1$ to $\left(\begin{array}[]{c}N\\ G\end{array}\right)$. This is achieved through the clever use of the $\mathit{reverse}$ function, which reverses the order of elements in a list. With this modification, combinations are arranged in revolving-door order He and Little (2025a), avoiding explicit representation of combinations. Instead, their “ranks” are used, allowing the information of each hyperplane to be indexed in $O\left(1\right)$ time using the integer representation.
5.2.2 Generating mixed polynomial degree hypersurface splitting rules
  1. 1.
    Input: $\mathit{xs}$: input data list of length $N$; $K$: Outer combination size (nested combination); $\mathit{Gs}=\left[G_{1},G_{2},\ldots G_{l}\right]$: a list of inner-combination sizes (ordinary combination), corresponds to a list of polynomial degree $\mathit{Ms}=\left[M_{1},M_{2},\ldots M_{l}\right]$
  2. 2.
    Output: List of $\left(K,Gs\right)$-nested-combinations arrays
  3. 3.
    $\mathit{css}=\left[\left[\left[\>\right]\right],\left[\right]^{k}\right]$ // initialize combinations
  4. 4.
    $\mathit{ncss}=\left[\left[\left[\>\right]\right],\left[\right]^{k}\right]$ // initialize nested-combinations
  5. 5.
    $\mathit{asgn}^{+},\mathit{asgn}^{-}=\mathit{empty}\left(\left(\begin{array}[]{c}N\\ G_{1}\end{array}\right)+\left(\begin{array}[]{c}N\\ G_{2}\end{array}\right)+\ldots+\left(\begin{array}[]{c}N\\ G_{l}\end{array}\right),N\right)$ // initialize $\mathit{asgn}^{+},\mathit{asgn}^{-}$ as two empty $\left(\begin{array}[]{c}N\\ D\end{array}\right)\times N$ matrix
  6. 6.
    for $n\leftarrow\mathit{range}\left(0,N\right)$ do: $\mathit{range}\left(0,N\right)=\left[0,1,\ldots,N-1\right]$
  7. 7.
    $\mathit{candInds}=\left[\>\right]$
  8. 8.
    for $j\leftarrow\mathit{reverse}\left(\mathit{range}\left(G_{l},n+1\right)\right)$ do:
  9. 9.
      $\mathit{updates}=\mathit{reverse}\left(\mathit{map}\left(\cup\rho_{M}\left(\mathit{xs}\right)\left[n\right],\mathit{css}\left[j-1\right]\right)\right)$
  10. 10.
      $\mathit{css}\left[j\right]=\mathit{css}\left[j\right]\cup\mathit{reverse}\left(\mathit{map}\left(\cup\rho_{M}\left[n\right],\mathit{updates}\right)\right)$ // update $\mathit{css}$ to generate combinations in revolving door ordering,
  11. 11.
      if $G\in\mathit{Gs}$ do:
  12. 12.
       $\mathit{asgn}^{+},\mathit{asgn}^{-}=\mathit{genModels}\left(\mathit{updates},\mathit{asgn}^{+},\mathit{asgn}^{-}\right)$ // use $G$-combination to generate the positive prediction and negative prediction of each hyperplanes
  13. 13.
       $C_{1}=\left(\begin{array}[]{c}G\\ D\end{array}\right)$, $C_{2}=\left(\begin{array}[]{c}G+1\\ D\end{array}\right)$
  14. 14.
       for $G^{\prime}\leftarrow\mathit{Gs}$ and $G^{\prime}<G$ do:
  15. 15.
        $C_{1}=C_{1}+\left(\begin{array}[]{c}N\\ G^{\prime}\end{array}\right)$, $C_{2}=C_{2}+\left(\begin{array}[]{c}N\\ G^{\prime}\end{array}\right)$
  16. 16.
       $\mathit{candInds}=\mathit{candInds}+\!\!+\mathit{range}\left(C_{1},C_{2}\right)$
  17. 17.
    for $i\leftarrow\mathit{candInds}$ do:
  18. 18.
       for $k\leftarrow\mathit{reverse}\left(\mathit{range}\left(K,i+1\right)\right)$ do:
  19. 19.
        $\mathit{ncss}\left[k\right]=\mathit{map}\left(\cup\left[i\right],\mathit{ncss}\left[k-1\right]\right)\cup\mathit{ncss}\left[k\right]$ // update nested combinations
  20. 20.
    $\mathit{ncss}\left[K\right]=\left[\>\right]$
  21. 21.
    return $\mathit{ncss}\left[K\right]$
Algorithm 2 $\mathit{nestedCombs}^{\prime}$
One advantage of constructing algorithms based on a generic framework is the flexibility it provides. In theory, since the input to $\mathit{odt}_{\text{size}}$ is a list of rules rather than raw data, arbitrary splitting rules satisfying the decision tree axioms can be supplied. This allows the construction of decision trees with mixed splitting rules. The main challenge lies in generating these splitting rules efficiently from a given data list. Fortunately, axis-parallel hyperplanes, general hyperplanes, and hypersurfaces can all be represented as combinations of data items.
A key observation about Algorithm 1 is that, it can generate not only $G$-combinations but all $G^{\prime}$-combinations for $0\leq G^{\prime}\leq G$. However, the nested combination generator $\mathit{nestedCombs}\left(K,G\right)$ uses only the $G$-combinations and ignores smaller $G^{\prime}$-combination ($G^{\prime}<G$). To accomodate this, we need to modify specification 47 to the following
$\mathit{nestedCombs}^{\prime}\left(K,\mathit{Gs}\right)=\left\langle\mathit{setEmpty}\left(G_{l}\right),\mathit{kcombs}\left(K\right)\cdot!!\left(\mathit{Gs}\right)\right\rangle\cdot\mathit{kcombs}\left(G\right),$ (48)
where $\left(\mathit{Gs},\mathit{xss}\right)=\mathit{concat}\left[!\left(G,\mathit{xss}\right)\mid G\leftarrow\mathit{Gs}\right]$ and $\mathit{Gs}=\left[G_{1},G_{2},\ldots G_{l}\right]$ is an ascending list of integers ($G_{1}<G_{2}\ldots<G_{l}$), with $G_{l}$ being the largest, corresponds to a list of polynomial degree $\mathit{Ms}=\left[M_{1},M_{2},\ldots M_{l}\right]$. Specification 48 selects several $G$-combinations, such that $G\in\mathit{Gs}$, and constructs $K$-nested combinations from them.
As usual, the naive specification 48 is inefficient. He and Little (2025a) provide an efficient solution by defining the recursive pattern of $\mathit{nestedCombs}^{\prime}\left(K,\mathit{Gs}\right)$ as
$\bigg{<}\mathit{setEmpty}\left(G_{l}\right)\cdot\mathit{KcombsAlg}\left(K\right)\cdot\mathit{Ffst},$
$\quad\quad\mathit{\mathit{\mathit{KcombsAlg}}\left(K\right)}\cdot\left\langle\mathit{Kcombs}\left(K\right)\cdot!!\left(\mathit{Gs}\right)\cdot convol_{\text{new}}\left(\cup,\circ\right)\cdot\mathit{Ffst},\mathit{KcombsAlg}\left(K\right)\cdot\mathit{Fsnd}\right\rangle\bigg{>}$
where the detailed definitions for $\mathit{Fst}$, $\mathit{Fsnd}$, $\mathit{setEmpty}$, $\mathit{KcombsAlg}$ and $\mathit{setEmpty}$ are provided in He and Little (2025a) and are omitted here due to space constraints. Specifically, $convol_{\text{new}}$ is defined as
$convol_{\text{new}}\left(\cup,\circ,\mathit{css}_{1},\mathit{css}_{2}\right)=\left[\mathit{concat}\left(\mathit{zipWith}\left(\circ,\mathit{init}\left(\mathit{tail}\left(\mathinner{css}\right)\right),\mathit{init}\left(\mathit{tail}\left(\mathit{css}_{2}\right)\right)\right)\right)\mid\mathinner{css}\leftarrow\mathit{inits}\left(\mathit{css}_{1}\right)\right].$ (49)
The sequential definition of $\mathit{nestedCombs}^{\prime}$ is provided in Algorithm 2. Unlike the original $\mathit{nestedCombs}$, which represents inner combinations using integers ranging from $1$ to $\left(\begin{array}[]{c}N\\ G\end{array}\right)$. $\mathit{nestedCombs}^{\prime}$must account for chains of $G_{1}$-$G_{2}$…-ombinations connected consecutively. For a $G$-combinations with rank $r$, its index in the algorithm is $i=r+\mathit{sum}\left[\left(\begin{array}[]{c}N\\ G^{\prime}\end{array}\right)\mid G\in\mathit{Gs},G^{\prime}\leq G\right]$, which corresponds to the calculation in lines 13–16 of the algorithm.
5.3 Complexity of solving optimal decision tree problems
5.3.1 Complexity of size-constrained trees
We present a theoretical analysis of the computational complexity of size-constrained decision trees. The complexity of depth-constrained trees can be upper-bounded by noting that a tree of depth $d$ has at most $2^{d+1}-1$ splitting rules.
Complexity of the hyperplane decision tree problem
The computational complexity of the optimal hypersurface decision tree problem is relatively straightforward to analyze. There are $O\left(N^{GK}\right)$ possible combinations of degree $M$ hypersurfaces. The subprogram $\mathit{sodt}$ has a worst-case complexity of $O\left(K!\right)$. Thus, solving the HODT problem requires
$O\left(N^{G}\times D^{3}\times N+K!\times N\times N^{GK}\right)=O\left(N^{GK+1}\right).$ (50)
where $O\left(N^{G}\times D^{3}\times N\right)$ represents the time required to compute hyperplane predictions. Although this complexity appears daunting, Part II demonstrates that the actual complexity of the hypersurface decision tree problem is significantly lower.
Furthermore, the average-case complexity of $\mathit{sodt}_{\text{rec}}$ is substantially lower than its theoretical upper bound. The worst-case complexity of $\mathit{sodt}_{{}^{\prime}\text{rec}}$ occurs only when the tree forms a “chain” structure, which, as shown in the empirical analysis in Part II, is a rare occurrence.
Complexity of the axis-parallel decision tree problem
For the axis-parallel decision tree (AODT) problem, there are $O\left(N\times D\right)$ possible splits defined by axis-parallel hyperplanes. Consequently, the complexity of the AODT problem is
$O\left(N\times D+K!\times N\times\left(N\times D\right)^{K}\right)=O\left(N^{K}\right).$ (51)
Notably, when $K$ splitting rules are fixed, both hypersurface and axis-parallel decision trees have a worst-case complexity bounded by $O\left(K!\right)$. However, the number of feasible trees in the hypersurface case is significantly smaller than in the axis-parallel case. This difference arises because, in axis-parallel decision trees, every pair of hyperplanes is mutually ancestral (a pair of hyperplanes is mutually ancestral if both $h_{i}\left(\swarrow\vee\searrow\right)h_{j}$ and $h_{j}\left(\swarrow\vee\searrow\right)h_{i}$ are viable, a case that rarely occurs in hypersurface decision trees).
Complexity of the decision tree problem over binary feature data
The complexity of decision trees with binary feature data differs significantly. Due to axiom 4, if $h_{1}\swarrow h_{2}$ or $h_{1}\searrow h_{2}$, then $h_{1}\searrow h_{2}$ or $h_{1}\swarrow h_{2}$ is impossible, i.e. if $h_{2}$ is in the left branch of $h_{1}$, it cannot lie in the right branch of it.
In contrast, for binary feature data, hyperplanes represent the selection or non-selection of features, allowing any pair of branch nodes to be mutually ancestral and symmetric. Symmetry implies that any branch node can serve as either the left or right child of the root. For example, when selecting features $x_{i}$ and $x_{j}$, both $x_{i}\swarrow x_{j}$ and $x_{i}\searrow x_{j}$ are possible when $x_{i}$ is the root, and $x_{j}\swarrow x_{i}$ and $x_{j}\searrow x_{i}$ are possible with $x_{j}$ as root. This holds because, in all previous cases, hyperplanes are characterized by data points, and similarly, in the axis-parallel tree case, only one data item is needed to define each hyperplane, and this item always lies on one side of another hyperplane. However, for binary feature data, this method of characterization splitting rules as data points are no longer applicable. In the case of binary feature data, it is only necessary to consider whether a particular feature is selected or not and there are $K$ features in total used to define decision trees of size $K$. Hence, for binary feature data, the number of possible decision trees depends not only on the shape of the tree (which can be counted using the Catalan number) but also the labels.
Therefore, assume $D$ is much larger than $K$, the worst-case complexity for the decision tree problem with binary feature data is
$O\left(\left(\begin{array}[]{c}D\\ K\end{array}\right)+K!\times\mathrm{Catalan}\left(K\right)\times\left(\begin{array}[]{c}D\\ K\end{array}\right)\right)=O\left(D^{K}\right),$ (52)
where $\mathrm{Catalan}\left(n\right)$ is the Catalan number of $n$.
5.3.2 Complexity of depth-constrained trees
The combinatorial complexity of the depth-constrained ODT problem is challenging to compute precisely. A key observation is that, for arbitrary decision tree of depth-$d$ decision tree it should contains at least $d$ internal nodes (splitting rules). Consequently, the search space of depth-$d$ decision trees with respect to data list $\mathit{xs}$, denoted $\mathcal{S}_{\text{depth}}\left(d,\mathit{xs}\right)$, is at least as large as the search space of decision trees with $d$ splitting rules, $\mathcal{S}_{\text{size}}\left(d,\mathit{xs}\right)$. In other words, we have
$\mathcal{S}_{\text{depth}}\left(d,\mathit{xs}\right)\geq\mathcal{S}_{\text{size}}\left(d,\mathit{unique}\left(\mathit{gen}_{\mathit{splits}}\left(\mathit{xs}\right)\right)\right).$ (53)
where $\mathit{unique}\left(\mathit{gen}_{\mathit{splits}}\left(\mathit{xs}\right)\right)$ denote the all possible unique splitting rules generated by $\mathit{gen}_{\mathit{splits}}\left(\mathit{xs}\right)$, as there might have duplicate splitting rules (for categorical data, distinct data points may have same feature values).
For the axis-parallel decision tree problem, any pair of splitting rules $r_{i}$ and $r_{j}$ defined by two data points $x_{i}$ and $x_{j}$, satisfies both $r_{i}\swarrow r_{j}$ and $r_{i}\searrow r_{j}$. Consequently, by Axiom 4, the relation $r_{i}\overline{\left(\swarrow\vee\searrow\right)}r_{j}$ does not hold for any pair of splitting rules. This implies that any $d$-combination of splitting rules can construct at least one valid decision tree, thus
$\mathcal{S}_{\text{size}}\left(d,\mathit{unique}\left(\mathit{gen}_{\mathit{splits}}\left(\mathit{xs}\right)\right)\right)\geq\mathcal{S}_{\text{kcombs}}\left(d,\mathit{unique}\left(\mathit{gen}_{\mathit{splits}}\left(\mathit{xs}\right)\right)\right).$ (54)
Therefore, the following inequality holds:
$\mathcal{S}_{\text{depth}}\left(d,\mathit{xs}\right)\geq\mathcal{S}_{\text{size}}\left(d,\mathit{unique}\left(\mathit{gen}_{\mathit{splits}}\left(\mathit{xs}\right)\right)\right)\geq\mathcal{S}_{\text{kcombs}}\left(d,\mathit{unique}\left(\mathit{gen}_{\mathit{splits}}\left(\mathit{xs}\right)\right)\right)=\left(\begin{array}[]{c}\left|\mathit{unique}\left(\mathit{gen}_{\mathit{splits}}\left(\mathit{xs}\right)\right)\right|\\ d\end{array}\right)$ (55)
As discussed, for axis parallel decision tree problem, $\left|\mathit{gen}_{\mathit{splits}}\left(\mathit{xs}\right)\right|$ is the sum of all distinct features with respect to the training data list $\mathit{xs}$, which is at most $N\times D$ for a size $N$ data list in $\mathbb{R}^{D}$.
Remarks on Quant-BnB algorithm
In the original source code of Mazumder et al. (2022)’s Quant-BnB algorithm, there is a function count_remain_3D, which computes the combinatorial complexity of the search space, i.e., the number of remaining trees in each iteration. The initial call to count_remain_3D, calculates this complexity directly from the input data without bounding parameters. For almost all datasets we have tested, count_remain_3D computed results that is strictly smaller than $\left(\begin{array}[]{c}\left|\mathit{gen}_{\mathit{splits}}\left(\mathit{xs}\right)\right|\\ d\end{array}\right)$. Interested readers are encouraged to test this result on their own datasets.
As explained above, this should be impossible because every $d$-combination of splitting rule should construct at least one possible proper decision tree. We hypothesize that Mazumder et al. (2022) either employed an incorrect algorithmic procedure for solving the AODT problem or addressed a specialized ODT problem. However, their ambiguous problem definition hinders verification of their claims. Consequently, there is an urgent need for Mazumder et al. (2022) to clarify their counting formula to validate the correctness of their algorithms.
6 Conclusion
The primary distinction between our research and other studies on ODT algorithms using combinatorial methods lies in philosophy: we argue that deriving optimal algorithms requires the highest level of rigor and clarity. Consequently, in this first of two papers, we introduce a generic algorithmic framework for deriving optimal algorithms directly from problem specifications. We begin by establishing a formal standard for designing optimal algorithms, identifying several key principles that highlight the level of rigor required for problem definitions in the study of optimal algorithms. These principles resemble the notion of a “standard form” in the study of MIP solvers. Importantly, when a problem specification satisfies these principles, proving optimality becomes straightforward. This is analogous to general-purpose MIP solvers: once a problem is expressed in a standardized form, the solver can compute the solution directly, while the challenging part lies in producing the appropriate specification.
To demonstrate the importance and effectiveness of these standards, we present a concrete example by outlining the algorithmic and geometric foundations for solving the ODT problem. In the algorithmic section, we provide four generic and formal definitions of optimal decision tree problems using the “proper decision tree” datatype, a recursive datatype with four associated axioms. Building on this, we develop four formal definitions of the ODT problem over arbitrary splitting rules and objective functions within a specified form. Furthermore, these formulations yield four solutions to the ODT problem: we prove the existence of a dynamic programming solution for two and the non-existence of such a solution for the remaining two. To the best of our knowledge, these definitions and their corresponding solutions are both generic and novel.
Note that in this part, we have not yet presented a concrete implementation of any algorithm for solving optimal decision tree problems. Although most algorithms presented in this part are ready for implementation, a critical assumption remains unaddressed—the generation of the ancestry relation matrix $\boldsymbol{K}$ for a given $K$ splitting rules. In the discussion of $\mathit{sodt}_{\text{rec}}$, we assume $\boldsymbol{K}$ is pre-stored in memory; however, generating it actually requires an additional computational process. However, the construction of the ancestry relation matrix is problem-dependent, which is why we have not yet defined it, as our goal is to provide a generic solution rather than algorithms for special cases.
Indeed, in the second paper (Part II) of this series, one of the central focuses is to develop an incremental algorithm for efficiently constructing the ancestry relation matrix for hypersurface splitting rules, which lays the foundation for solving the hypersurface optimal decision tree problem. Building on the algorithmic and geometric foundations established here, Part II will present the first two algorithms designed to address optimal hypersurface decision tree problems.
Specifically, we focus on two algorithms, $\mathit{odt}_{\text{vec}}$ and $\mathit{odt}_{\text{rec}}$—the former offering superior hardware compatibility, while the latter achieving stronger theoretical efficiency. We provide comprehensive experimental results analyzing the computational efficiency of $\mathit{odt}_{\text{vec}}$ and $\mathit{odt}_{\text{rec}}$. Additionally, we conduct a thorough assessment of the generalization capabilities of hypersurface decision tree models and axis-parallel decision tree algorithms, including those learned by CART and Brița et al. (2025)’s ConTree algorithm, using both synthetic and real-world datasets.

### v2

@@
 In 2017, Bertsimas and Dunn proposed the first optimal decision tree (ODT) algorithm using general-purpose solvers, sparking significant interest in subsequent studies.
-However, existing combinatorial methods—primarily based on branch-and-bound (BnB) method—focus on much simpler variants, such as trees with binary feature data or axis-parallel hyperplanes.
-Although the first combinatorial algorithm appeared in 1989, little subsequent work has directly built upon it.Researchers have proposed various algorithms with different data structures and frameworks, but few adopt a unified framework comparable to general-purpose solvers.
-Many so-called optimal algorithms rely on ad-hoc processes and bounding strategies, with complex pseudo-code reducing accessibility and ambiguously stated claims hindering independent validation.
-As a result, prior BnB algorithms often lack the rigor to prove optimality, and algorithms claiming to solve the same problem may yield inconsistent—or even incorrect—solutions.
-Ultimately, optimal algorithms demand rigorous scrutiny: they must provably guarantee the optimal solution for a given search space and objective.
+However, existing combinatorial methods—primarily based on branch-and-bound method—focus on much simpler variants, such as trees with binary feature data or axis-parallel hyperplanes.
+Although the first combinatorial algorithm appeared in 1989, little subsequent work has directly built upon it.
+Researchers have proposed various algorithms with different data structures and frameworks, but few adopt a unified framework comparable to general-purpose solvers.
+Many so-called optimal algorithms rely on ad-hoc algorithmic processes and bounding strategies, with complex pseudo-code reducing accessibility and ambiguously stated claims hindering independent validation.
+However, optimal algorithms require rigorous verification—they must provably guarantee optimal solutions for a given search space and objective.
+Previous BnB algorithms often fail to meet this level of rigor, rendering claims of fully reliable optimality, at best, premature.
 We argue that the root cause lies in the absence of a formal framework for designing optimal algorithms using combinatorial methods and standardized criteria for verifying correctness.
 Any algorithm claiming optimality must be formally proven.
-To address this, in the first paper (Part I) of this two-part series, we examined the ambiguity and informality in studies of ODT problems and discussed the consequences of relying on informal formalisms in optimal algorithm design.
+To address this, in the first paper (part I) of this series of two, we examined the ambiguity and informality in studies of ODT problems and discussed the consequences of relying on informal formalisms in optimal algorithm design.
 We then presented criteria that an optimal algorithm should satisfy to ensure a verifiable and transparent design process.
 Beyond identifying limitations, our primary goal—and the central contribution of this work—is to develop a foundational theory for solving ODT problems.
-In Part I, we introduce four novel definitions of the ODT problems: three for size-constrained trees and one for depth-constrained trees.
+We aim to establish a comprehensive algorithmic framework for all optimal decision tree problems studied in machine learning, encompassing axis-parallel, hyperplane, hypersurface decision trees, and those defined over binary feature data.
+In Part I, we introduce five rigorous formulations of ODT problems: three for size-constrained trees and one for depth-constrained trees, and one for trees over binary feature data.
 These definitions are stated unambiguously through executable recursive programs, satisfying all criteria we propose for a formal specification.
@@
 Grounded in algebraic programming theory—a relational formalism for deriving correct-by-construction algorithms from specifications—we can not only establish the existence or nonexistence of dynamic programming solutions but also derive them constructively whenever they exist.
-Consequently, the four generic problem definitions yield four novel optimal algorithms for ODT problems with arbitrary splitting rules satisfies the axioms and objective functions of a given form.
-These algorithms encompass the known depth-constrained, axis-parallel ODT algorithm as special case, while providing a unified, efficient, and elegant solution for the general ODT problem.
-In Part II, we present the first optimal hypersurface decision tree algorithm and provide comprehensive experiments against axis-parallel decision tree algorithms, including heuristic CART and state-of-the-art optimal methods.
-The results demonstrate the significant potential of decision trees with flexible splitting rules.
-Moreover, our framework is readily extendable to support algorithms for constructing even more flexible decision trees, including those with mixed splitting rules.
+Consequently, the five generic problem definitions yield five optimal algorithms for ODT problems with arbitrary splitting rules satisfies the axioms and objective functions of a given form.
+These algorithms encompass the known ODT algorithms as special case, while providing a unified, efficient, and elegant solution for the general ODT problem.
+Keywords: decision tree, dynamic programming, global optimal algorithm
 1 Introduction
@@
 For example, a generalization of the classical decision tree problem is the hyperplane (or oblique) decision tree, which employs hyperplane decision boundaries to potentially simplify complex boundary structures.
-However, the optimization problems in both approaches are computationally challenging.
-For the optimal axis-parallel decision tree (AODT) problem, Laurent and Rivest (1976) showed that determining whether there exists an axis-parallel decision tree with cost less than or equal to $w$ for a given finite dataset and rule set is NP-complete.
+However, the decision tree optimization is notoriously difficult.
+Laurent and Rivest (1976) showed that finding the smallest decision tree (defined by the path lengths) is NP-complete.
 Consequently, finding an optimal axis-parallel decision tree for a given dataset is NP-hard.
 Similarly, for hyperplane decision trees, even the greedy top-down optimization approach is NP-hard, as the number of possible hyperplane splits for a dataset of size $N$ in dimension $D$ is $O\left(N^{D}\right)$, which grows exponentially with $D$.
-Nevertheless, the stopping criteria of the CART algorithm recursively split a leaf until one of the following conditions is met: (1) the maximum tree depth is reached, (2) the maximum number of leaves is attained, (3) the minimum number of observations per leaf is violated, or (4) the minimum impurity reduction threshold is satisfied.
-Upon closer examination, the optimization problem that the CART algorithm attempts to solve differs from the one defined by Laurent and Rivest (1976), as the latter does not impose the constraints specified in CART.
-This raises an important question: is the optimization problem solved by CART also NP-hard?
-Interestingly, Ordyniak and Szeider (2021) showed that the decision tree problem with fixed size and depth is fixed-parameter tractable.
-In other words, if the size or depth of the tree is fixed, there exists a polynomial-time algorithm for solving the problem.
-While the result of Ordyniak and Szeider (2021) leaves certain aspects underspecified—for example, some key definitions regarding the problem formulation—it nevertheless offers a valuable perspective on the tractability of decision tree optimization.
-As a result, the fixed-parameter tractability of ODT problems has led to a significant growth of research interest in recent years.
-One major line of work focuses on formulating the ODT problem as a mixed-integer programming (MIP) problem, which can then be solved using general-purpose optimization solvers.
-Following the first MIP formulation proposed by Bertsimas and Dunn (2017), a surge of subsequent studies has investigated this approach (Boutilier et al., 2023; Verwer and Zhang, 2017, 2019; Boutilier et al., 2022; Zhu et al., 2020) leveraging general-purpose MIP solvers.
-Notably, Bertsimas and Dunn (2017)’s algorithm can not only solve the AODT problem but also the optimal hyperplane decision tree problem.
-An alternative line of research, which we refer to as combinatorial methods, seeks to employ approaches such as dynamic programming (DP), greedy strategies, or branch-and-bound (BnB) techniques to exhaustively explore the solution space while eliminating non-optimal solutions through bounding techniques.
-Unlike the advances made in the study of general-purpose solvers, where many practical solutions have been developed for both the AODT problem and the hyperplane decision tree problem.
-Although the first combinatorial algorithm for decision tree optimization was introduced as early as 1989 (Cox Jr et al., 1989), most subsequent studies have focused on simpler settings, such as the ODT problem over binary feature data.
-Only a few studies, such as Brița et al.
-(2025); Mazumder et al.
-(2022), have addressed the AODT problem, whereas algorithms for the hyperplane optimal decision tree problem remain unexplored.
-Unlike the work of Bertsimas and Dunn (2017) on the ODT problem, which has inspired a substantial body of subsequent research, many later studies on combinatorial algorithms adopt distinct problem definitions.
-For instance, the approaches of Demirović et al.
-(2022); Verwer and Zhang (2019); Nijssen and Fromont (2007); Aglin et al.
-(2020); Nijssen and Fromont (2010) construct optimal decision trees by constraining the tree depth.
-In contrast, Zhang et al.
-(2023); Lin et al.
-(2020); Hu et al.
-(2019) propose algorithms that constrain the number of leaves, referring to the resulting models as sparse decision trees.
-Additionally, the combinatorial methods developed to solve the ODT problem are highly diverse.
-For example, Nijssen and Fromont (2007); Aglin et al.
-(2021, 2020); Nijssen and Fromont (2010) characterize decision trees over binary feature data using the concept of itemsets from data, Demirović et al.
-(2022) represent decision trees via feature vectors, whereas Lin et al.
-(2020); Hu et al.
-(2019) define trees in terms of two sets of leaves: those that can be further split and those that are fixed.
-Unfortunately, the advantages and limitations of these data structures remain insufficiently studied.
-If these distinct representations of decision trees prove to be indispensable—playing central roles in their respective algorithms and serving different purposes in tree construction—their strengths and weaknesses warrant careful analysis.
-At the same time, while using general-purpose MIP solvers for the ODT problem is appealing and can be efficient in many cases, the overall computational complexity remains unpredictable.
-More importantly, MIP-based approaches provide limited insight into the combinatorial and geometric structure of the problem.
-Their performance largely depends on advances in general-purpose solvers and, potentially, more sophisticated MIP formulations.
-To advance theoretical understanding and develop practical algorithms, it is therefore essential to study combinatorial methods, which provide a more direct view of the underlying geometry and algorithmic structure.
-In this study, we focus on combinatorial methods for solving the ODT problem, following the axiomatic definition of decision trees established in our previous work He and Little (2025b), which we refer to as proper decision trees, we provide a detailed analysis of the algorithmic, combinatorial, and geometric properties of the ODT problems.
-Our generic solution not only offers several new perspectives on decision tree optimization but also provides a rigorous framework for future analysis.
-Building on the algorithmic and geometric results discussed in this paper, another key contribution of this work is demonstrating that the hypersurface decision tree (HODT) model satisfies the axioms we proposed and can therefore be solved using the four ODT algorithms introduced here.
-Moreover, our algorithms are readily applicable to more flexible decision tree models that incorporate mixed splitting rules.
-In Part II, we will present the detailed construction of the first optimal hypersurface decision tree (HODT) algorithms, highlighting their practical implementation and applicability.
-We believe this research lays a theoretical foundation for the study of ODT problems. A summary of our contributions is as follows:
-Achievement in Part I
-  • Critique and proposed solutions: We critically examine current approaches to designing algorithms and defining problems in the study of optimal decision trees, highlighting the consequences of relying on informal problem formulations paired with ad-hoc algorithmic solutions—a practice common in combinatorial studies of ODT problems, particularly those employing BnB methods.
+However, instead of addressing the general decision tree problem that seeks the smallest tree, successful algorithms such as CART and C4.5 were not designed with this objective.
+Rather, their stopping criteria explicitly constrain either the tree size (i.e., the number of branch nodes) or the tree depth.
+Interestingly, Ordyniak and Szeider (2021) demonstrated that the decision tree problem with fixed size or depth is fixed-parameter tractable.
+In other words, when either parameter is fixed, the problem can be solved in polynomial time.
+In this first paper of a two-part series, we present a framework for constructing efficient algorithms—polynomial in the worst case and trivially parallelizable—for obtaining globally optimal solutions to decision tree optimization problems.
+Our exposition offers solutions to nearly all ODT problems encountered in practice, including classical axis-parallel hyperplane, general hyperplane, polynomial hypersurface, and decision trees over binary feature data.
+In Part II (He, 2025a), we will detail the construction of the first optimal hypersurface decision tree algorithms, emphasizing their practical implementation and applicability.
+As our goal is to provide a comprehensive algorithmic framework for all decision tree problems relevant in practice, this paper is written in a pedagogical style.
+Nonetheless, it includes several novel contributions.
+A summary of our main contributions in Part I is as follows:
+  • Critique and proposed solutions: In Section 2.1, We critically examine current approaches to designing algorithms and defining problems in the study of optimal decision trees, highlighting the consequences of relying on informal problem formulations paired with ad-hoc algorithmic solutions—a practice common in combinatorial studies of ODT problems, particularly those employing BnB methods.
 Beyond this critique, the central distinction of our research lies in philosophy: we argue that deriving truly optimal algorithms demands the highest levels of rigor and clarity.
@@
 These discussions provide a foundational step for advancing research on ODT problems using combinatorial methods.
-  • Formal combinatorial definitions: We provide four formal combinatorial definitions for the ODT problem—covering both size- and depth-constrained trees—ensuring unambiguous communication for future research.
-All definitions are directly executable, allowing verification that the generated trees align with expectations.
-To our knowledge, no prior studies on using BnB algorithms for ODT have provided a formal problem definition with this level of rigor.
-  • Generic optimal decision tree algorithms: We present four novel and generic algorithms for solving ODT problems with arbitrary splitting rules that satisfy our axioms, capable of optimizing arbitrary objective functions of a specified form.
-All algorithms are constructively derived in a formal manner.
-Specifically, we provide three algorithms for size-constrained trees and one for depth-constrained trees, formally proving and extending results from previous studies on axis-parallel ODT problems Brița et al.
-(2025).
-  • Filter and thinning fusion: We propose two fusion theorems—filtering and thinning—to enhance the flexibility and practical applicability of ODT algorithms while maintaining formal rigor.
+  • Formal combinatorial definitions: Following the proper decision tree framework examined in He and Little (2025b), which defines decision trees through recursive datatypes and establishes four axioms, In Subsection 4.2, we established four formal combinatorial definitions for the ODT problem—covering both size- and depth-constrained trees.
+Importantly, we rigorously defined the search space and the objective function of the ODT problems as directly executable programs.
+This is vitally important because it allows researchers to inspect the search space by running the program and verifying the resulting decision trees.
+Such transparency not only validates our solution but also fosters unambiguous communication.
+To our knowledge, no prior work on BnB algorithms for ODT problems has provided a problem definition with this level of rigor, combining axiomatic definitions of decision trees with programmatic definitions of the search space and objective function.
+  • Generic optimal decision tree algorithms: In Subsection 4.3, We present four generic algorithms for solving ODT problems with arbitrary splitting rules over proper decision trees, capable of optimizing arbitrary objective functions of a specified form.
+One algorithm was previously introduced in our earlier work He and Little (2025b), while the remaining three are new contributions in this exposition.
+All algorithms are derived constructively and presented in a formal manner.
+Specifically, we provide three algorithms for size-constrained trees and one for depth-constrained trees.
+Notably, our depth-constrained ODT algorithm not only offers a formal proof of the results in Brita et al.
+(2025) for the axis-parallel ODT problem but also generalizes their findings to broader settings.
+  • Extension to decision tree over binary feature data: He and Little (2025b) showed that the ODT-BF problem is a non-proper decision tree problem, requiring a distinct formulation of its search space.
+Based on this formulation, we derived a provably optimal algorithm for solving it.
+To the best of our knowledge, this is the first optimal algorithm that guarantees exhaustive exploration of the ODT-BF search space, thereby establishing a solid foundation for addressing this problem.
+  • Filter and thinning fusion: In Subsection 4.5, We propose two fusion theorems—
+    filtering and thinning—to enhance the flexibility and practical applicability of ODT algorithms while maintaining formal rigor.
 The filtering fusion theorem allows additional constraints, such as depth and leaf-size constraints, to be incorporated into ODT algorithms defined over size-constrained trees.
@@
 Together, these theorems extend the practical utility of the algorithms without compromising optimality.
-  • Proving the properness of hypersurface decision trees: By generalizing Xi’s 0-1 loss linear/hypersurface classification theorem from linear models to decision tree models, and combining it with the geometric characterization of polynomial hypersurfaces as combinations of data points, we demonstrate that hypersurface decision trees are proper decision trees and can therefore be solved exactly using our proposed algorithms.
-  • Splitting rule generators: Leveraging the sophisticated combination generator introduced by He and Little (2025a), we develop efficient algorithms for generating both hypersurface splitting rules and mixed-splitting rules for a given dataset.
+  • Proving the properness of hypersurface decision trees: In Section 5, by generalizing He and Little (2023)’s 0-1 loss linear/hypersurface classification theorem from linear models to decision tree models, and combining it with the geometric characterization of polynomial hypersurfaces as combinations of data points, we demonstrate that hypersurface decision trees are proper decision trees and can therefore be solved exactly using our proposed algorithms.
+  • Splitting rule generators: In Subsection 5.2, leveraging the sophisticated combination generator introduced by He and Little (2025a), we develop efficient algorithms for generating both hypersurface splitting rules and mixed-splitting rules for a given dataset.
 The hypersurface splitting rule generator provides the foundation for constructing efficient hypersurface decision tree algorithms developed in Part II, while the mixed-splitting rule generator establish a foundation for future work on constructing optimal ODT algorithms with mixed splitting rules.
-  • Complexity analysis for optimal decision tree problems: Most studies on ODT algorithms either do not analyze worst-case complexity or analyze it using incorrect combinatorial characterizations.
+  • Complexity analysis for optimal decision tree problems: In Subsection 5.3, Most studies on ODT algorithms either do not analyze worst-case complexity or analyze it using incorrect combinatorial characterizations.
 Our analysis provides an upper bound for size-constrained decision trees, which also implies trivial lower bounds for depth-constrained axis-parallel decision tree problem.
-Moreover, this lower bound contradicts the function for counting the combinatorial complexity of the problem in Mazumder et al.
-(2022)’s source code, which may indicate a potential flaw regarding the exhaustiveness of their algorithm.
-  • Challenge to existing algorithms: Because our algorithms are derived from a formal axiomatic definition of decision trees, they challenge prior studies that applied a similar algorithmic process to ODT problems with binary feature data (Demirović et al., 2022; Hu et al., 2019; Lin et al., 2020; Zhang et al., 2023; Aglin et al., 2021, 2020).
-These studies do not satisfy the axioms we propose, which are essential for justifying the correctness of a critical component, yet this component remain unspecified or ambiguously described in their constructions.
-It is therefore unclear whether their ambiguous algorithmic processes can support the claimed optimality of these algorithms, as none formally specified their problem or proved correctness.
-Our results call into question the validity of these claims.
-Achievement Part II
-  • Crossed-hyperplane properties: We identify the crossed-hyperplane properties, which significantly reduce the size of the search space for the HODT problem.
-  • Incremental algorithm for ancestry matrix generation: Leveraging the prefix-closed property of crossed hyperplanes, we introduce an incremental algorithm for efficiently generating the ancestry relation matrix, which is crucial for solving size-constrained decision tree problems.
-This algorithm can be fully vectorized and massively parallelized.
-  • Application of size-constrained ODT algorithms to HODT problem: We apply the two rigorously proven algorithms from Part I to solve the size-constrained HODT problem, encompassing classical axis-parallel hyperplanes, general hyperplanes, and hypersurfaces.
-Both algorithms have polynomial complexity (for fix tree size and data dimension), are easily parallelizable, and amenable to speed-up techniques.
-  • Experiments on synthetic data: Using the two heuristic methods we proposed, we conduct comprehensive experiments on hyperplane decision algorithms with synthetic datasets, where the ground truth is generated by hyperplane decision trees (i.e., polygonal regions).
-The results show that hyperplane decision trees recover the ground truth more accurately than both heuristic methods and optimal axis-parallel tree algorithms under various noise conditions.
-Notably, our algorithms can even achieve solutions that are more accurate than the ground truth on noisy training data, refuting the notion that optimal algorithms necessarily overfit.
-  • Benchmarking on real-world datasets: We perform a comprehensive analysis on 30 real-world datasetss, comparing hypersurface decision tree algorithms with CART and Brița et al.
-(2025)’s AODT algorithm (ConTree algorithm).
-The results demonstrate that hyperplane decision trees achieve superior generalization accuracy compared with axis-parallel decision trees when tree size is properly controlled.
+Our lower bound analysis reveals a discrepancy with the function used to count the combinatorial complexity in the source code of Mazumder et al.
+(2022).
+This observation suggests a possible limitation in the exhaustiveness of their algorithm, which could warrant further investigation to refine the approach.
 A summary of Part I of this paper is as follows.
@@
 However, his argument was unsatisfactory to Gödel.
-As (Wadler, 2015) noted:“ Church merely presented the definition of $\lambda$-definability and baldly claimed it corresponded to effective calculability.”
+As Wadler (2015) noted:“ Church merely presented the definition of $\lambda$-definability and baldly claimed it corresponded to effective calculability.”
 In contrast, Turing conducted a detailed analysis of the capabilities of a “computer”—at the time, the term referred to a human performing a computation assisted by paper and pencil.
@@
 However, establishing the optimality of an algorithm requires more than just validating a bounding technique.
-In essence, much of this work amounts to the design of “optimal bounding techniques” rather than the development of truly “optimal algorithms” for the decision tree problem.
+In essence, much of these works amounts to the design of “optimal bounding techniques” rather than the development of truly “optimal algorithms” for the decision tree problem.
 These are fundamentally different research directions, as a bounding technique is meaningful only when it supports a provably correct algorithm in the first place.
@@
 In some cases, the space is not even ambiguously defined, as in specification (1), making it unclear which ODT problem they are actually attempting to solve.
-(Lin et al., 2020; Zhang et al., 2023; van der Linden et al., 2023; Demirović et al., 2022; Hu et al., 2019) or specify their problem ambiguously, as in (Mazumder et al., 2022; Brița et al., 2025), where the problem is expressed in a form similar to
+(Lin et al., 2020; Zhang et al., 2023; van der Linden et al., 2023; Demirovic et al., 2022; Hu et al., 2019) or specify their problem ambiguously, as in (Mazumder et al., 2022; Brita et al., 2025), where the problem is expressed in a form similar to
 $s^{*}=\text{argmin}_{s\in\mathcal{S}}\sum_{n\in\mathcal{N}}E\left(s\left(x_{n}\right)\neq t\right)$ (1)
@@
 First, the search space $\mathcal{S}$ usually denoted by a symbol and defined through text description.
-For example, Brița et al.
+For example, Brita et al.
 (2025) wrote:“ $\mathcal{T}\left(\mathcal{D},d\right)$ describe the set of all decision trees for the datasets $\mathcal{D}$ with a maximum depth of $d$ .” Adopting such a textual description for $\mathcal{T}\left(\mathcal{D},d\right)$ is inherently ambiguous and dangerous.
@@
 In contrast, MIP specifications force researchers to define precisely what they mean by a decision tree by specifying the constraints first, enabling an unambiguous definition.
-As a result, studies that adopt such informal problem definitions cannot verify the most important step in proving the optimality of an algorithm—showing that the algorithm provably searches all possible solutions in $\mathcal{S}$ ($\mathcal{T}\left(\mathcal{D},d\right)$ in Brița et al.
+As a result, studies that adopt such informal problem definitions cannot verify the most important step in proving the optimality of an algorithm—showing that the algorithm provably searches all possible solutions in $\mathcal{S}$ ($\mathcal{T}\left(\mathcal{D},d\right)$ in Brita et al.
 (2025)), because $\mathcal{S}$ is not formally defined.
@@
 For example, although Mazumder et al.
-(2022) and Brița et al.
+(2022) and Brita et al.
 (2025) claim to solve the same problem, we found that their algorithms return different trees with significantly different objective values on the same datasets—differences too large to be attributed to numerical issues.
@@
   3. 3.
-    Algorithms should be intellectually manageable: Although a common strategy in algorithm design is to first construct a program and then verify that it satisfies the required properties, this process is often cumbersome and error-prone.
+    Algorithms should be intellectually manageable: For instance, Demirovic et al.
+(2022)’s algorithm spans nearly five pages of pseudocode, making it intellectually challenging to understand, implement, and verify.
+Although a common strategy in algorithm design is to first construct a program and then verify that it satisfies the required properties, this process is often cumbersome and error-prone.
 As Dijkstra (1972) famously quote:“ Program testing can be a very effective way to show the presence of bugs, but it is hopelessly inadequate for showing their absence.
@@
 However, this approach is problematic.
-Complex algorithms typically result in complex proofs, which not only make the algorithms less accessible but also harder to verify.
-For instance, Demirović et al.
-(2022)’s algorithm spans nearly five pages of pseudocode, making it intellectually challenging to understand, implement, and verify.
-To address this, Dijkstra—and likewise, this work—advocates a constructive approach to programming: derive programs from well-defined specifications, thereby ensuring correctness by design.
+Complex algorithms typically result in complex proofs, which not only make the algorithms less accessible but also harder to verify To address this, Dijkstra—and likewise, this work—advocates a constructive approach to programming: derive programs from well-defined specifications, thereby ensuring correctness by design.
 In combinatorial optimization, we begin with an unambiguous brute-force specification of the optimization problem and systematically derive a more efficient algorithm.
@@
 The key is to balance the cost of applying an acceleration method against the potential reduction in search space.
-For example, Brița et al.
+For example, Brita et al.
 (2025) claim that caching is a major factor behind their algorithm’s efficiency.
 However, He and Little (2025b) provide a simple analysis showing that overlapping subproblems in decision tree construction are rare—particularly for the ODT problem with axis-parallel or more complex splitting rules—making caching or memoization largely ineffective.
-In our experiments, we even observed cases where Brița et al.
+In our experiments, we even observed cases where Brita et al.
 (2025)’s algorithm explored over a million decision trees without a single cache hit, suggesting that in such scenarios, disabling caching entirely may improve performance.
@@
 Similarly, we argue that designing hardware-friendly algorithms is as important as achieving theoretical efficiency.
+For example, in Part I, we show that solving the ODT problem with intractable splitting rules depends heavily on the algorithm’s ability to exploit parallelism.
+Furthermore, in Part II, we demonstrate that an algorithm with superior theoretical performance may still run slower than a less theoretically efficient one if the latter is better aligned with parallel hardware.
 Rather than merely offering criticism or vacuous suggestions, we aim to demonstrate the effectiveness of the criteria we proposed.
@@
 We have previously formalized this notion by defining a proper decision tree, associating $\mathit{DTree}\left(\mathcal{A},\mathcal{B}\right)$with four axioms introduced in He and Little (2025b), as described below.
-Definition 1.
-[…]
-We call a decision tree that satisfies the following axioms, a proper decision tree with respect to $rs$:
+Definition 1
+Axioms for proper decision trees. We call a decision tree consists of splitting rules that satisfies the following axioms, a proper decision tree:
   1. 1.
-    Each branch node is defined by a single splitting rule $r:\mathcal{R}$, and each splitting rule subdivides the ambient space into two disjoint and connected subspaces, $r^{+}$ and $r^{-}$.
-The descendant rules of $r$ can be generated only from within $r^{+}$ and $r^{-}$,
+    Structural constraint one (Ambient space partition): Each branch node is defined by a single splitting rule $r:\mathcal{R}$, and each splitting rule subdivides the ambient space into two disjoint and connected subspaces, $r^{+}$ and $r^{-}$.
   2. 2.
-    Each leaf $L$ is defined by the intersection of subspaces $\bigcap_{p\in P_{L}}r_{p}^{\pm}$ for all the splitting rules $\left\{r_{p}\mid p\in P_{L}\right\}$ in the path $P_{L}:\mathcal{P}\left(\mathcal{A}\right)$ from the root to leaf $L$.
-The connected region (subspace) defined by $\bigcap_{p\in P_{L}}r_{p}^{\pm}$ is referred to as the decision region,
+    Structural constraint two (Path intersection in leafs): Each leaf $L$ is defined by the intersection of subspaces $\bigcap_{p\in P_{L}}r_{p}^{\pm}$ for all the splitting rules $\left\{r_{p}\mid p\in P_{L}\right\}$ in the path $P_{L}$ from the root to leaf $L$.
+The connected region (subspace) defined by $\bigcap_{p\in P_{L}}r_{p}^{\pm}$ is referred to as the decision region.
   3. 3.
-    The ancestry relation between any pair of splitting rules $r_{i}\left(\swarrow\vee\searrow\right)r_{j}$ is transitive; in other words, if $r_{i}\left(\swarrow\vee\searrow\right)r_{j}$ and $r_{j}\left(\swarrow\vee\searrow\right)r_{k}$ then $r_{i}\left(\swarrow\vee\searrow\right)r_{k}$,
+    Structural constraint three (Partition transitivity): Part I: The ancestry relation between any pair of splitting rules $r_{i}\left(\swarrow\vee\searrow\right)r_{j}$ is transitive; in other words, if $r_{i}\left(\swarrow\vee\searrow\right)r_{j}$ and $r_{j}\left(\swarrow\vee\searrow\right)r_{k}$ then $r_{i}\left(\swarrow\vee\searrow\right)r_{k}$.
+Part II: Moreover, $\boldsymbol{K}_{ij}=\pm 1$ if $r_{j}$ can be generated from $r_{i}^{\pm}$.
+As a result, any new decision rule $r$ added to a leaf must be generated within its corresponding decision region.
   4. 4.
-    For any pair of splitting rules $r_{i}$ and $r_{j}$, only one of the following three cases is true: $r_{i}\swarrow r_{j}$, $r_{i}\searrow r_{j}$, and $r_{i}\overline{\left(\swarrow\vee\searrow\right)}r_{j}$; additionally, $r_{i}\overline{\left(\swarrow\vee\searrow\right)}r_{i}$ is always true;.
+    Ancestral constraint (Uniqueness of the ancestry relation): For any pair of splitting rules $r_{i}$ and $r_{j}$, only one of the following three cases is true: $r_{i}\swarrow r_{j}$, $r_{i}\searrow r_{j}$, and $r_{i}\overline{\left(\swarrow\vee\searrow\right)}r_{j}$; additionally, $r_{i}\overline{\left(\swarrow\vee\searrow\right)}r_{i}$ is always true; in other words, the possible value of $\boldsymbol{K}_{ij}\in\left\{1,0,-1\right\}$ is unique determined for all $i,j$ , and $\boldsymbol{K}_{ii}=0$ for all $i$.
+A key observation in He and Little (2025b) is that if a tree’s splitting rules satisfy the structural constraints (Axioms 1-3), they define a class of decision tree problems.
+However, the ancestral constraint (Axiom 4), which specifies pairwise relationships between splitting rules, implicitly determines the subclass of decision tree problems under consideration.
+The proper decision tree framework adopts Axiom 4 as the core ancestral constraint because it encompasses most decision tree models in machine learning, including axis-parallel hyperplane, general hyperplane, and polynomial hypersurface decision trees.
+For decision trees over binary feature data, however, Subsection 4.4 explains how this problem violates Axiom 4, necessitating a different algorithmic approach.
+Fortunately, thanks to the deep algorithmic insights provided by our framework, this adaptation is achieved with minimal changes.
 Consider a tree with splitting rules $r$ defined as hyperplanes $h$, referred as the decision hyperplanes $h_{1}$, $h_{2}$, and $h_{3}$ defined by normal vector $\boldsymbol{w}_{1}$, $\boldsymbol{w}_{2}$, and $\boldsymbol{w}_{3}$, respectively.
@@
 $h_{1}$$h_{2}$$h_{1}^{+}\cap h_{2}^{+}$$\bar{\boldsymbol{w}}_{2}^{T}\boldsymbol{x}\geq 0$$h_{1}^{+}\cap h_{2}^{-}$$\bar{\boldsymbol{w}}_{2}^{T}\boldsymbol{x}<0$$\bar{\boldsymbol{w}}_{1}^{T}\boldsymbol{x}\geq 0$$h_{3}$$h_{1}^{-}\cap h_{3}^{+}$$\bar{\boldsymbol{w}}_{3}^{T}\boldsymbol{x}\geq 0$$h_{1}^{-}\cap h_{3}^{-}$$\bar{\boldsymbol{w}}_{3}^{T}\boldsymbol{x}<0$$\bar{\boldsymbol{w}}_{1}^{T}\boldsymbol{x}<0$
-Each hyperplane $h_{i}$ partitions $\mathbb{R}^{D}$ into two regions: the positive region $h_{i}^{+}=\left\{x\mid x\in\mathbb{R}^{D},\boldsymbol{w}_{i}^{T}\bar{x}\geq 0\right\}$ and the negative region $h_{i}^{-}=\left\{x\in\mathbb{R}^{D},\boldsymbol{w}_{i}^{T}\bar{x}<0\right\}$, where $\bar{x}=\left(x,1\right)\in\mathbb{R}^{D+1}$ denotes the data points in homogeneous coordinates.
+Each hyperplane $h_{i}$ partitions $\mathbb{R}^{D}$ into two regions: the positive region
+$h_{i}^{+}=\left\{x\mid x\in\mathbb{R}^{D},\boldsymbol{w}_{i}^{T}\bar{x}\geq 0\right\}$ and the negative region $h_{i}^{-}=\left\{x\in\mathbb{R}^{D},\boldsymbol{w}_{i}^{T}\bar{x}<0\right\}$, where $\bar{x}=\left(x,1\right)\in\mathbb{R}^{D+1}$ denotes the data points in homogeneous coordinates.
 For brevity, we omit the labels on the branch arrows $\boldsymbol{w}^{T}x<0$ and $\boldsymbol{w}^{T}x\geq 0$, assuming by default that the left branch corresponds to the “less than” case and the right branch to the “greater than or equal to” case.
@@
 […]
-Definition 2.
+Definition 2
 Ancestry relation matrix.
@@
 4.2.1 Defining combinatorial optimization problem formally through brute-force enumeration
-In the theory of transformational programming (also known as constructive algorithmics) (Bird and De Moor, 1996), combinatorial optimization problems such as 1 is solved using the following, generic, generate-filter-select paradigm,
+In the theory of transformational programming (also known as constructive algorithmics) (Bird and De Moor, 1996), combinatorial optimization problems such as (1) is solved using the following, generic, generate-filter-select paradigm,
 $s^{*}=\mathit{min}_{E}\left(\mathit{filter}_{p}\left(\mathit{gen}\left(\mathit{xs}\right)\right)\right)$ (4)
@@
 Bird and De Moor (1996) formalized classical combinatorial methods, such as dynamic programming, greedy algorithms, and D&C, using a generic recursive program (hylomorphism) and derived efficient solutions as theorems.
-Building on this, He (2025) extended the framework to include BnB by incorporating search strategies into Bird’s theorems, further broadening the applicability of this approach.
+Building on this, He (2025b) extended the framework to include BnB by incorporating search strategies into Bird’s theorems, further broadening the applicability of this approach.
 In practice, one only needs to verify the conditions of these theorems, after which program fusion follows automatically.
-Furthermore, He (2025) established the following inclusion relationships among different combinatorial optimization methods:
+Furthermore, He (2025b) established the following inclusion relationships among different combinatorial optimization methods:
 $\text{SDP}\subseteq\text{Greedy algorithm}\subseteq\text{BnB}\subseteq\text{General SDP}\subseteq\text{ DP},\text{Classical D\&C}\subseteq\text{\text{General D\&C}},$ (7)
@@
 Greedy algorithms represent the simplest instance, corresponding to the fused program of an SDP generator with the selector $\mathit{min}$.
-Importantly, this inclusion is not based on intuitive classification but on a rigorous formal abstraction of each method using abstract recursion, and these inclusion relations naturally emerge from the abstract level of these recursions.
-See Section II.2.9 of He (2025) for precise definitions of these terms and explanations.
+Importantly, the inclusion (7) is not based on intuitive classification but on a rigorous formal abstraction of each method using abstract recursion, and these inclusion relations naturally emerge from the abstract level of these recursions.
+See Section II.2.9 of He (2025b) for precise definitions of these terms and explanations.
 In order to construct an efficient combinatorial optimization algorithm using constructive approach, two key aspects must be considered in the design of the generator: (i) the efficiency of the generator itself, and (ii) its feasibility for program fusion, i.e., the compatibility between the structures of $\mathit{gen}$ and the objective function $E$.
@@
 For a decision tree satisfying Axiom 1, He and Little (2025b) demonstrated the following important theorem, which reveals a deep connection between the combinatorics of decision trees and permutations.
-Theorem 1.
+Theorem 3
 A decision tree consisting of $K$ splitting rules corresponds to a unique $K$-permutation if and only if it is proper.
 In other words, there exists an injective mapping from proper decision trees to valid $K$-permutation (i.e., $K$-permutations that satisfies the proper decision tree axiom).
-Proof.
-See He and Little (2025b) for more details. ∎
+Proof See He and Little (2025b) for more details.
 As a consequence, the following lemma is immediate.
-Lemma 1.
+Lemma 4
 The search space of decision trees with $K$ splitting rule $\mathcal{S}_{\text{size}}\left(K,\mathit{rs}\right)$ with respect to a list of rules $rs$ is included in $\mathcal{S}_{\text{kperms}}\left(K,\mathit{rs}\right)$, the space of $K$-permutations with respect to $rs$.
@@
 $\mathit{odt}_{\text{rec}}\left(K\right)$ $:\left[\mathcal{R}\right]\to\mathit{DTree}\left(\mathcal{R},\mathcal{D}\right)$ (18)
-$\mathit{odt}_{\text{rec}}\left(K\right)$ $=\mathit{min}_{E_{\text{0-1}}}\circ\mathit{genDTKs}_{K}.$
+$\mathit{odt}_{\text{rec}}\left(K\right)$ $=\mathit{min}_{E_{\text{0-1}}}\circ\mathit{genDTSs}_{K}.$
 $=\mathit{min}_{E_{\text{0-1}}}\circ\mathit{concatMapL}{}_{\mathit{genDTs}_{\text{rec}}\left(\mathit{xs}\right)}\circ\mathit{kcombs}\left(K\right).$
@@
 which call $\mathit{genDTs}_{\text{vec}}$ recursively, and append a new rule $r$ to every tree $t$ generated by $\mathit{genDTs}_{\text{vec}}\left(\mathit{rs}\backslash r,xs\right)$ by using $\mathit{update}$ function
-$\mathit{update}$ $:\mathcal{R}\to\mathit{DTree}\left(\mathcal{R},\mathcal{D}\right)\to\mathit{Maybe}\left(\mathit{DTree}\left(\mathcal{R},\mathcal{D}\right)\right)$
+$\mathit{update}$ $:\mathcal{R}\to\mathit{DTree}\left(\mathcal{R},\mathcal{D}\right)\to\mathit{Maybe}\left(\mathit{DTree}\left(\mathcal{R},\mathcal{D}\right)\right)$ (21)
 $\mathit{update}$ $\left(r,\mathit{DL}\left(\mathit{xs}\right)\right)=\mathit{Just}\left(DN\left(\mathit{DL}\left(xs^{+}\right),r,\mathit{DL}\left(xs^{-}\right)\right)\right)$
@@
 Although the definition of e $\mathit{update}$ update appears complex, it is applied recursively along the path of the old tree $\mathit{DN}\left(u,s,v\right)$ to determine whether adding a new splitting rule $r$ would result in a non-proper tree.
-If the tree remains proper, rris added to the leaf; otherwise, $\mathit{Nothing}$ is returned, indicating a non-proper decision tree.
+If the tree remains proper, $r$ is added to the leaf; otherwise, $\mathit{Nothing}$ is returned, indicating a non-proper decision tree.
 Because the computation in $\mathit{genDTs}_{\text{vec}}$ is sequential, it can be fully vectorized.
@@
 We discuss this advantage in detail in Subsection 4.6.
-One can check $\mathit{genDTs}_{\text{vec}}$ will generate the same set of proper decision trees as $\mathit{genDTs}_{\text{rec}}$, but will produce different list of trees.
+One can check $\mathit{genDTs}_{\text{vec}}$ will generate the same set of proper decision trees as $\mathit{genDTs}_{\text{rec}}$, but will produce different lists of trees.
 This is because different valid $K$-permutation can correspondent to the same tree (vice is not true).
@@
 Definition based on tree datatype for depth-constrained tree
-Depth-constrained decision trees are often used in the study of ODT problems with smaller combinatorial complexity, such as axis-parallel decision trees (Mazumder et al., 2022; Brița et al., 2025) or decision trees over binary feature data Demirović et al.
+Depth-constrained decision trees are often used in the study of ODT problems with smaller combinatorial complexity, such as axis-parallel decision trees (Mazumder et al., 2022; Brita et al., 2025) or decision trees over binary feature data Demirovic et al.
 (2022); Verwer and Zhang (2019); Nijssen and Fromont (2007); Aglin et al.
@@
 $\mathit{genDTDs}_{\text{depth}}$ $\left(1,\mathit{xs}\right)=\left[DN\left(\mathit{DL}\left(\mathit{xs}^{+}\right),r,\mathit{DL}\left(\mathit{xs}^{-}\right)\right)\mid r\leftarrow\mathit{gen_{splits}}\left(\mathit{xs}\right)\right]$
-$\mathit{genDTDs}_{\text{depth}}$ $\left(d,\mathit{xs}\right)=\left[\mathit{DN}\left(u,r,v\right)\mid u\leftarrow\mathit{genDTDs}_{\text{depth}}\left(\mathit{xs}^{+},d-1\right),v\leftarrow\mathit{genDTDs}_{\text{depth}}\left(\mathit{xs}^{-},d-1\right),r\leftarrow\mathit{gen_{splits}}\left(\mathit{xs}\right)\right]$
+$\mathit{genDTDs}_{\text{depth}}$ $\left(d,\mathit{xs}\right)=\big[\mathit{DN}\left(u,r,v\right)\mid$
+$u\leftarrow$ $\mathit{genDTDs}_{\text{depth}}\left(\mathit{xs}^{+},d-1\right),v\leftarrow\mathit{genDTDs}_{\text{depth}}\left(\mathit{xs}^{-},d-1\right),r\leftarrow\mathit{gen_{splits}}\left(\mathit{xs}\right)\big]$
 This definition states that all possible decision trees of depth $d$ with respect to data list $\mathit{xs}$ can be constructed from decision trees of depth d1d-1 on the subsets $\mathit{xs}^{+}$ and $\mathit{xs}^{-}$.
@@
 For the classical axis-parallel ODT problem, $\mathit{gen_{splits}}$ is straightforward: splitting rules can be generated from axis-parallel hyperplanes based on the data points.
-Equivalently, Brița et al.
+Equivalently, Brita et al.
 (2025); Mazumder et al.
 (2022) use the medians of adjacent pairs of data points for each dimension $D$.
+A more detailed discussion of the various definitions of $\mathit{gen_{splits}}$ for solving different ML problems is provided in Subsection 5.2.
 For a data list of size $N$, there are $N\times D$ possible splitting rules.
@@
 $\mathit{odt}_{\text{depth}}$ $\left(d\right)=\mathit{min}{}_{E}\circ\mathit{genDTDs}_{\text{depth}}\left(d\right)$
-Some readers may find the recursion 23 closely related to the algorithm given by Brița et al.
+Some readers may find the recursion 24 closely related to the algorithm given by Brita et al.
 (2025), which has similar recursive structure.
-We will see shortly that Brița et al.
-(2025)’s algorithm can indeed be derived from the specification (24), which proves why Brița et al.
+We will see shortly that Brita et al.
+(2025)’s algorithm can indeed be derived from the specification (25), which proves why Brita et al.
 (2025)’s algorithm is optimal.
-4.2.3 Remarks of potential flaws for optimal decision tree problems over binary feature data
-The recursions 23 and 18 are valid because, according to Axiom 1 of decision trees, subtrees can only be constructed from the subregions defined by $r$, which partition the data list $\mathit{xs}$ to $\mathit{xs}^{+}$ and $\mathit{xs}^{-}$.
-This allows us to fix a root $r$, and safely divide the remaining splitting rules into two disjoint lists of splitting rules, they correspond to $\mathit{gen_{splits}}\left(\mathit{xs}^{\pm}\right)$ in $\mathit{odt}_{\text{depth}}$ and $r^{\pm}$ in $\mathit{odt}_{\text{rec}}$.
-In particular, we do not need to worry about elements of $\mathit{gen_{splits}}\left(\mathit{xs}^{\pm}\right)$ or $r^{\pm}$ appearing in $\mathit{gen_{splits}}\left(\mathit{xs}^{\mp}\right)$ or $r^{\mp}$ because of Axiom 4, which ensures that only one of the three possible cases occurs.
-However, in the decision tree problem over binary feature data, a selected feature may appear in either the left or right subtree, depending on the presence or absence of the feature, regardless of the root $r$.
-Consequently, for any chosen root $r\in\mathit{gen_{splits}}\left(\mathit{xs}\right)$ (or $r\in\mathit{rs}$ for $\mathit{odt}_{\text{rec}}$) any subset of $\r\mathit{gen_{splits}}\left(\mathit{xs}\right)\backslash$ r or $\r\mathit{rs}\backslash$ r can form a feasible subtree.
-This substantially increases the complexity of splitting in binary-feature decision trees.
-It is therefore unclear why existing studies on ODT over binary feature data—such as Zhang et al.
-(2023); Demirović et al.
-(2022); Aglin et al.
-(2021); Lin et al.
-(2020); Hu et al.
-(2019); Verwer and Zhang (2019); Nijssen and Fromont (2010, 2007)—define splitting rules in ways that are either ambiguous or entirely unspecified.
-For instance, Demirović et al.
-(2022) provide a relatively clear definition using a recursive formula (Equation (1) in their paper).
-Their method proceeds by sequentially scanning the features (denoted as $i\in\left[0,n-1\right]$ in Equation (1), where $n$ is the number of features) and partitioning the remaining elements by selecting a data point with index $i$.
-Here is our hypothesis: for each fix index $i$, Demirović et al.
-(2022) partition the index list $\left[0,n-1\right]$ into subsets of indices smaller than $i$ (size $i$) and greater than $i$ (size ni1n-i-1), reminiscent of Catalan-style splitting.
-However, the trees constructed over binary feature data do not satisfy the standard decision tree axioms: a descendant feature $j$ can simultaneously appear as both the left and right child of its ancestor $i$, and fixing a root $i$ does not uniquely determine how the remaining n1n-1 features are split.
-The only plausible explanation for their algorithmic process is that they implicitly fix an ordering of features by arranging them as a list $\left[0,n-1\right]$, which explains their use of list notation rather than set notation.
-In this sense, a fixed-order labeled set (i.e., a list) is equivalent to an unlabeled set, which justifies our description of their recursion as “Catalan-style splitting,” originally designed for unlabeled trees.
-Moreover, line 9 of Algorithm 2 in their paper explicitly uses a for-loop to scan all features, which further supports our interpretation.
-Therefore, unless Demirović et al.
-(2022) can demonstrate that such splitting rules exhaustively enumerate all possible labeled decision trees over binary feature data, their claim of optimality remains, at best, premature.
-Also, Demirović et al.
-(2022) state that their algorithms can be considered starting points for their work, which may suggest that the same splitting technique was employed.
-Verwer and Zhang (2019) also adopt a splitting principle similar to that of Demirović et al.
-(2022).
-Hu et al.
-(2019) define leaves as either unsplittable ($d_{\text{un}}$) or splittable ($d_{\text{split}}$), but the criteria for assigning leaves to $d_{\text{un}}$ remain unclear.
-Their recursive process first scans all possible combinations of $d_{\text{split}}$ and then scans subsets of these combinations.
-It is unclear why this two-step approach is necessary, as considering all subsets directly would be equivalent.
-Moreover, Hu et al.
-(2019) present a highly complex formula (Equation (1) in their paper) for counting the combinatorics of the decision tree problem over binary feature data, yet they provide neither explanation nor justification.
-Consequently, it remains unclear why the recursive process in their algorithm can exhaustively enumerate the recursion it describes.
-In their subsequent study, Lin et al.
-(2020) modify the splitting function to resemble the approach of Aglin et al.
-(2021); Nijssen and Fromont (2007, 2010) scanning all possible splitting rules sequentially and partitioning the remaining data using an undefined $\mathit{split}$ function (line 13-14 in Algorithm 1 GOSDT).
-Without formal problem definitions and rigorous proofs of algorithmic optimality, these studies on ODT over binary feature data may be incorrect.
-Even if their final algorithms or implementations produce correct results, ambiguities in definitions can mislead newcomers and impede progress in the field.
-Therefore, it is highly desirable to provide formal proofs demonstrating the correctness—or incorrectness—of these algorithms.
-Unfortunately, this is currently impossible due to ambiguous problem definitions in these works
 4.3 Four solutions for optimal decision tree problems
@@
 This reduction not only makes the resulting algorithm significantly more efficient but also provides greater flexibility for designing parallel algorithms.
-Theorem 2.
+Theorem 5
 Simplified decision tree problem. Given a list of rules $\mathit{rs}:\left[\mathcal{R}\right]$ and a size constraint $K:\mathbb{N}$, we have following equivalence
@@
 where $\mathit{sodt}\left(\mathit{xs}\right)=\mathit{min}_{E_{\text{0-1}}}\circ\mathit{genDTs}\left(\mathit{xs}\right)$.
-Proof.
-The result is a simple consequence of distributivity
+Proof The result is a simple consequence of distributivity
 $\mathit{min}{}_{E}\circ\mathit{mapL}{}_{genDTs\left(\mathit{xs}\right)}\circ\mathit{kcombs}{}_{K}$
@@
 $\mathit{perms}$ $\left(r:\mathit{rs}\right)=\mathit{ins}\left(r,\mathit{perms}\left(\mathit{rs}\right)\right)$
-where function $\mathit{ins}$ insert $r$ into all possible positions for each partial permutations in $\mathit{perms}\left(\mathit{rs}\right)$ ( see subsection II.1.2.4 in He (2025) for details).
+where function $\mathit{ins}$ insert $r$ into all possible positions for each partial permutations in $\mathit{perms}\left(\mathit{rs}\right)$ ( see subsection II.1.2.4 in He (2025b) for details).
 Comparing (27) with (26), we can define $\mathit{sodt}$ using $\mathit{perms}$ through the following equivalence
@@
 The following lemma trivially holds.
-Lemma 2.
+Lemma 6
 Monotonicity in the decision tree problem. Given left subtrees $u$ and $u^{\prime}$ and right subtrees $v$ and $v^{\prime}$ rooted at $r$, the implication
-$E\left(\mathit{mapD}{}_{\cap_{r^{+}}}\left(u\right)\right)\leq E\left(\mathit{mapD}{}_{\cap_{r^{\prime+}}}\left(u^{\prime}\right)\right)\wedge E\left(\mathit{mapD}{}_{\cap_{r^{-}}}\left(v\right)\right)\leq E\left(\mathit{mapD}{}_{\cap_{r^{\prime-}}}\left(v^{\prime}\right)\right)\Longrightarrow E\left(DN\left(u,r,v\right)\right)\leq E\left(DN\left(u^{\prime},r^{\prime},v^{\prime}\right)\right)$ (32)
+$E\left(\mathit{mapD}_{\cap_{r^{+}}}(u)\right)\leq$ $E\left(\mathit{mapD}_{\cap_{r^{\prime+}}}(u^{\prime})\right)\wedge E\left(\mathit{mapD}_{\cap_{r^{-}}}(v)\right)\leq E\left(\mathit{mapD}_{\cap_{r^{\prime-}}}(v^{\prime})\right)$ (34)
+$\quad\Longrightarrow E\left(DN(u,r,v)\right)\leq E\left(DN(u^{\prime},r^{\prime},v^{\prime})\right)$
 only holds, in general, if $r=r^{\prime}$.
@@
 We now show that this alignment implies a dynamic programming (DP) solution.
-By leveraging the monotonicity property 2, we can fuse the minimizer $\mathit{min}_{E}$ with $\mathit{genDTs}_{\text{rec}}$ in the simplified decision tree problem defined in 28.
-This yields a DP solution $\mathit{sodt}_{\text{rec}}$ for the problem $\mathit{min}_{E}\circ\mathit{genDTs}_{\text{rec}}$ 28.
+By leveraging the monotonicity property 6, we can fuse the minimizer $\mathit{min}_{E}$ with $\mathit{genDTs}_{\text{rec}}$ in the simplified decision tree problem defined in 30.
+This yields a DP solution $\mathit{sodt}_{\text{rec}}$ for the problem $\mathit{min}_{E}\circ\mathit{genDTs}_{\text{rec}}$ 30.
 Thus providing an efficient algorithm for solving the optimal decision tree problem by substituting the DP solution $\mathit{sodt}_{\text{rec}}$ into $\mathit{odt}_{\text{rec}}$.
-Theorem 3.
+Theorem 7
 Dynamic programming algorithm for size-constrained decision tree.
@@
 provides a solution to the brute-force specification (30). In other words, we have following relation
-$\mathit{sodt}_{\text{rec}}\left(\mathit{xs}\right)\subseteq\mathit{min}_{E}\circ\mathit{genDTs}_{\text{rec}}\left(\mathit{xs}\right),$
+$\mathit{sodt}_{\text{rec}}\left(\mathit{xs}\right)\subseteq\mathit{min}_{E}\circ\mathit{genDTs}_{\text{rec}}\left(\mathit{xs}\right),$ (36)
 where the symbol $\subseteq$” means the solution on the left is one of the solution on the right.
-Proof.
-Due to the monotonicity of the problem, we can now derive the program by following equational reasoning
+Proof Due to the monotonicity of the problem, we can now derive the program by following equational reasoning
 $\mathit{min}_{E}\circ genDTs_{\text{rec}}\left(\mathit{xs}\right)$
@@
 A similar reasoning applies to the ODT problem with depth constraints, commonly used in axis-parallel and binary-feature decision trees.
-Brița et al.
+Brita et al.
 (2025) present an algorithm for this problem; however, instead of formally deriving it, they present it directly.
 While their approach can be seen as a special case of the algorithm derived here, a formal derivation is important for understanding the underlying principles and for establishing a rigorous foundation.
-The following theorem explains why the algorithm of Brița et al. (2025) is correct and how it can be systematically derived from (24).
-Theorem 4.
+The following theorem explains why the algorithm of Brita et al. (2025) is correct and how it can be systematically derived from (25).
+Theorem 8
 Dynamic programming algorithm for depth-constrained decision tree.
@@
 $\mathit{odt}_{\text{depth}}\subseteq\mathit{min}_{E}\circ\mathit{genDTDs}_{\text{depth}}\left(d\right).$
-Proof.
-The base cases ($d=1$ and $d=0$) follow directly from the definition of an optimal decision tree:
+Proof The base cases ($d=1$ and $d=0$) follow directly from the definition of an optimal decision tree:
 For $d=0$, no classification is performed; all data points are assigned to a single leaf.
@@
 $=$ $\{\text{definition of $\mathit{genDTDs}_{\text{depth}}$}\}$
-$\mathit{min}_{E}\left[\mathit{DN}\left(u,r,v\right)\mid u\leftarrow\mathit{genDTDs}_{\text{depth}}\left(d-1,\mathit{xs}^{+}\right),v\leftarrow\mathit{genDTDs}_{\text{depth}}\left(d-1,\mathit{xs}^{-}\right),r\leftarrow\mathit{gen_{splits}}\left(\mathit{xs}\right)\right]$
+$\mathit{min}_{E}\big[\mathit{DN}\left(u,r,v\right)\mid u\leftarrow\mathit{genDTDs}_{\text{depth}}\left(d-1,\mathit{xs}^{+}\right),$
+$\quad\quad v\leftarrow\mathit{genDTDs}_{\text{depth}}\left(d-1,\mathit{xs}^{-}\right),r\leftarrow\mathit{gen_{splits}}\left(\mathit{xs}\right)\big]$
 $\subseteq$ { monotonicity (6) }
-$\left[\mathit{DN}\left(\mathit{min}_{E}\left(\mathit{genDTDs}_{\text{depth}}\left(d-1,\mathit{xs}^{+}\right)\right),r,\mathit{min}_{E}\left(\mathit{genDTDs}_{\text{depth}}\left(d-1,\mathit{xs}^{-}\right)\right)\right)\mid r\leftarrow\mathit{gen_{splits}}\left(\mathit{xs}\right)\right]$
+$\big[\mathit{DN}\big(\mathit{min}_{E}\left(\mathit{genDTDs}_{\text{depth}}\left(d-1,\mathit{xs}^{+}\right)\right),$
+$\quad\quad r,\mathit{min}_{E}\left(\mathit{genDTDs}_{\text{depth}}\left(d-1,\mathit{xs}^{-}\right)\right)\big)\mid r\leftarrow\mathit{gen_{splits}}\left(\mathit{xs}\right)\big]$
 $=$ { definition of $\mathit{odt}_{\text{depth}}$ }
 $\left[\mathit{DN}\left(\mathit{odt}_{\text{depth}}\left(d-1,\mathit{xs}^{+}\right),r,\mathit{odt}_{\text{depth}}\left(d-1,\mathit{xs}^{-}\right)\right)\mid r\leftarrow\mathit{gen_{splits}}\left(\mathit{xs}\right)\right]$
-This proved that $\mathit{odt}_{\text{depth}}\left(d,\mathit{xs}\right)$ can be computed recursively using optimal subtrees for $\mathit{xs}^{+}$ and $\mathit{xs}^{-}$ of depth d1d-1. ∎
-Interestingly, for AODT and ODT over binary-feature data, Brița et al.
-(2025) and Demirović et al.
+This proved that $\mathit{odt}_{\text{depth}}\left(d,\mathit{xs}\right)$ can be computed recursively using optimal subtrees for $\mathit{xs}^{+}$ and $\mathit{xs}^{-}$ of depth d1d-1.
+Interestingly, for AODT and ODT over binary-feature data, Brita et al.
+(2025) and Demirovic et al.
 (2022) observed that $\mathit{odt}_{\text{depth}}\left(2,\mathit{xs}\right)$ can be solved more efficiently than the classical recursive method.
-Expressing $\mathit{odt}_{\text{depth}}\left(2,\mathit{xs}\right)$ in a more compact form and substituting it into the recursion 34, may help convey their algorithmic insights more clearly.
+Expressing $\mathit{odt}_{\text{depth}}\left(2,\mathit{xs}\right)$ in a more compact form and substituting it into the recursion 37, may help convey their algorithmic insights more clearly.
 4.3.3 Non-existence of monotonicity
@@
 The following theorem formalizes the non-existence of monotonicity for $\mathit{sodt}_{\text{kperms}}$ and $\mathit{sodt}_{\text{rec}}$.
-Theorem 5.
+Theorem 9
 Non-existence of monotonicity.
 For objective function defined according to (32), the one-step recursive process in $\mathit{sodt}_{\text{kperms}}$ and $\mathit{sodt}_{\text{rec}}$ does not have the monotonicity.
-Proof.
-It suffices to show that the following implication does not hold:
+Proof It suffices to show that the following implication does not hold:
 $E\left(t\right)\leq E\left(t^{\prime}\right)\nRightarrow E\left(\mathit{update}\left(r,t\right)\right)\leq E\left(\mathit{update}\left(r,t^{\prime}\right)\right),$
@@
 $E\left(\mathit{rs}\right)\leq E\left(\mathit{rs}^{\prime}\right)\nRightarrow E\left(\mathit{insR}\left(r,\mathit{rs}\right)\right)\leq E\left(\mathit{insR}\left(r,\mathit{rs}^{\prime}\right)\right)$
-where $\mathit{insR}:\left[\mathcal{R}\right]\to\left[\mathcal{R}\right]$ is the relational version of the $\mathit{ins}$ operation He (2025), which inserts $r$ to a random position in a existing permutation of rules $\mathit{rs}$.
+where $\mathit{insR}:\left[\mathcal{R}\right]\to\left[\mathcal{R}\right]$ is the relational version of the $\mathit{ins}$ operation He (2025b), which inserts $r$ to a random position in a existing permutation of rules $\mathit{rs}$.
 The implication fails because the newly inserted rule $r$ can become the root of the resulting tree, which can change the error non-monotonically.
@@
 For instance, if the objective function is incrementally definable—such as the tree size—then a greedy solution exist.
-∎
 4.3.4 The use of memoization or caching
@@
 For instance, Aglin et al.
-(2021) employ caches over itemsets (collections of Boolean features), while Demirović et al.
+(2021) employ caches over itemsets (collections of Boolean features), while Demirovic et al.
 (2022) note a trade-off between branch and dataset caching, as introduced by Nijssen and Fromont (2007).
@@
 Thus, no trade-off exists between branch and dataset caching: both subtrees (depth- or size-constrained) and datasets must be stored.
-In this context, the caching strategy in Brița et al.
+In this context, the caching strategy in Brita et al.
 (2025) appears correct, as their ConTree algorithm reuses cached solutions defined both by sub-dataset and subtree depth.
@@
 Denote the set of possible splitting rules with respect to data list $\mathit{xs}$ as $\mathcal{S}_{\mathit{\text{split}}}\left(\mathinner{xs}\right)$, to enable full memoization for constructing an optimal decision tree of size $K$, one would need to store at least:
-$O\left(\left|\mathcal{S}_{\mathit{\text{split}}}\left(\mathinner{xs}\right)\right|\times\sum_{k=0}^{K-1}\left|\mathcal{S}_{\text{size}}\left(k,\mathinner{xs}\right)\right|\right)$
+$O\left(\left|\mathcal{S}_{\mathit{\text{split}}}\left(\mathinner{xs}\right)\right|\times\sum_{k=0}^{K-1}\left|\mathcal{S}_{\text{size}}\left(k,\mathinner{xs}\right)\right|\right)$ (38)
 Similarly, for depth-constrained trees, the combinatorial complexity becomes:
-$O\left(\left|\mathcal{S}_{\mathit{\text{split}}}\left(\mathinner{xs}\right)\right|\times\sum_{i=0}^{d-1}\left|\mathcal{S}_{\text{depth}}\left(i,\mathinner{xs}\right)\right|\right)$
+$O\left(\left|\mathcal{S}_{\mathit{\text{split}}}\left(\mathinner{xs}\right)\right|\times\sum_{i=0}^{d-1}\left|\mathcal{S}_{\text{depth}}\left(i,\mathinner{xs}\right)\right|\right)$ (39)
 This results in formidable combinatorial complexity for most ODT problems.
@@
 For hyperplane ODT problems, the number of possible splits grows to $\mathcal{S}_{\mathcal{H}}=O\left(N^{{}^{D}}\right)$ (see Section 5).
-Consequently, the caching approach in Brița et al.
+Consequently, the caching approach in Brita et al.
 (2025), which stores only a few thousand optimal sub-trees, has a very low probability of cache hits.
@@
 This suggests that the efficiency claimed for their ConTree algorithm due to caching is overstated; most gains likely arise from factors other than caching.
+4.4 Extension to decision tree problem over binary feature data (non-proper decision tree)
+4.4.1 Definition of the search space
+A key advantage of adapting a general axiomatic framework for defining decision trees is that it provides an unambiguous method for classifying decision tree problems.
+The optimal decision tree problem over binary feature data (ODT-BF) aims to construct a tree that encodes a mapping over binary feature datasets, i.e., $t:\left\{-1,1\right\}^{D}\to L$, where $L$ denotes the set of labels.
+The ODT-BF problem is perhaps the most extensively studied case of applying combinatorial methods Hu et al.
+(2019); Lin et al.
+(2020); Zhang et al.
+(2023); Demirovic et al.
+(2022); Aglin et al.
+(2021, 2020); Nijssen and Fromont (2007, 2010).
+The tree structure of ODT-BF is defined in the same way as proper decision trees; in other words, ODT-BF also satisfies the structural constraints of decision trees (Axioms 1–3) and is therefore a valid decision tree problem.
+However, this problem has much simpler combinatorics, and indeed He (2025b) showed that its complexity is independent of the input data size, and empirical results from Hu et al.
+(2019) further indicate that, when the feature dimension is fixed, the algorithm scales linearly with data size.
+Characterization of splitting rules
+Unlike decision tree problems over continuous datasets, where splitting rules can often be interpreted as geometric partitions—such as a hyperplane dividing the space into two regions—ODT-BF cannot naturally be viewed this way.
+Some researchers classify ODT-BF within the class of axis-parallel hyperplanes, but this characterization is misleading, as it obscures the geometric meaning of axis-parallel splits.
+Binary feature data correspond to vertices of a hypercube, and splits cannot be defined by midpoints between vertices, since the data are discrete rather than continuous.
+PA more natural interpretation is to view each splitting rule in ODT-BF as a logical question.
+Specifically, given a data point $x\in\mathbb{R}^{D}$, each splitting rule in ODT-BF problem corresponds to determining whether the $i$-th feature of a data point $x\in\mathbb{R}^{D}$ is 0 or 1—equivalently, whether $x$ contains feature $i$?
+An immediate consequence of this characterization is that the number of splitting rules is independent of the dataset size, as a hypercube in $\mathbb{R}^{D}$ has at most $2^{D}$ vertexes.
+Indeed, in the ODT-BF problem, many data points are duplicates and occupy the same vertices of the cube, and there are many bounding techniques are developed based on this observation (Lin et al., 2020; Zhang et al., 2023; Hu et al., 2019).
+Ancestral constraints and splitting function
+Having characterized the splitting rules in ODT-BF as simple logical questions, we now analyze the ancestral constraints between pairs of rules.
+Suppose a rule $r_{i}$ is an ancestor of another rule $r_{j}$.
+The question of whether $x$ contains feature $i$ is independent of whether it contains feature $j$.
+This means that fixing an ancestor rule $r_{i}$ does not influence whether $r_{j}$ should go to the left or right subtree of $r_{i}$.
+In other words, if $r_{i}$ is the ancestor of $r_{j}$, both $r_{i}\swarrow r_{j}$ and $r_{i}\searrow r_{j}$ are valid.
+Consequently, Axiom 4 of the proper decision tree no longer holds, and the ODT-BF problem cannot be solved using the $\mathit{odt}_{K}$ program.
+As a result, ODT-BF satisfies only the structural constraints of decision trees, with no ancestral constraints.
+It therefore reduces to an ordinary labeled binary tree problem with no explicit constraints, yielding a complexity of $K!\times\mathit{Catalan}\left(K!\right)$ for an input of size $K$, since there are $\mathit{Catalan}\left(K!\right)$ possible tree shapes and $K!$ possible labelings.
+Without ancestral constraints, the split function for ODT-BF is straightforward: once a root $r$ is fixed, how should the two sublists $\mathit{rs}^{+}$ and $\mathit{rs}^{-}$ for its subtrees be determined from $\mathit{rs}/r$?
+The answer is straightforward: any subset of $\mathit{rs}/r$ is valid!
+The following function produces all possible sublists $\mathit{xs}^{\prime}$ along with their complementary subsets $\mathit{xs}/\mathit{xs}^{\prime}$ as a pair.
+$\mathit{subsPair}$ $\left(\left[\>\right],\mathit{ys}\right)=\left[\left(\left[\>\right],\mathit{ys}\right)\right]$
+$\mathit{subsPair}$ $\left(x:\mathit{xs},\mathit{ys}\right)=\mathit{subsPair}\left(\mathit{xs},\mathit{ys}\right)\cup\mathit{mapL}\left(f,\mathit{subsPair}\left(\mathit{xs},\mathit{ys}\right)\right)$
+where $f\left(a,\mathit{as},\mathit{bs}\right)=\left(a:\mathit{as},\mathit{bs}/a\right)$.
+For instance, running $\mathit{subsPair}\left(\left[1,2,3\right],\left[1,2,3\right]\right)$ returns
+$\left[\left(\left[\>\right],\left[1,2,3\right]\right),\left(\left[3\right],\left[1,2\right]\right),\left(\left[2\right],\left[1,3\right]\right),\left(\left[2,3\right],\left[1\right]\right),\left(\left[1\right],\left[2,3\right]\right),\left(\left[1,3\right],\left[2\right]\right),\left(\left[1,2\right],\left[3\right]\right),\left(\left[1,2,3\right],\left[\>\right]\right)\right]$, the $2^{3}$ possible sublists of $\left[1,2,3\right]$.
+Definition 10
+The split function for the ODT-BF problem is defined as
+$\mathit{splits}_{\text{BF}}\left(\mathit{rs}\right)=\mathit{subsPair}\left(\mathit{rs},\mathit{rs}\right)$ (40)
+Definition of the search space
+With the help of splitting function, He and Little (2025b) proposed the following generator to exhaustively generate the search space for decision trees over binary feature data, as formalized below.
+Definition 11
+The generator for the search space of the ODT-BF problem over the binary tree datatype is defined as
+$\mathit{genDTBFs}$ $:\mathcal{D}\times\left[\mathcal{R}\right]\to\left[\mathit{DTree}\left(\mathcal{R},\mathcal{D}\right)\right]$ (41)
+$\mathit{genDTBFs}$ $\left(\mathit{xs},\left[\>\right]\right)=\left[\>\right]$
+$\mathit{genDTBFs}$ $\left(\mathit{xs},\left[r\right]\right)=\left[\mathit{DN}\left(\mathit{DL}\left(xs^{+}\right),r,\mathit{DL}\left(xs^{-}\right)\right)\right]$
+$\mathit{genDTBFs}$ $\left(\mathit{xs},\mathit{rs}\right)=\big[\mathit{DN}\left(\mathit{mapD}{}_{\cap_{\mathit{xs}^{+}}}\left(u\right),r_{i},\mathit{mapD}{}_{\cap_{\mathit{xs}^{-}}}\left(v\right)\right)\mid r_{i}\leftarrow\mathit{rs},$
+$\left(\mathit{rs}^{+},\mathit{rs}^{-}\right)\leftarrow\mathit{splits}_{\text{BF}}\left(\mathit{rs}/r_{i}\right),u\leftarrow\mathit{\mathit{genDTBFs}}\left(\mathit{rs}^{+}\right),v\leftarrow\mathit{\mathit{genDTBFs}}\left(\mathit{rs}^{-}\right)\big].$
+The combinatorial analysis in He and Little (2025b) showed that $\mathit{genDTBFs}$ will generate $f_{\text{BF}}\left(K\right)=K!\times\mathit{Catalan}\left(K\right)$, align with our interpretation of labeled binary tree’s complexity.
+Consequently the ODT-BF problem can be defined by following brute-force program
+$\mathit{odt}_{\text{BF}}\left(\mathit{xs}\right)=\mathit{min}_{E}\circ\mathit{genDTBFs}\left(\mathit{xs}\right)$ (42)
+4.4.2 optimal decision tree algorithm over binary feature data
+As a result, following an exactly same derivation process as the Theorem 24 by changing $\mathit{splits}$ as $\mathit{splits}_{\text{BF}}$, we can derive the following dynamic programming algorithm for solving the 42 problem.
+Theorem 12
+Dynamic programming algorithm for size-constrained decision tree.
+Given a list of $K$ rules $\mathit{rs}_{K}:\left[\mathcal{R}\right]$.
+If the objective function $E$ can be defined according to (32)—that is, if Lemma Lemma 6 holds—then the solution obtained by
+$\mathit{odt}_{\text{BF}}$ $:\mathcal{D}\times\left[\mathcal{R}\right]\to\mathit{DTree}\left(\mathcal{R},\mathcal{D}\right)$ (43)
+$\mathit{odt}_{\text{BF}}$ $\left(\mathit{xs},\left[\;\right]\right)=\left[\mathit{DL}\left(\mathit{xs}\right)\right]$
+$\mathit{odt}_{\text{BF}}$ $\left(\mathit{xs},\left[r\right]\right)=\left[\mathit{DN}\left(\mathit{DL}\left(xs^{+}\right),r,\mathit{DL}\left(xs^{-}\right)\right)\right]$
+[…]
+$r\leftarrow\mathit{rs},\left(\mathit{rs}^{+},r,\mathit{rs}^{-}\right)\leftarrow\mathit{splits}_{\text{BF}}\left(\mathit{rs}\right)\big].$
+provides a solution to the brute-force specification (42).
 4.5 Further acceleration—Thinning and filtering
@@
 Specifically, the following condition that help identify how filtering can be fused into the generation process.
-Definition 3.
+Definition 13
 Prefix-closed predicate.
@@
 In particular, the following filter fusion theorem enables the integration of depth- or size-constrained conditions into leaf- or size-constrained ODT problems.
-Theorem 6.
+Theorem 14
 Filtering fusion through prefix-predicate. If a predicate $p$ is prefix-closed, Then, the solution obtained using $\mathit{sodt}_{\text{filtRec}}$, defined as
@@
 […]
+$\quad\quad\left(\mathit{rs}^{+},r,\mathit{rs}^{-}\right)\leftarrow\mathit{splits}\left(\mathit{rs}\right)\big]\big).$
 provides a solution to $\mathit{min}_{E}\circ\mathit{filter}_{p}\circ\mathit{genDTs}_{\text{rec}}$. Similarly, the solution obtained using $\mathit{odt}_{\text{filtDep}}$, defined as
@@
 $\mathit{odt}_{\text{filtDep}}$ $\left(1,\mathit{xs}\right)=\mathit{min}_{E}\left[DN\left(\mathit{DL}\left(\mathit{xs}^{+}\right),r,\mathit{DL}\left(\mathit{xs}^{-}\right)\right)\mid r\leftarrow\mathit{gen_{splits}}\left(\mathit{xs}\right)\right]$
-$\mathit{odt}_{\text{filtDep}}$ $\left(d,\mathit{xs}\right)=\mathit{min}_{E}\left(\mathit{filter}_{p}\left[\mathit{DN}\left(\mathit{odt}_{\text{filtDep}}\left(d-1,\mathit{xs}^{+}\right),r,\mathit{odt}_{\text{filtDep}}\left(d-1,\mathit{xs}^{-}\right)\right)\mid r\leftarrow\mathit{gen_{splits}}\left(\mathit{xs}\right)\right]\right)$
+$\mathit{odt}_{\text{filtDep}}$ $\left(d,\mathit{xs}\right)=\mathit{min}_{E}\big(\mathit{filter}_{p}\big[\mathit{DN}\left(\mathit{odt}_{\text{filtDep}}\left(d-1,\mathit{xs}^{+}\right),r,\mathit{odt}_{\text{filtDep}}\left(d-1,\mathit{xs}^{-}\right)\right)\mid$
+$\quad\quad r\leftarrow\mathit{gen_{splits}}\left(\mathit{xs}\right)\big]\big)$
 provides a solution to (45).
-Proof.
-The proof follows by substituting the fused generator and following the procedure of Theorems (3) and (4). For more general results and formal proofs, see He (2025). ∎
+Proof The proof follows by substituting the fused generator and following the procedure of Theorems (7) and (8). For more general results and formal proofs, see He (2025b).
 Theorem 14 provides an efficient and convenient method to incorporate any prefix-closed constraint into the ODT algorithm.
@@
 The derivation for $\mathit{sodt}_{\text{vec}}$ and $\mathit{sodt}_{\text{kperms}}$ similar to Theorem 14, but requires a different prefix-closed predicate definition, as the recursive structure of this two algorithms are sequential.
-S $e$ He (2025) for more general results.
+See He (2025b) for more general results.
 4.5.2 Thinning algorithm (bounding technique)
@@
 Indeed, the $\mathit{min}_{E}$ function can be understood as a special case of thinning with respect to a total order defined by the objective function $E$.
-In contrast, thinning is based on a preorder $R$, because transitivity is required, and total orders are too restrictive for the purposes of thinning (He, 2025; Bird and Gibbons, 2020).
+In contrast, thinning is based on a preorder $R$, because transitivity is required, and total orders are too restrictive for the purposes of thinning (He, 2025b; Bird and Gibbons, 2020).
 In a preorder relation, some configurations may be incomparable.
@@
 Without this property, thinning could discard potentially optimal solutions.
-Definition 4.
+Definition 15
 Monotonicity in thinning.
@@
 Similarly, in $\mathit{sodt}_{\text{kperms}}$, we have: $R\left(\mathit{rs},\mathit{rs}^{\prime}\right)\Longrightarrow R\left(\mathit{insR}\left(r,\mathit{rs}\right),\mathit{insR}\left(r,\mathit{rs}^{\prime}\right)\right)$.
-Likewise, in $\mathit{sodt}_{\text{rec}}$, the semantic interpretation of monotonicity in thinning can be expressed as $R\left(u,u^{\prime}\right)\wedge R\left(v,v^{\prime}\right)\subseteq R\left(\mathit{DN}\left(u,r,v\right),\mathit{DN}\left(u^{\prime},r,v^{\prime}\right)\right)$.
+Likewise, in $\mathit{sodt}_{\text{rec}}$, the semantic interpretation of monotonicity in thinning can be expressed as
+$R\left(u,u^{\prime}\right)\wedge R\left(v,v^{\prime}\right)\subseteq R\left(\mathit{DN}\left(u,r,v\right),\mathit{DN}\left(u^{\prime},r,v^{\prime}\right)\right)$.
 Finite dominance relation and global upper bound
-There are many ways to define a thinning algorithm He (2025), and its effectiveness depends on the context.
+There are many ways to define a thinning algorithm He (2025b), and its effectiveness depends on the context.
 For our purposes—the study of optimal decision tree problems—the dominance relation allows us to implement $\mathit{thin}$ simply using the $\mathit{filter}$ function.
 Therefore, we will not discuss the general construction of thinning, but instead focus on formalizing the specific dominance relations studied in ODT algorithms for machine learning research.
-He (2025) showed that dominance relations commonly used in ML can be abstracted into two types: finite dominance relation (FDR) and global upper bound (GUB).
+He (2025b) showed that dominance relations commonly used in ML can be abstracted into two types: finite dominance relation (FDR) and global upper bound (GUB).
 Both relations rely on two auxiliary functions defined as follows.
-Definition 5.
+Definition 16
 Pessimistic upper bound and optimistic lower bound.
@@
 Many existing studies on ODT problems focus on developing effective optimistic lower bounds or pessimistic upper bounds.
-He (2025) demonstrated that the following two generic dominance relations, derived using $\mathit{ub}$ and $\mathit{lb}$, satisfy the thinning laws.
+He (2025b) demonstrated that the following two generic dominance relations, derived using $\mathit{ub}$ and $\mathit{lb}$, satisfy the thinning laws.
 These abstract the dominance relations used in most optimal decision tree algorithms for machine learning.
-Theorem 7.
+Theorem 17
 Finite dominance relations and global upper bound. Given two partial configuration $a,b:\mathcal{A}$, and a global upper-bound constant $\mathit{ub}^{*}:\mathcal{A}$, defines
@@
 Then both $\mathit{fdom}$ and $\mathit{gubdom}$ satisfies the monotonicity.
-Proof.
-To prove the dominance valid, we nee to verify the thinning algorithm defined using $\mathit{fdom}$ satisfies the universal property, which is equivalent to verifying the following monotonicity (Bird and De Moor, 1996):
+Proof To prove the dominance valid, we nee to verify the thinning algorithm defined using $\mathit{fdom}$ satisfies the universal property, which is equivalent to verifying the following monotonicity (Bird and De Moor, 1996):
 $\mathit{ub}\left(a\right)\leq\mathit{lb}\left(b\right)\Longrightarrow\mathit{ub}\left(f^{n}\left(a\right)\right)\leq\mathit{lb}\left(f^{n}\left(b\right)\right)$
@@
 $\mathit{ub}^{*}\leq\mathit{lb}\left(a\right)\Longrightarrow\mathit{ub}^{*}\leq\mathit{lb}\left(f^{n}\left(a\right)\right)$
-which holds because $\mathit{ub}^{*}\leq\mathit{lb}\left(a\right)\leq\mathit{lb}\left(f^{n}\left(a\right)\right)$. ∎
-For the global upper bound dominance relation, the thinning function $\mathit{thin}:\left[\mathcal{A}\right]\to\left[\mathcal{A}\right]$ can be efficiently implemented using a simple $\mathit{filter}$ He (2025).
+which holds because $\mathit{ub}^{*}\leq\mathit{lb}\left(a\right)\leq\mathit{lb}\left(f^{n}\left(a\right)\right)$.
+For the global upper bound dominance relation, the thinning function $\mathit{thin}:\left[\mathcal{A}\right]\to\left[\mathcal{A}\right]$ can be efficiently implemented using a simple $\mathit{filter}$ He (2025b).
 Examples of lower bound in literature
-We now introduce two useful and generic lower bounds proposed by Demirović et al. (2022) and Zhang et al. (2023) for classification and regression problems, respectively.
-One particularly interesting and widely applicable bounding technique was introduced by Demirović et al.
+We now introduce two useful and generic lower bounds proposed by Demirovic et al. (2022) and Zhang et al. (2023) for classification and regression problems, respectively.
+One particularly interesting and widely applicable bounding technique was introduced by Demirovic et al.
 (2022); Lin et al.
-(2020); Brița et al.
+(2020); Brita et al.
 (2025) for classification problems.
 This technique is based on the following observation, although it relies on an assumption that has not yet been formally proven.
-Definition 6.
+Definition 18
 Similarity lower bound.
@@
 Then:
-$E_{\text{0-1}}\left(\mathit{odt}_{\text{depth}}\left(\mathit{xs}_{1}\cap\mathit{xs}_{\text{2}}\right),\mathit{xs}_{1}\cap\mathit{xs}_{\text{2}}\right)=E_{\text{0-1}}\left(\mathit{odt}_{\text{depth}}\left(\mathit{xs}_{\text{1}}\right),\mathit{xs}_{\text{1}}\right)-\left|\mathit{xs}_{1}-\mathit{xs}_{2}\right|$
+$E_{\text{0-1}}\left(\mathit{odt}_{\text{depth}}\left(\mathit{xs}_{1}\cap\mathit{xs}_{\text{2}}\right),\mathit{xs}_{1}\cap\mathit{xs}_{\text{2}}\right)=E_{\text{0-1}}\left(\mathit{odt}_{\text{depth}}\left(\mathit{xs}_{\text{1}}\right),\mathit{xs}_{\text{1}}\right)-\left|\mathit{xs}_{1}-\mathit{xs}_{2}\right|$ (54)
 This equality implicitly suggests that the optimal decision tree trained on the larger data list $\mathit{xs}_{\text{1}}$ is also optimal on the smaller data list $\mathit{xs}_{1}\cap\mathit{xs}_{\text{2}}$.
 If this were not true, then there would exist another tree $t_{\text{opt}}^{\prime}$ such that
-$E_{\text{0-1}}\left(t_{\text{opt}}^{\prime},\mathit{xs}_{1}\cap\mathit{xs}_{\text{2}}\right)<E_{\text{0-1}}\left(\mathit{odt}_{\text{depth}}\left(\mathit{xs}_{\text{1}}\right),\mathit{xs}_{\text{1}}\right)-\left|\mathit{xs}_{1}-\mathit{xs}_{2}\right|,$
+$E_{\text{0-1}}\left(t_{\text{opt}}^{\prime},\mathit{xs}_{1}\cap\mathit{xs}_{\text{2}}\right)<E_{\text{0-1}}\left(\mathit{odt}_{\text{depth}}\left(\mathit{xs}_{\text{1}}\right),\mathit{xs}_{\text{1}}\right)-\left|\mathit{xs}_{1}-\mathit{xs}_{2}\right|,$ (55)
 contradicting the bound.
@@
 It captures the key insight that the regression error within a leaf cannot be lower than the optimal solution to the one-dimensional $K$-means clustering problem over the corresponding labels.
-Theorem 8.
+Theorem 19
 $K$-means lower bound.
@@
 Then, we have the following lower bound:
-$E_{2}\left(\mathit{sodt}\left(\mathit{xs},\mathit{rs}_{K}\right),\mathit{ys}\right)\geq\text{kmeans-1D}\left(K+1,\mathit{ys}\right)$
-Proof.
-The result follows directly, as a decision tree can be viewed as a method of clustering a one-dimensional data list—a label vector.
+$E_{2}\left(\mathit{sodt}\left(\mathit{xs},\mathit{rs}_{K}\right),\mathit{ys}\right)\geq\text{kmeans-1D}\left(K+1,\mathit{ys}\right)$ (56)
+Proof The result follows directly, as a decision tree can be viewed as a method of clustering a one-dimensional data list—a label vector.
 Its clustering performance cannot surpass that of exact 1D $K$-means clustering, which is optimal with respect to squared loss.
 Consequently, any other clustering over $\mathit{ys}$, including that induced by the decision tree, will exhibit a greater squared loss.
-∎
 Fusion for thinning algorithm
@@
 The thinning algorithm for $\mathit{sodt}_{\text{rec}}$ can be derived in a similar manner.
-Lemma 3.
+Lemma 20
 Thinning fusion by finite dominance relation.
@@
 $\mathit{odt}_{\text{thindep}}$ $\left(0,\mathit{xs}\right)=\mathit{DL}\left(\mathit{xs}\right)$
-$\mathit{odt}_{\text{thindep}}$ $\left(1,\mathit{xs}\right)=\mathit{min}_{E}\left[DN\left(\mathit{DL}\left(\mathit{xs}^{+}\right),r,\mathit{DL}\left(\mathit{xs}^{-}\right)\right)\mid r\leftarrow\mathit{gen}_{\mathit{splits}}\left(\mathit{xs}\right)\right]$
+$\mathit{odt}_{\text{thindep}}$ $\left(1,\mathit{xs}\right)=\mathit{min}_{E}\left[\mathit{DN}\left(\mathit{DL}\left(\mathit{xs}^{+}\right),r,\mathit{DL}\left(\mathit{xs}^{-}\right)\right)\mid r\leftarrow\mathit{gen}_{\mathit{splits}}\left(\mathit{xs}\right)\right]$
 […]
 where $E\left(t^{*}\right)=\inf$.
-Proof.
-We prove 45 by following equational reasoning
+Proof We prove 57 by following equational reasoning
 $\mathit{min}_{E}\left(\left[\mathit{DN}\left(\mathit{odt}_{\text{depth}}\left(d-1,\mathit{xs}^{+}\right),r,\mathit{odt}_{\text{depth}}\left(d-1,\mathit{xs}^{-}\right)\right)\mid r\leftarrow\mathit{gen_{splits}}\left(\mathit{xs}\right)\right]\right)$
@@
 Its performance depends on both the implementation of $\mathit{thin}$ and the effectiveness of the relation $R$.
-He (2025) provides various recursive definitions of $\mathit{thin}$ and the effectiveness of each definition is problem-dependent.
+He (2025b) provides various recursive definitions of $\mathit{thin}$ and the effectiveness of each definition is problem-dependent.
 4.6 Comparison of four methods
 4.6.1 Comparison between size- and depth-constrained trees algorithms
-As noted, depth-constrained decision trees are commonly used in ODT studies with relatively small combinatorics, such as axis-parallel trees (Mazumder et al., 2022; Brița et al., 2025) trees over binary-feature data (Demirović et al., 2022; Verwer and Zhang, 2019; Nijssen and Fromont, 2007; Aglin et al., 2020; Nijssen and Fromont, 2010).
-These existing methods typically process recursion in a depth-first manner, unlike the breadth-first formulation we defined in 34.
+As noted, depth-constrained decision trees are commonly used in ODT studies with relatively small combinatorics, such as axis-parallel trees (Mazumder et al., 2022; Brita et al., 2025) trees over binary-feature data (Demirovic et al., 2022; Verwer and Zhang, 2019; Nijssen and Fromont, 2007; Aglin et al., 2020; Nijssen and Fromont, 2010).
+These existing methods typically process recursion in a depth-first manner, unlike the breadth-first formulation we defined in 37.
 In contrast, the size-constrained ODT algorithm has been far less studied; Hu et al.
@@
 The primary reason is that such problems require parallelization to achieve reasonable solution times, yet depth-constrained trees are difficult to implement in an embarrassingly parallel manner—parallelization requires minimal or no communication between processors.
-We first analyze the size-constrained decision tree and explain why it is suitable for parallel implementation. In 25, the original size-constrained ODT problem can be factorized as:
-$\mathit{min}_{E}\circ\mathit{concatMapL}{}_{\mathit{sodt}\left(\mathit{xs}\right)}\circ\mathit{kcombs}_{K}=\mathit{odt}_{K}$
+We first analyze the size-constrained decision tree and explain why it is suitable for parallel implementation. In 26, the original size-constrained ODT problem can be factorized as:
+$\mathit{min}_{E}\circ\mathit{concatMapL}{}_{\mathit{sodt}\left(\mathit{xs}\right)}\circ\mathit{kcombs}_{K}=\mathit{odt}_{K}$ (58)
 This factorization indicates that solving the challenging $\mathit{odt}$ problem can be achieved by solving many smaller instances of $\mathit{sodt}$.
@@
 Unfortunately, no similar factorization has yet been found for the depth-constrained ODT problem that would enable embarrassingly parallel execution.
-Consequently, the optimal decision tree algorithm derived from 23 can only be executed sequentially or parallelized post hoc, making it difficult to achieve the same level of speed-up as embarrassingly parallel algorithms.
+Consequently, the optimal decision tree algorithm derived from 24 can only be executed sequentially or parallelized post hoc, making it difficult to achieve the same level of speed-up as embarrassingly parallel algorithms.
 For hyperplane decision trees, each recursive step may involve up to $O\left(N^{D}\right)$ possible rules.
@@
 Previous research on ODT problems with smaller combinatorics remains tractable even without parallelization because the number of possible splits at each recursive step is limited to $O\left(N\times D\right)$ for axis-parallel trees.
-However, this approach becomes ineffective when applied to the larger combinatorics of 23 where splitting rules are defined as hyperplanes.
+However, this approach becomes ineffective when applied to the larger combinatorics of 24 where splitting rules are defined as hyperplanes.
 Hence, parallelization is likely the only viable solution for problems with such large combinatorial complexity.
@@
 The total degree of this monomial is the sum $\left|\boldsymbol{\alpha}\right|=\alpha_{1}+\cdot\cdot\cdot+\alpha_{n}$.
-Definition 7.
+Definition 21
 Polynomial. A polynomial $P\left(x\right)$, or $P$ in short, is a finite linear combination (with coefficients in the field $\mathbb{R}$) of monomials in the form
-$P\left(x\right)=\sum_{i}w_{i}x^{\boldsymbol{\alpha}_{i}},w_{i}\in\mathbb{R}.$
+$P\left(x\right)=\sum_{i}w_{i}x^{\boldsymbol{\alpha}_{i}},w_{i}\in\mathbb{R}.$ (59)
 We denote the set of all polynomials with variables $x_{1},x_{2}\ldots,x_{D}$ and coefficients in $\mathbb{R}$ is denoted by $\mathbb{R}\left[x_{1},x_{2}\ldots,x_{D}\right]$ or $\mathbb{R}\left[x\right]$.
@@
 This lead to the following fact.
-Fact 1.
+Fact 1
 If polynomial $P$ in $\mathbb{R}\left[x_{1},x_{2}\ldots,x_{D}\right]$ has maximal degree $M$, then polynomial $P$ has $\left(\begin{array}[]{c}D+M\\ D\end{array}\right)$ monomial terms (including constant term) at most.
@@
 Given $N$ data points in general position, the number of linearly separable dichotomies is:
-$\mathcal{S}_{\mathrm{Cover}}\left(N,D+1\right)=2\sum_{d=0}^{D}\left(\begin{array}[]{c}N-1\\ d\end{array}\right).$
+$\mathcal{S}_{\mathrm{Cover}}\left(N,D+1\right)=2\sum_{d=0}^{D}\left(\begin{array}[]{c}N-1\\ d\end{array}\right).$ (60)
 This result implies that the optimal 0-1 loss solution obtained by exhaustively searching over Cover’s space $\mathcal{S}_{\mathrm{Cover}}$ is equivalent to the optimal solution obtained by searching the entire continuous space $\mathbb{R}^{D}$.
@@
 In particular, we have following results.
-Theorem 9.
+Theorem 22
 Under the general position assumption and when optimizing the 0-1 loss, for any partition induced by a hyperplane $h$ in $\mathbb{R}^{D}$ with respect to a dataset $\mathit{xs}$, there exists $D$ points in $\mathit{xs}$ such that the partition induced by hyperplanes passing through these $D$ points coincides with the partition induced by $h$
@@
 More formally, we have the following theorem.
-Theorem 10.
+Theorem 23
 0-1 loss linear classification theorem. Consider a data list $\mathit{xs}:\left[\mathbb{R}^{D}\right]$ of $N$ points in general position, along with their associated labels. Then
-$\mathit{minR}{}_{E_{\text{0-1}}}\left\{min_{E_{\text{0-1}}}\left(\boldsymbol{w}_{s},-\boldsymbol{w}_{s}\right)\mid s\in\mathcal{S}_{\text{kcombs}}\left(D,\mathit{xs}\right)\right\}\subseteq\mathit{minR}{}_{E_{\text{0-1}}}\left\{\boldsymbol{w}\mid\boldsymbol{w}\in\mathcal{S}_{\mathrm{Cover}}\left(N,D+1\right)\right\}\subseteq\mathit{minR}{}_{E_{\text{0-1}}}\left\{\boldsymbol{w}\mid\boldsymbol{w}\in\mathbb{R}^{D}\right\}$
+$\mathit{minR}_{E_{\text{0-1}}}\left\{\min_{E_{\text{0-1}}}\left(\boldsymbol{w}_{s},-\boldsymbol{w}_{s}\right)\mid s\in\mathcal{S}_{\text{kcombs}}\left(D,\mathit{xs}\right)\right\}\subseteq$ (62)
+$\quad\mathit{minR}_{E_{\text{0-1}}}\left\{\boldsymbol{w}\mid\boldsymbol{w}\in\mathcal{S}_{\mathrm{Cover}}\left(N,D+1\right)\right\}\subseteq\mathit{minR}_{E_{\text{0-1}}}\left\{\boldsymbol{w}\mid\boldsymbol{w}\in\mathbb{R}^{D}\right\}$
 and $\boldsymbol{w}_{s}\in\mathbb{R}^{D}$ denotes the normal vector determined uniquely by $D$ data points (in general position, a unique hyperplane passes through any such $D$ points, which can be computed via matrix inversion).
-Proof.
-See He and Little (2023). ∎
+Proof See He and Little (2023).
 5.1.3 0-1 loss polynomial hypersurface classification theorem
@@
 Axis-parallel hyperplane $\left(0,\ldots,x_{i},\ldots 0\right)$ $x_{i}=a$, $\forall i\in\left[D\right]$ 1
-Affine hyperplane $\left(x_{1},\ldots,x_{D},1\right)$ $w_{1}x_{1}+w_{2}x_{2}+\ldots+w_{D}x_{D}+1=0$, $D$
+Affine hyperplane $\left(x_{1},\ldots,x_{D},1\right)$ $w_{1}x_{1}+w_{2}x_{2}+\ldots+w_{D}x_{D}+1=0$$w\bar{\boldsymbol{x}}=0$, $D$
 Hypersphere $\left(x_{1},\ldots,x_{D},x_{1}^{2},\ldots,x_{D}^{2},1\right)$ $w_{1}x_{1}+\ldots w_{D}x_{D}\ldots+w_{2D}x_{D}^{2}+1=0$ $2D$
-Conic section $\left(x_{1},\ldots,x_{1}x_{D},\ldots,x_{D}^{2},1\right)$ $w_{1}x_{1}+\ldots w_{g}x_{1}x_{2}\ldots+w_{G}x_{D}^{2}+1=0$, $G=\left(\begin{array}[]{c}D+2\\ 2\end{array}\right)-1$
+Conic section $\left(x_{1},\ldots,x_{1}x_{D},\ldots,x_{D}^{2},1\right)$ $w_{1}x_{1}+\ldots w_{g}x_{1}x_{2}\ldots+w_{G}x_{D}^{2}+1=0$$w\bar{\boldsymbol{x}}=0$, $G=\left(\begin{array}[]{c}D+2\\ 2\end{array}\right)-1$
 Arbitrary degree-$M$ polynomials $\left(x_{1},\ldots,x_{1}x_{2}^{M-1},\ldots,x_{D}^{M},1\right)$ $w_{1}x_{1}+\ldots w_{g}x_{1}x_{2}^{M-1}\ldots+w_{G}x_{D}^{M}+1=0$ $G=\left(\begin{array}[]{c}D+M\\ D\end{array}\right)-1$
@@
 If both $M$ and $D$ are fixed constants, this again gives a polynomial-time algorithm for solving the 0-1 loss hypersurface classification problem.
-Theorem 11.
+Theorem 24
 Given variables $x_{1},\ldots x_{D}$ in space $\mathbb{R}^{D}$ (which is isomorphic to the affine space $\mathbb{R}^{D}$ when ignoring the points at infinity (Cox et al., 1997)), let $M_{0},M_{1},\ldots M_{G}$ be all monomials of degree $M$ with variables $x_{0},x_{1},\ldots x_{D}$, where $G\leq\left(\begin{array}[]{c}D+M\\ D\end{array}\right)-1$.
@@
 The hyperplane classification over the embedded data lists $\rho_{M}\left(\mathit{xs}\right)$ is isomorphic to the polynomial hypersurface classification (defined by a degree $M$ polynomial) over the original data list $\mathit{xs}$.
-Proof.
-First, we need to prove the existence of the hypersurface.
+Proof First, we need to prove the existence of the hypersurface.
 For any $G$ distinct points $\mathit{s}$ in $\mathbb{R}^{D+1}$, there exist a polynomial hypersurface $S$ defined by a polynomial $P$ with degree at most $M$ such that $\mathit{s}$ lies on $S$.
 This is because the vector space $V$ of a polynomial $P$ in $\mathbb{R}\left[x_{1},\ldots x_{D}\right]$of degree $M$ has a dimension $\left(\begin{array}[]{c}D+W\\ D\end{array}\right)$.
-For any data list $ds$ of size $G<\left(\begin{array}[]{c}D+M\\ D\end{array}\right)$, the linear map from $V$ to $\mathbb{R}^{G}$ must has non-trivial kernel (the linear system has more variables than equations), thus the polynomial hypersurface $S$ must exist.
+For any data list $xs$ of size $G<\left(\begin{array}[]{c}D+M\\ D\end{array}\right)$, the linear map from $V$ to $\mathbb{R}^{G}$ must has non-trivial kernel (the linear system has more variables than equations), thus the polynomial hypersurface $S$ must exist.
 Next, consider a hyperplane $\boldsymbol{w}^{T}\bar{y}=0$ defined over variables $\bar{y}=\left(1,y_{1},\ldots y_{G}\right)\in\mathbb{R}^{G+1}$, we have $\boldsymbol{w}^{T}\bar{y}\geq$0 or $\boldsymbol{w}^{T}\bar{y}$¡0 implies $\boldsymbol{w}^{T}\rho_{M}\left(\bar{x}\right)\geq 0$ or $\boldsymbol{w}^{T}\rho_{M}\left(\bar{x}\right)<0$, because the closure of the real numbers under polynomial operation, thus $M\left(\bar{x}\right)\in\mathbb{R}$ if $\bar{x}\in\mathbb{R}^{D+1}$, for any monomial $M$.
-∎
 Theorem 24 show that the $M$-degree polynomial hypersurface classification problem over data lists in $\mathbb{R}^{D}$ is equivalent to a linear classification problem in the embedded space $\mathbb{R}^{G}$.
 By analogy with the linear classification result in Theorem 23, we can formalize the 0-1 Loss Polynomial Hypersurface Classification Theorem as follows.
-Theorem 12.
+Theorem 25
 0-1 loss polynomial hypersurface classification theorem.
@@
 Denote $E_{\text{0-1}}\left(\boldsymbol{w}\right)$ as the 0-1 loss for evaluating normal vector $\boldsymbol{w}\in\mathbb{R}^{G}$ with respect to data lists $\rho_{M}\left(\mathit{xs}\right)$.
-Then the hyper degree $M$ polynomial
-$\mathit{minR}{}_{E_{\text{0-1}}}\left\{min_{E_{\text{0-1}}}\left(\boldsymbol{w}_{s},-\boldsymbol{w}_{s}\right)\mid s\in\mathcal{S}_{\text{kcombs}}\left(G,\rho_{M}\left(\mathit{xs}\right)\right)\right\}\subseteq\mathit{minR}{}_{E_{\text{0-1}}}\left\{\boldsymbol{w}\mid\boldsymbol{w}\in\mathbb{R}^{G}\right\}$
+Then we have
+$\mathit{minR}{}_{E_{\text{0-1}}}\left\{min_{E_{\text{0-1}}}\left(\boldsymbol{w}_{s},-\boldsymbol{w}_{s}\right)\mid s\in\mathcal{S}_{\text{kcombs}}\left(G,\rho_{M}\left(\mathit{xs}\right)\right)\right\}\subseteq\mathit{minR}{}_{E_{\text{0-1}}}\left\{\boldsymbol{w}\mid\boldsymbol{w}\in\mathbb{R}^{G}\right\}$ (63)
 where $\boldsymbol{w}_{s}\in\mathbb{R}^{G}$ denote as the normal vector determined by $G$ data points in $\mathcal{S}_{\text{kcombs}}\left(G,\mathit{xs}\right)$.
@@
 We now show that the hyperplane classification tree satisfies the axioms of a proper decision tree, as verified by the following lemma.
-Lemma 4.
+Lemma 26
 Given a set of data $\mathit{xs}$ and list of $K$ hyperplanes $\mathit{hs}_{K}=\left[h_{1},h_{2},\ldots h_{K}\right]$ defined by normal vectors $ws=\left[w_{1},w_{2},\ldots w_{K}\right]$ in $\mathbb{R}^{D+1}$, where each hyperplane passes through exactly $D$ data points.
@@
 Then, the decision tree constructed using these hyperplanes as splitting rules is a proper decision tree.
-Proof.
-Axioms 1 and 2 follow directly from the definition of $h^{\pm}$. Axioms 3 and 4 hold because, for any hyperplane $h$ defined by a $D$-tuple of data points $\mathit{xs}_{D}$ we have
+Proof Axioms 1 and 2 follow directly from the definition of $h^{\pm}$. Axioms 3 and 4 hold because, for any hyperplane $h$ defined by a $D$-tuple of data points $\mathit{xs}_{D}$ we have
 $\mathit{xs}_{D}\subseteq\bigcap_{p\in P}h_{p}^{\pm}\implies\mathit{xs}_{D}\subseteq h_{p}^{\pm},\forall p\in P$
@@
 $\mathit{xs}_{D}\subseteq h^{\pm}\implies\mathit{xs}_{D}\nsubseteq h^{\mp}.$
-Thus axiom 3, 4 holds. ∎
+Thus axiom 3, 4 holds.
 By the equivalence between hyperplane and hypersurface, as established in Theorem 24, it follows immediately that decision trees whose splitting rules are characterized by hypersurfaces passing through $G$-combinations of data points are also proper.
 We next show that the optimal decision tree constructed from these hypersurfaces, with $G$ data points lying on them, is also optimal among all decision trees constructed from general hypersurfaces in $\mathbb{R}^{D}$.
-Lemma 5.
+Lemma 27
 0-1 classification theorem for hypersurfaces decision tree model.
@@
 The optimal decision tree constructed using $M$-degree polynomial hypersurfaces, each passing through $G$ points, achieves the same minimum 0-1 loss as the optimal decision tree constructed using arbitrary hypersurfaces in $\mathbb{R}^{D}$.
-Proof.
-We prove this by induction.
+Proof We prove this by induction.
 The base case, involving a single linear hypersurface, has already been established in Theorem 25.
@@
 Therefore, the optimal decision tree constructed using hypersurfaces through $G$ points attains the same 0-1 loss as the optimal tree constructed from arbitrary hypersurfaces.
-∎
 Although we assume that the data are in general position—which excludes axis-parallel hyperplanes, as they do not satisfy the general position assumption—the results also hold for axis-parallel hyperplanes, since they can be infinitesimally shifted to become hyperplanes in general position without affecting the predictions.
@@
 5.2 Splitting rule generators
+Note that both $\mathit{odt}_{\text{size}}:\left[\mathcal{R}\right]\to\mathit{DTree}\left(\mathcal{R},\mathcal{D}\right)$ and $\mathit{odt}_{\text{depth}}:\mathcal{D}\to\mathit{DTree}\left(\mathcal{R},\mathcal{D}\right)$ cannot yet be directly applied to solve the optimal decision tree problem in machine learning.
+This is because we have not explained how to generate the input $\mathit{rs}:\left[\mathcal{R}\right]$ for $\mathit{odt}_{\text{size}}$ or the function $\mathit{gen}_{\mathit{splits}}:\mathcal{D}\to\left[\mathcal{R}\right]$ used in $\mathit{odt}_{\text{depth}}$.
+It turns out that both functions are the same function.
+In this section, we describe how to construct $\mathit{gen}_{\mathit{splits}}$ for solving various ML problems, including axis-parallel, hyperplane, hypersurface, and mixed-splitting ODT problems.
+Specifically, the general paradigm for solving the size-constrained optimal decision tree problem requires composing $\mathit{odk}_{\text{size}}\left(K\right)$ with $\mathit{gen}_{\mathit{splits}}^{M}$.
+In other words, we obtain the following specification:
+$\mathit{odt}_{\text{size}}\left(K\right)\circ\mathit{gen}_{\mathit{splits}}=\mathit{min}_{E_{\text{0-1}}}\circ\mathit{concatMap}_{\mathit{sodt}\left(\mathit{xs}\right)}\circ\mathit{kcombs}\left(K\right)\circ\mathit{gen}_{\mathit{splits}}$ (64)
+where $\mathit{gen}_{\mathit{splits}}$ is precisely the splitting rule generator used in the definition of $\mathit{odk}_{\text{depth}}$.
+In principle, as long as the rules $\mathit{rs}:\left[\mathcal{R}\right]$ generated by $\mathit{gen}_{\mathit{splits}}^{M}$ satisfy the decision tree axioms 1, $\mathit{odt}_{\text{size}}\left(K\right)$ can solve the size-constrained ODT problem exactly with respect to $\mathit{rs}$.
+This modularized design provides substantial flexibility for generalizing our algorithm to different problems.
+In practice, we need only redefine $\mathit{gen}_{\mathit{splits}}^{M}$ to adapt the program to new problems, while the main program $\mathit{odt}_{\text{size}}\left(K\right)$ remains unchanged, which demonstrates one of the primary advantage of adopting a general formalism in algorithm design.
+For example, in the axis-parallel decision tree problem, $\mathit{gen}_{\mathit{splits}}^{0}:\mathcal{D}\to\left[\mathcal{H}^{0}\right]$ can be defined as the set of all possible axis-parallel hyperplanes passing through each data point in $\mathit{xs}:\mathcal{D}$.
+Thus, there are $N\times D$ such hyperplanes in total.
+More concretely, given a dataset $\mathit{xs}:\mathcal{D}$ in $\mathbb{R}^{D}$,
+$\mathit{gen}_{\mathit{splits}}^{0}\left(\mathit{xs}\right)=\left[x_{i}\mid x_{i}\leftarrow x,x\leftarrow\mathit{xs}\right],$ (65)
+By substituting $\mathit{gen}_{\mathit{splits}}^{0}$ into 1
+$\mathit{aodt}_{K}=\mathit{odt}_{\text{size}}\left(K\right)\circ\mathit{gen}_{\mathit{splits}}^{0},$
+where $\mathit{aodt}$ is short for “axis-parallel optimal decision tree problem,” which yields an efficient solution for solving the classical optimal decision tree problem in machine learning.
+Constructing generators for more complex splitting rules is more involved, but their structured form allows the composed function $\mathit{kcombs}\left(K\right)\circ\mathit{gen}_{\mathit{splits}}$ to be fused into a more efficient algorithm.
+In this section, we detail the construction of fused programs for both r $\mathit{gen}_{\mathit{splits}}^{M}$—the generator of hypersurface splitting rules—and $\mathit{gen}_{\mathit{splits}}^{m}$ ($0\leq m\leq M$) —the generator of mixed splitting rules that combine hypersurface, hyperplane, and axis-parallel rules.
 5.2.1 Generating fixed polynomial degree hypersurface splitting rules
@@
   11. 11.
-    $C_{1}=\left(\begin{array}[]{c}n\\ G\end{array}\right)$, $C_{2}=\left(\begin{array}[]{c}n+1\\ G\end{array}\right)$
+      $C_{1}=\left(\begin{array}[]{c}n\\ G-1\end{array}\right)$, $C_{2}=\left(\begin{array}[]{c}n\\ G\end{array}\right)$
   12. 12.
@@
 Algorithm 1 $\mathit{nestedCombs}$
-To solve the original decision tree problem $\mathit{odt}_{\text{size}}$ using $\mathit{sodt}$, a separate generator is required to produce possible splitting rules for a given data list $\mathit{xs}$.
-Specifically, we need an efficient definition for the following specification:
-$\mathit{odk}_{\text{size}}\left(K\right)\circ\mathit{gen}_{\mathit{splits}}^{M}=\mathit{min}_{E_{\text{0-1}}}\circ\mathit{concatMap}_{\mathit{sodt}\left(\mathit{xs}\right)}\circ\mathit{kcombs}\left(K\right)\circ\mathit{gen}_{\mathit{splits}}^{M}$
-where $\mathit{gen}_{\mathit{splits}}^{M}:\mathcal{D}\to\left[\mathcal{R}\right]$ denotes the splitting rule generator, precisely the splitting rule generator used in the definition of $\mathit{odk}_{\text{depth}}$.
-For the axis-parallel decision tree problem, $\mathit{gen}_{\mathit{splits}}^{0}\left(\mathit{xs}\right)$ can be simply defined as the all possible axis-parallel hyperplanes that go through every data points in $\mathit{xs}$, so there are $N\times D$ such hyperplanes in total.
-For more complex splitting rules, as established in Theorems 10 and 12, both hyperplane and hypersurface decision boundaries can be exhaustively enumerated by generating $G$-combinations of data points.
+Due to the isomorphism between hyperplane classification and hypersurface classification established in Theorem 24, we can define the optimal hypersurface decision tree algorithm $\mathit{hodt}^{M}:\mathbb{N}\times\mathcal{D}\to\mathit{DTree}\left(\mathcal{H}^{M},\mathcal{D}\right)$ as the composition of $\mathit{odk}_{\text{size}}\left(K\right)$ with a hypersurface splitting rule generator $\mathit{gen}_{\mathit{splits}}^{M}$, together with an embedding function $\mathit{embed}_{M}$ (the programmatic definition of $\rho_{M}$).
+In other words, we obtain the following specification:
+$\mathit{hodt}_{K}^{M}$ $=\mathit{odk}_{\text{size}}\left(K\right)\circ\mathit{gen}_{\mathit{splits}}^{M}\circ\mathit{embed}_{M}$ (66)
+$=\mathit{min}_{E_{\text{0-1}}}\circ\mathit{concatMap}_{\mathit{sodt}\left(\mathit{xs}\right)}\circ\mathit{kcombs}\left(K\right)\circ\mathit{gen}_{\mathit{splits}}^{M}\circ\mathit{embed}_{M}$
+As established in Theorems 23 and 25, both hyperplane and hypersurface decision boundaries can be exhaustively enumerated by generating $G$-combinations of data points.
 Consequently, the search space of hypersurface decision boundaries can be represented as $\mathcal{S}_{\text{kcombs}}\left(G,\mathit{ds}\right)$, and the generator for this search space can be implemented directly using a combination generator:
-$\mathit{gen}_{\mathit{splits}}^{M}\left(\mathit{ds}\right)=!\left(G\right)\circ\mathit{kcombs}\left(G,\mathit{ds}\right),$
+$\mathit{gen}_{\mathit{splits}}^{M}\left(\mathit{ds}\right)=!\left(G\right)\circ\mathit{kcombs}\left(G,\mathit{ds}\right),$ (67)
 where $\mathit{kcombs}\left(G,\mathit{ds}\right):\mathbb{N}\times\mathcal{D}\to\left[\left[\mathcal{D}\right]\right]$ generate all combinations with size smaller than $G$, and $\left(G\right)$ extract combinations of size exactly $G$.
 Hence, the composition
-$\mathit{kcombs}\left(K\right)\circ\mathit{gen}_{\mathit{splits}}^{M}=\mathit{kcombs}\left(K\right)\circ!\left(G\right)\circ\mathit{kcombs}\left(G,\mathit{ds}\right),$
+$\mathit{kcombs}\left(K\right)\circ\mathit{gen}_{\mathit{splits}}^{M}=\mathit{kcombs}\left(K\right)\circ!\left(G\right)\circ\mathit{kcombs}\left(G,\mathit{ds}\right),$ (68)
 corresponds to a combination-of-combinations generator, i.e., nested combinations, denote as $\left(K,G\right)$-nested combinations.
@@
 where $\left\langle f,g\right\rangle\left(x\right)=\left(f\left(x\right),g\left(x\right)\right)$, and $G=1$ if $M=0$ and $G=\left(\begin{array}[]{c}M+D\\ D\end{array}\right)-1$ if $M\geq 1$.
-However, the naive implementation 47 is inefficient due to the overhead of storing intermediate results: $\mathit{kcombs}\left(G\right)$ produces $O\left(N^{G}\right)$ combinations for an input of size $N$, which is both memory-intensive and computationally costly.
+Then we can rewrite 66 by following
+$\mathit{hodt}_{K}^{M}$ $=\mathit{min}_{E_{\text{0-1}}}\circ\mathit{concatMap}_{\mathit{sodt}\left(\mathit{xs}\right)}\circ\mathit{kcombs}\left(K\right)\circ\mathit{gen}_{\mathit{splits}}^{M}\circ\mathit{embed}_{M}$
+$\subseteq\mathit{min}_{E_{\text{0-1}}}\circ\mathit{concatMap}_{\mathit{sodt}\left(\mathit{xs}\right)}\circ\mathit{nestedCombs}\left(K,G\right)\circ\mathit{embed}_{M}$
+However, the naive implementation 69 is inefficient due to the overhead of storing intermediate results: $\mathit{kcombs}\left(G\right)$ produces $O\left(N^{G}\right)$ combinations for an input of size $N$, which is both memory-intensive and computationally costly.
 Additionally, its non-recursive nature prevents the application of thinning techniques.
-Nevertheless, He (2025) presents both D&C and sequential versions of $\mathit{nestedCombs}$, while the D&C version is efficient, it complicates the ordering of elements.
-Element ordering is critical for our applications because combinations represent hyperplanes, which carry essential predictive information.
-Therefore, we adopt the sequential version of $\mathit{nestedCombs}$ developed by He and Little (2025a).
-The pseudo-code for this sequential version is presented in Algorithm 1.
+Our goal is now to fuse the two-step definition of $\mathit{nestedCombs}\left(K,G\right)$, given in 69 into a single recursive function.
+He (2025b) presents both divide-and-conquer (D&C) and sequential versions for the fused program of $\mathit{nestedCombs}$.
+While the D&C version is more easy to be parallelized, it complicates element ordering.
+Since ordering is critical in our applications—where combinations represent hyperplanes that encode essential predictive information—we adopt the sequential version of $\mathit{nestedCombs}$ developed by He (2025b).
+The pseudo-code for this sequential version is shown in Algorithm 1.
 One source of efficiency in $\mathit{nestedCombs}$ is the representation of inner combinations as integers ranging from $1$ to $\left(\begin{array}[]{c}N\\ G\end{array}\right)$.
@@
   13. 13.
-       $C_{1}=\left(\begin{array}[]{c}G\\ D\end{array}\right)$, $C_{2}=\left(\begin{array}[]{c}G+1\\ D\end{array}\right)$
+          $C_{1}=\left(\begin{array}[]{c}n\\ G-1\end{array}\right)$, $C_{2}=\left(\begin{array}[]{c}n\\ G\end{array}\right)$
   14. 14.
@@
 Algorithm 2 $\mathit{nestedCombs}^{\prime}$
-One advantage of constructing algorithms based on a generic framework is the flexibility it provides.
-In theory, since the input to $\mathit{odt}_{\text{size}}$ is a list of rules rather than raw data, arbitrary splitting rules satisfying the decision tree axioms can be supplied.
-This allows the construction of decision trees with mixed splitting rules.
-The main challenge lies in generating these splitting rules efficiently from a given data list.
-Fortunately, axis-parallel hyperplanes, general hyperplanes, and hypersurfaces can all be represented as combinations of data items.
+SiSince axis-parallel hyperplanes, general hyperplanes, and hypersurfaces can all be represented as combinations of data items, it is natural to ask whether we can construct a single generator capable of producing all these rules.
+A decision tree built from such a rule list would incorporate mixed splitting rules, thereby offering much greater flexibility.
+Consider a list of polynomial degrees $\mathit{Ms}=\left[M_{1},M_{2},\ldots M_{l}\right]$, which corresponds to a list of integers $\mathit{Gs}=\left[G_{1},G_{2},\ldots G_{l}\right]$ ordered in ascending fashion ($G_{1}<G_{2}\ldots<G_{l}$), with $G_{l}$ the largest.
+We can specify $\mathit{modt}^{\mathit{\mathit{Ms}}}:\mathbb{N}\times\mathcal{D}\to\mathit{DTree}\left(\mathcal{H}^{Ms},\mathcal{D}\right)$ (where $\mathcal{H}^{Ms}=\mathcal{H}^{M_{1}}\cup\mathcal{H}^{M_{1}}\ldots\cup\mathcal{H}^{M_{1}}$ ) as
+$\mathit{modt}_{M}=\mathit{min}_{E_{\text{0-1}}}\circ\mathit{concatMap}_{\mathit{sodt}\left(\mathit{xs}\right)}\circ\mathit{kcombs}\left(K\right)\circ\mathit{gen}_{\mathit{splits}}^{\mathit{Ms}}\circ\mathit{embed}_{\mathit{Ms}}$
+where $\mathit{gen}_{\mathit{splits}}^{\mathit{Ms}}:\mathcal{D}\to\left[\mathcal{H}^{Ms}\right]$ is the generator that produces all possible splitting rules in $\mathcal{H}^{Ms}$.
+Similar to the hypersurface generator, the composed function $\mathit{kcombs}\left(K\right)\circ\mathit{gen}_{\mathit{splits}}^{\mathit{Ms}}$ can also be fused.
 A key observation about Algorithm 1 is that, it can generate not only $G$-combinations but all $G^{\prime}$-combinations for $0\leq G^{\prime}\leq G$.
 However, the nested combination generator $\mathit{nestedCombs}\left(K,G\right)$ uses only the $G$-combinations and ignores smaller $G^{\prime}$-combination ($G^{\prime}<G$).
-To accomodate this, we need to modify specification 47 to the following
+To accomodate this, we need to modify specification 69 to the following
 $\mathit{nestedCombs}^{\prime}\left(K,\mathit{Gs}\right)=\left\langle\mathit{setEmpty}\left(G_{l}\right),\mathit{kcombs}\left(K\right)\cdot!!\left(\mathit{Gs}\right)\right\rangle\cdot\mathit{kcombs}\left(G\right),$ (70)
 where $\left(\mathit{Gs},\mathit{xss}\right)=\mathit{concat}\left[!\left(G,\mathit{xss}\right)\mid G\leftarrow\mathit{Gs}\right]$ and $\mathit{Gs}=\left[G_{1},G_{2},\ldots G_{l}\right]$ is an ascending list of integers ($G_{1}<G_{2}\ldots<G_{l}$), with $G_{l}$ being the largest, corresponds to a list of polynomial degree $\mathit{Ms}=\left[M_{1},M_{2},\ldots M_{l}\right]$.
-Specification 48 selects several $G$-combinations, such that $G\in\mathit{Gs}$, and constructs $K$-nested combinations from them.
-As usual, the naive specification 48 is inefficient.
+Specification 70 selects several $G$-combinations, such that $G\in\mathit{Gs}$, and constructs $K$-nested combinations from them.
+As usual, the naive specification 70 is inefficient.
 He and Little (2025a) provide an efficient solution by defining the recursive pattern of $\mathit{nestedCombs}^{\prime}\left(K,\mathit{Gs}\right)$ as
-$\bigg{<}\mathit{setEmpty}\left(G_{l}\right)\cdot\mathit{KcombsAlg}\left(K\right)\cdot\mathit{Ffst},$
+$\bigg<\mathit{setEmpty}\left(G_{l}\right)\cdot\mathit{KcombsAlg}\left(K\right)\cdot\mathit{Ffst},$ (71)
 $\quad\quad\mathit{\mathit{\mathit{KcombsAlg}}\left(K\right)}\cdot\left\langle\mathit{Kcombs}\left(K\right)\cdot!!\left(\mathit{Gs}\right)\cdot convol_{\text{new}}\left(\cup,\circ\right)\cdot\mathit{Ffst},\mathit{KcombsAlg}\left(K\right)\cdot\mathit{Fsnd}\right\rangle\bigg$>
@@
 Specifically, $convol_{\text{new}}$ is defined as
-$convol_{\text{new}}\left(\cup,\circ,\mathit{css}_{1},\mathit{css}_{2}\right)=\left[\mathit{concat}\left(\mathit{zipWith}\left(\circ,\mathit{init}\left(\mathit{tail}\left(\mathinner{css}\right)\right),\mathit{init}\left(\mathit{tail}\left(\mathit{css}_{2}\right)\right)\right)\right)\mid\mathinner{css}\leftarrow\mathit{inits}\left(\mathit{css}_{1}\right)\right].$ (49)
+$\mathrm{convol}_{\text{new}}\left(\cup,\circ,\mathit{css}_{1},\mathit{css}_{2}\right)=$ (72)
+$\quad\left[\mathit{concat}\left(\mathit{zipWith}\left(\circ,\mathit{init}\left(\mathit{tail}\left(\mathinner{css}\right)\right),\mathit{init}\left(\mathit{tail}\left(\mathit{css}_{2}\right)\right)\right)\right)\mid\mathinner{css}\leftarrow\mathit{inits}\left(\mathit{css}_{1}\right)\right].$
 The sequential definition of $\mathit{nestedCombs}^{\prime}$ is provided in Algorithm 2.
@@
 For a $G$-combinations with rank $r$, its index in the algorithm is $i=r+\mathit{sum}\left[\left(\begin{array}[]{c}N\\ G^{\prime}\end{array}\right)\mid G\in\mathit{Gs},G^{\prime}\leq G\right]$, which corresponds to the calculation in lines 13–16 of the algorithm.
+Composing $\mathit{nestedCombs}^{\prime}\left(K,\mathit{Gs}\right)$ with $\mathit{min}_{E_{\text{0-1}}}\circ\mathit{concatMap}_{\mathit{sodt}\left(\mathit{xs}\right)}$, we obtain the algorithm for solving the optimal decision tree problem with mixed splitting rules (MODT)
+$\mathit{modt}_{K}^{\mathit{Ms}}=\mathit{min}_{E_{\text{0-1}}}\circ\mathit{concatMap}_{\mathit{sodt}\left(\mathit{xs}\right)}\circ\mathit{kcombs}\left(K\right)\circ\mathit{nestedCombs}^{\prime}\left(K,\mathit{Gs}\right)\circ\mathit{embed}_{\mathit{Ms}}$ (73)
 5.3 Complexity of solving optimal decision tree problems
@@
 Furthermore, the average-case complexity of $\mathit{sodt}_{\text{rec}}$ is substantially lower than its theoretical upper bound.
-The worst-case complexity of $\mathit{sodt}_{{}^{\prime}\text{rec}}$ occurs only when the tree forms a “chain” structure, which, as shown in the empirical analysis in Part II, is a rare occurrence.
+The worst-case complexity of $\mathit{sodt}_{\text{rec}}$ occurs only when the tree forms a “chain” structure, which, as shown in the empirical analysis in Part II, is a rare occurrence.
 Complexity of the axis-parallel decision tree problem
@@
 In other words, we have
-$\mathcal{S}_{\text{depth}}\left(d,\mathit{xs}\right)\geq\mathcal{S}_{\text{size}}\left(d,\mathit{unique}\left(\mathit{gen}_{\mathit{splits}}\left(\mathit{xs}\right)\right)\right).$ (53)
-where $\mathit{unique}\left(\mathit{gen}_{\mathit{splits}}\left(\mathit{xs}\right)\right)$ denote the all possible unique splitting rules generated by $\mathit{gen}_{\mathit{splits}}\left(\mathit{xs}\right)$, as there might have duplicate splitting rules (for categorical data, distinct data points may have same feature values).
+$\mathcal{S}_{\text{depth}}\left(d,\mathit{xs}\right)\geq\mathcal{S}_{\text{size}}\left(d,\mathit{unique}\left(\mathit{gen}_{\mathit{splits}}^{0}\left(\mathit{xs}\right)\right)\right).$ (77)
+where $\mathit{unique}\left(\mathit{gen}_{\mathit{splits}}^{0}\left(\mathit{xs}\right)\right)$ denote the all possible unique splitting rules generated by $\mathit{gen}_{\mathit{splits}}^{0}\left(\mathit{xs}\right)$, as there might have duplicate splitting rules (for categorical data, distinct data points may have same feature values).
 For the axis-parallel decision tree problem, any pair of splitting rules $r_{i}$ and $r_{j}$ defined by two data points $x_{i}$ and $x_{j}$, satisfies both $r_{i}\swarrow r_{j}$ and $r_{i}\searrow r_{j}$.
@@
 This implies that any $d$-combination of splitting rules can construct at least one valid decision tree, thus
-$\mathcal{S}_{\text{size}}\left(d,\mathit{unique}\left(\mathit{gen}_{\mathit{splits}}\left(\mathit{xs}\right)\right)\right)\geq\mathcal{S}_{\text{kcombs}}\left(d,\mathit{unique}\left(\mathit{gen}_{\mathit{splits}}\left(\mathit{xs}\right)\right)\right).$ (54)
+$\mathcal{S}_{\text{size}}\left(d,\mathit{unique}\left(\mathit{gen}_{\mathit{splits}}^{0}\left(\mathit{xs}\right)\right)\right)\geq\mathcal{S}_{\text{kcombs}}\left(d,\mathit{unique}\left(\mathit{gen}_{\mathit{splits}}^{0}\left(\mathit{xs}\right)\right)\right).$ (78)
 Therefore, the following inequality holds:
-$\mathcal{S}_{\text{depth}}\left(d,\mathit{xs}\right)\geq\mathcal{S}_{\text{size}}\left(d,\mathit{unique}\left(\mathit{gen}_{\mathit{splits}}\left(\mathit{xs}\right)\right)\right)\geq\mathcal{S}_{\text{kcombs}}\left(d,\mathit{unique}\left(\mathit{gen}_{\mathit{splits}}\left(\mathit{xs}\right)\right)\right)=\left(\begin{array}[]{c}\left|\mathit{unique}\left(\mathit{gen}_{\mathit{splits}}\left(\mathit{xs}\right)\right)\right|\\ d\end{array}\right)$ (55)
-As discussed, for axis parallel decision tree problem, $\left|\mathit{gen}_{\mathit{splits}}\left(\mathit{xs}\right)\right|$ is the sum of all distinct features with respect to the training data list $\mathit{xs}$, which is at most $N\times D$ for a size $N$ data list in $\mathbb{R}^{D}$.
+$\mathcal{S}_{\text{depth}}\left(d,\mathit{xs}\right)\geq\mathcal{S}_{\text{size}}\left(d,\mathit{unique}\left(\mathit{gen}_{\mathit{splits}}^{0}\left(\mathit{xs}\right)\right)\right)\geq$ (79)
+$\quad\mathcal{S}_{\text{kcombs}}\left(d,\mathit{unique}\left(\mathit{gen}_{\mathit{splits}}^{0}\left(\mathit{xs}\right)\right)\right)=\left(\begin{array}[]{c}\left|\mathit{unique}\left(\mathit{gen}_{\mathit{splits}}^{0}\left(\mathit{xs}\right)\right)\right|\\ d\end{array}\right)$
+As discussed, for axis parallel decision tree problem, $\left|\mathit{unique}\left(\mathit{gen}_{\mathit{splits}}\left(\mathit{xs}\right)\right)\right|$ is the sum of all distinct features with respect to the training data list $\mathit{xs}$, which is at most $N\times D$ for a size $N$ data list in $\mathbb{R}^{D}$.
 Remarks on Quant-BnB algorithm
@@
 The initial call to count_remain_3D, calculates this complexity directly from the input data without bounding parameters.
-For almost all datasets we have tested, count_remain_3D computed results that is strictly smaller than $\left(\begin{array}[]{c}\left|\mathit{gen}_{\mathit{splits}}\left(\mathit{xs}\right)\right|\\ d\end{array}\right)$.
-Interested readers are encouraged to test this result on their own datasets.
-As explained above, this should be impossible because every $d$-combination of splitting rule should construct at least one possible proper decision tree.
-We hypothesize that Mazumder et al.
-(2022) either employed an incorrect algorithmic procedure for solving the AODT problem or addressed a specialized ODT problem.
-However, their ambiguous problem definition hinders verification of their claims.
-Consequently, there is an urgent need for Mazumder et al.
-(2022) to clarify their counting formula to validate the correctness of their algorithms.
+For almost all datasets we have tested, count_remain_3D computed results that is strictly smaller than $\left(\begin{array}[]{c}\left|\mathit{unique}\left(\mathit{gen}_{\mathit{splits}}^{0}\left(\mathit{xs}\right)\right)\right|\\ d\end{array}\right)$.
+For instance, in the bidding dataset from Mazumder et al.
+(2022)’s GitHub repository, we observed the numbers of unique values per dimension as $\mathit{l}=\left[452,377,3,4680,49,22,4589,72,5\right]$ for this nine-dimensional dataset.
+Thus, the total number of distinct splitting rules is $\mathit{unique}\left(\mathit{gen}_{\mathit{splits}}^{0}\left(\mathit{xs}\right)\right)=\mathit{sum}\left(\mathit{l}\right)=10249$.
+A depth-3 decision tree should contain at least three splitting rules, so the search space should contain at least $\left(\begin{array}[]{c}10249\\ 3\end{array}\right)=179376727124$ possible decision trees.
+However, the first call to count_remain_3D (before any search-space reduction) reports “Total number of trees = 3781512036”.
+This contradicts our lower bound, since every $3$-combination of splitting rule should construct at least one possible decision tree with depth smaller than 3.
+Our analysis suggests a potential discrepancy in the algorithmic approach used by Mazumder et al.
+(2022) for solving the AODT problem, or it may indicate that their work addresses a specialized variant of the ODT problem.
+The lack of clarity in their problem definition makes it challenging to verify their claims.
+Further clarification of their counting formula would be valuable to confirm the validity of their approach and advance the understanding of this problem.
 6 Conclusion
@@
 We provide comprehensive experimental results analyzing the computational efficiency of $\mathit{odt}_{\text{vec}}$ and $\mathit{odt}_{\text{rec}}$.
-Additionally, we conduct a thorough assessment of the generalization capabilities of hypersurface decision tree models and axis-parallel decision tree algorithms, including those learned by CART and Brița et al.
+Additionally, we conduct a thorough assessment of the generalization capabilities of hypersurface decision tree models and axis-parallel decision tree algorithms, including those learned by CART and Brita et al.
 (2025)’s ConTree algorithm, using both synthetic and real-world datasets.